### v1

# DEER: A Comprehensive and Reliable Benchmark for Deep-Research Expert Reports

Abstract
As large language models (LLMs) advance, deep research systems can generate expert-level reports via multi-step reasoning and evidence-based synthesis, but evaluating such reports remains challenging. Existing benchmarks often lack systematic criteria for expert reporting, evaluations that rely heavily on LLM judges can fail to capture issues that require expert judgment, and source verification typically covers only a limited subset of explicitly cited statements rather than report-wide factual reliability. We introduce DEER, a benchmark for evaluating expert-level deep research reports. DEER comprises 50 report-writing tasks spanning 13 domains and an expert-grounded evaluation taxonomy (7 dimensions, 25 sub-dimension) operationalized into 130 fine-grained rubric items. DEER further provides task-specific expert guidance to help LLM judges assess expert-level report quality more consistently. Complementing rubric-based assessment, we propose a document-level fact-checking architecture that extracts and verifies all claims across the entire report, including both cited and uncited ones, and quantifies external-evidence quality. DEER correlates closely with human expert judgments and yields interpretable diagnostics of system strengths and weaknesses.
DEER: A Comprehensive and Reliable Benchmark
for Deep-Research Expert Reports
Janghoon Han1, Heegyu Kim1, Changho Lee1, Dahm Lee1, Min Hyung Park1, Hosung Song1, Stanley Jungkyu Choi1, Moontae Lee1,2, Honglak Lee1,3 1LG AI Research, 2University of Illinois Chicago, 3University of Michigan, Ann Arbor Correspondence: hanjh9439@gmail.com
1 Introduction
Driven by the rapid advancement of large language models (LLMs), automated deep research systems are emerging as a core technology in both academia and industry openai2025deepresearch; google2025gemini; anthropic2025claude; qwen3technicalreport; webthinker; manusearch; webweaver. Unlike simple information retrieval, deep research systems tackle complex research queries by decomposing them into multiple steps and dynamically seeking additional information based on intermediate results. This enables them to integrate information from diverse sources and analyze different perspectives in a balanced manner, producing reliable, evidence-based research reports deepresearch_survey; deepresearch_survey2; characterizing. These advanced capabilities enable deep research systems to achieve strong performance even on challenging benchmark tasks.
Figure 1: Deep Research System Performance Comparison. Shows the performance by type for 5 major models in the proposed benchmark.
Early evaluation of deep research systems relied on complex reasoning benchmarks gpqa; hle, which indirectly assessed information-gathering, hypothesis testing, and multi-step reasoning through performance on these tasks. Subsequently, deep web search QA benchmarks were developed to directly assess systems’ web browsing and information retrieval abilities—core capabilities of deep research—by measuring multi-step search, information integration, and answer derivation browsecomp; frames; gaia; browsecomp_plus; mind2web2. More recently, deep research report benchmarks have emerged to evaluate the quality of generated reports from multiple perspectives, going beyond simple short-answer evaluation webthinker; deepbench; deepconsult; deepgym.
Despite these advancements, current methodologies for evaluating deep research systems still face key limitations in assessing expert reports. First, while prior work has proposed high-level dimensions for report evaluation, these frameworks have yet to comprehensively specify the diverse elements required for expert-level reports. In particular, many frameworks deepbench; DeepScholar-Bench; LiveResearchBench; DeepResearchArena rely on LLMs—rather than human experts—to instantiate detailed sub-dimension and concrete evaluation items, which can lead to inconsistencies and reduced reliability. Second, even with a fixed rubric, evaluations that rely on LLM judges can fail to capture issues that require expert judgment. Third, current approaches to source verification are typically limited to sections with explicit citation markers, leaving report-wide claim reliability insufficiently examined. Together, these limitations hinder comprehensive and reliable evaluation of deep research systems.
To address these limitations, we propose the Deep rEsearch Expert Report benchmark (DEER), consisting of 50 report-generation tasks spanning 13 domains, for comprehensive and reliable evaluation. DEER analyzes a wide range of expert report assessment criteria and consolidates them through expert consensus into a Deep Research Report Evaluation Taxonomy with 7 dimensions and 25 sub-dimension. Based on this taxonomy, DEER evaluates each report along two complementary components: (i) report-quality dimensions that require holistic document-level judgment and (ii) external-information dimensions that can be assessed at the claim level via fact verification. For report quality, we operationalize the taxonomy into a fixed set of 130 fine-grained rubric items and supplement it with report-specific Expert Evaluation Guidance authored by domain experts to improve the consistency and validity of LLM-judge-based scoring. For external information, we introduce a fact-verification module that checks all cited and uncited claims across the report and produces quantitative measures of external-evidence quality. Finally, DEER combines these metric-based signals with rubric-based LLM evaluations to yield an integrated, multi-dimensional assessment of each expert report. Figure 1 presents an overview of dimension-level performance for major deep research systems evaluated on DEER.
The main contributions of this study are as follows:
  • We present DEER, a systematic and interpretable benchmark for deep research, built on a hierarchical evaluation taxonomy.
  • We instantiate this taxonomy into a fixed set of 130 fine-grained rubric items and supplement it with report-specific Expert Evaluation Guidance, improving the consistency and reliability of LLM-judge-based evaluation.
  • We propose a fact verification architecture that backtracks inter-claim dependency relations to retrieve and verify relevant citations, enabling more complete report-wide evaluation of all claims.
2 Related Works
2.1 LLM-based Evaluation Automation
With the proliferation of Large Language Models, various approaches have been proposed to automate human-centric evaluation tasks using LLMs. Most notably, the LLM-as-a-Judge method, which directly uses LLMs as evaluators, has been widely researched. JudgeLM zhu2023judgelm and G-Eval liu2023g demonstrate that models can score generated responses based on multiple criteria such as accuracy, consistency, and logic. In addition, the scope of automated LLM evaluation is expanding to pairwise comparison-based evaluations such as Chatbot Arena chiang2024chatbot, CheckList-style robustness and bias checks checkeval, and various automated scoring frameworks. However, these approaches often struggle to control bias or hallucinations in the LLM’s own judgments and do not adequately address evaluations that require complex evidence structures, such as long-form reports.
2.2 Deep Research and High-Difficulty QA Evaluation
Recently, Deep Research benchmarks have emerged that go beyond simple knowledge recall to evaluate complex reasoning, web browsing, and information integration capabilities. High-difficulty QA benchmarks such as HLE hle, GPQA gpqa, and FRAMES frames reveal the limitations of existing models by requiring long-context reasoning, expert-level thinking, and information synthesis skills. Furthermore, ManuSearch manusearch and Mind2Web-2 mind2web2 employ multi-agent web browsing and agent-as-a-judge architectures to evaluate models’ ability to replicate real-world research processes. Additionally, DeepResearchBench deepbench and WebThinker Eval webthinker present Deep Research report-centric evaluations that assess literature-based report generation and multi-document reasoning, classifying them as measuring "research capability" beyond single-answer QA. However, while these benchmarks evaluate correctness, they cannot precisely verify the validity of evidence, citations, and arguments within the report, focusing more on whether the correct answer is reached than on factuality.
2.3 Claim Verification
In automated claim verification, various approaches have been proposed to mitigate LLM hallucinations. TrustGPT huang2023trustgpt attempted to improve factuality based on RAG, and InFact rothermel-etal-2024-infact established a structural verification system by decomposing claim verification into a 6-step pipeline. Additionally, MAPLE zeng2024maple, FactDetect jafari2024robustclaimverificationfact, and ClaimCheck putta-etal-2025-claimcheck implemented sophisticated claim-level fact-checking by combining claim type classification, evidence extraction, and verification automation. Furthermore, research analyzing citation quality in detail, such as ALiiCE xu2024aliiceevaluatingpositionalfinegrained and CiteEval xu2025citeevalprincipledrivencitationevaluation, which evaluate citation accuracy and evidence alignment in academic texts, is increasing. However, most approaches focus on verifying explicit citations or single claims and fail to address issues such as implicit evidence, contextual citation inheritance, and Fair Use that appear in long-form reports.
3 Data Construction
To construct an evaluation dataset that reflects real-world usage, we analyzed 5,045 user queries collected from a deep research system and estimated the resulting topic distribution. For topic classification, we used the taxonomy of weborganizer, following prior work deepbench.
We use Humanity’s Last Exam (HLE) hle as a source of seed questions because its expert-written, high-difficulty, multi-disciplinary questions align with the expert-level topics addressed in deep research reports. To match our target topic distribution, we mapped our topic domains to the 13 HLE subject domains and sampled 50 seed questions accordingly.
Because HLE items are posed in a QA format, they require reformulation before they can serve as deep research report-generation tasks. We therefore asked domain experts (with at least a master’s degree or equivalent expertise in the relevant field) to review each item’s question, answer, and rationale to identify the underlying concepts, theories, and phenomena, and to reformulate the items as report- or paper-style topics for report generation. Each reformulated prompt was drafted by one domain expert and cross-reviewed by another expert from the same field, with iterative revisions as needed. During reformulation, we removed answer-revealing elements (e.g., factual conclusions, proofs, or specific answers) and instead provided only high-level writing directions, such as the intended scope of analysis and comparative perspectives, so that the model must develop the reasoning and narrative on its own. This reformulation shifts the task from reproducing a short answer to generating a report that requires expert-level analysis, reasoning, and interpretation. Detailed content is presented in Appendix B.
4 Approach
4.1 Overview
Figure 2: Overview of the DEER evaluation framework. (a) Research question and expert guidance generation from real-world deep research queries. (b) Construction of the Deep Research Evaluation Taxonomy consisting of 7 dimensions, 26 criteria, and 130 granular rubrics. (c) The DEER evaluation pipeline, integrating expert-guided LLM-as-a-judge scoring with claim extraction and information verification to assess deep research reports.
To systematically evaluate deep research reports, we establish an evaluation framework grounded in two core aspects of deep research: external evidence acquisition and report-level synthesis characterizing; deepresearch_survey; deepresearch_survey2. Based on these aspects, we construct a Deep Research Report Evaluation Taxonomy with 7 major dimensions (§ 4.2), which we group into 5 report-quality dimensions and 2 external-information dimensions. Report-quality dimensions primarily assess synthesis and presentation via holistic, document-level judgment, whereas external-information dimensions assess external evidence acquisition and use via claim-level verification.
Based on these differences in evaluation characteristics, we propose a hybrid architecture that integrates methodologies tailored to each component, as illustrated in Figure 2. Expert-Guided Report Quality Assessment (§ 4.3) uses an LLM-as-a-judge approach, combining a fixed set of granular evaluation items with report-specific Expert Evaluation Guidance authored by domain experts, to assess report quality against expert-designed criteria. Information Verification (§ 4.4) performs metric-based evaluation by automatically extracting claims and citations and verifying claims against external sources, producing quantitative metrics of external evidence sufficiency and integrity. Scores from the rubric items and verification metrics are aggregated into 7 dimension scores, which are then combined into an overall report score.
Major Dimensions Sub-criteria
Request Completeness Request Completeness, Scope Boundary, Value Validity
Evidence Validity Numeric Accuracy, Logical Support
Structure Consistency Structure Intro, Structure Body, Structure Conclusion, Structure Section
Narration Style Report Form, Writing Quality, Paragraph, Reader Friendliness
Ethics Compliance Sensitive Issues, Safety Impact, Perspective Balance
Information Sufficiency Source Support, Information, Citations, References
Information Integrity Claim Factuality, Citation Validity, Reference Accuracy, Reference Quality, Reference Diversity
Table 1: Deep Research Report Evaluation Taxonomy: 7 Major Dimensions and 25 Sub-criteria
4.2 Deep Research Report Evaluation Taxonomy
In the absence of systematic evaluation criteria for deep research reports, we construct an evaluation taxonomy by synthesizing expert report assessment standards from multiple fields. We draw on internationally established standards spanning 20 domains of expertise, including systematic research reporting guidelines, technical and professional report writing and evaluation guidelines, and academic publishing norms, for a total of 80 standards. A panel of experts with experience in deep research system development and academic reviewing iteratively analyzes and consolidates these standards to derive report-quality dimensions. We then introduce complementary external-information dimensions that reflect deep research–specific characteristics of external evidence use and verification, resulting in 7 major dimensions and 25 detailed criteria in total (Table 1). We validate the taxonomy through independent review by 10 experts from diverse domains. This hierarchical taxonomy organizes diverse quality elements into structured evaluation dimensions and sub-dimension, enabling systematic, multi-dimensional evaluation. It facilitates diagnosis of report strengths and weaknesses and guides systematic improvement. Full descriptions, validation procedures, and the mapping from source standards to our criteria are provided in Appendix C.
4.3 Expert-Guided Report Quality Assessment
Existing evaluation methods have two limitations. First, existing methods apply broad, high-level criteria to long-form reports, giving LLM judges substantial discretion over what to focus on. As a result, judges may (i) examine only parts of a report, (ii) consider only a subset of the many sub-aspects implied by each criterion, and (iii) even focus on non-critical or superficial sub-aspects, increasing variance across judges and reducing evaluation consistency webthinker; deepconsult; deepgym. Second, even with a clear evaluation focus, assessing correctness and completeness often requires domain knowledge, which LLM judges may lack. As a result, they may miss subtle but important issues such as non-obvious logical leaps, domain-specific misinterpretations, and fine-grained inaccuracies deepbench. This study addresses these issues by (1) making the evaluation well-specified through a fixed set of fine-grained rubric items and (2) providing task-specific Expert Evaluation Guidance that includes domain-specific context and reference points, helping LLM judges identify hard-to-detect errors and omissions.
Granular Rubric Design. Broad evaluation criteria can lead to inconsistent LLM-judge-based scoring. Recent work addresses this by decomposing broad criteria into granular rubrics checkeval; rocketeval; expertlongbench. For deep research report evaluation, deepbench follows a similar direction by using an LLM to generate task-specific evaluation criteria. While this narrows the evaluation focus, it raises two concerns: whether the generated criteria are sufficiently complete and capture what matters most in expert reports, and whether dynamically changing criteria enable interpretable, comparable scores that support systematic diagnosis of system strengths and weaknesses across tasks.
This study uses a fixed, expert-designed rubric to ensure reliable and interpretable evaluation. Building on the taxonomy in Table 1, a panel of experts operationalizes the 25 sub-dimensions into 40 evaluation criteria and translates each criterion into concrete, checkable rubric items. Rubric items are organized into two aspects: coverage (whether the required components are present at all relevant points in the report and addressed in full whenever they occur) and quality (the degree to which the targeted components are executed to a high standard in the report). This process yields 130 scorable rubric items in total. [^1] The 40 criteria were validated through independent review by 10 experts following the procedure in § 4.2. The rubric is applied identically to all deep research reports across 50 tasks, enabling diagnosis of the system’s overall strengths and weaknesses with richer, more interpretable signals (see Appendix E for the rubric structure and examples).
Expert Evaluation Guidance. Even with a well-specified, granular rubric, scoring individual items can still require judgment informed by deep domain knowledge, especially for specialized and complex tasks—for example, to assess whether the report’s reasoning is sound, its calculations are correct, and its coverage and depth are adequate for the topic. In such cases, evaluators may miss subtle domain-specific errors or omissions that human experts would catch. To mitigate this, we provide task-specific expert evaluation guidance that enumerates the required content elements and expert expectations for each task, presented as concrete, checkable statements that evaluators can verify.
For each task, an expert with at least a master’s degree in the relevant field was assigned to detail the key concepts, theories, data, and analytical perspectives that the report must cover from an expert perspective. The expert first modified the original query to naturally include the essential content of the topic (Section 3), and then wrote the Expert Evaluation Guidance based on the modified query, specifying the content elements that must be included. This approach is based on the logic that if a report responds appropriately to the modified query, it should naturally contain the key content of the topic, so the omission of key content can be clearly judged as a quality issue. Each written Guidance was verified through cross-review by another expert in the same field. See Appendix B for detailed methodology.
4.4 Information Verification
To automate fact verification in expert reports, we must overcome the limitations of existing systems that focus solely on explicit citations. In reality, many sentences implicitly inherit preceding evidence or construct new claims through contextual connections. Excluding these implicit claims from verification makes the reliability assessment of the entire document incomplete. Our Information Verification Module proposes an architecture that extends the verification scope to the entire document by tracking sentence-level logical flows.
Implicit Claim & Back-Tracking
This module analyzes contextual correlations to convert even uncited sentences into verifiable units. Specifically, when sentence $s_{i}$ implicitly cites the content of a previous sentence or paragraph, we use a Back-Tracking mechanism to trace the reference source. To this end, the LLM identifies the set of preceding sentences $R(s_{i})$ that $s_{i}$ depends on, and the valid citation set $\mathcal{V}(s_{i})$ used for actual verification is defined as follows:
$\mathcal{V}(s_{i})=\mathcal{C}(s_{i})\cup\bigcup_{k\in R(s_{i})}\mathcal{C}(s_{k})$ (1)
where $\mathcal{C}(s_{i})$ is the explicit citation of $s_{i}$. This structure allows verification by inheriting citations from referenced preceding sentences even if the sentence itself has no citation ($\mathcal{C}(s_{i})=\emptyset$). This is a key element that substantially expands the verification scope in policy and research reports containing many implicit claims.
Batch Processing for Claim Extraction
When processing long documents as a single input, LLMs suffer from the Lost-in-the-Middle problem, which leads to the omission of important claims liu2023lostmiddlelanguagemodels. To address this, we apply a Batch Extraction strategy that maintains the context of the entire document $D$ while parallel-processing only the target sentences divided into small batches $B_{j}$:
$\mathcal{A}=\bigcup_{j=1}^{M}E(D,B_{j}),\quad D=\bigcup_{j=1}^{M}B_{j}$ (2)
This increases the model’s focus to prevent missing important claims, and experimental results showed a significant improvement in extraction recall compared to the single processing ($E(D,D)$) method. Although the number of API calls increases, we confirmed that this is an essential strategy for securing high coverage in reports dealing with hundreds of claims.
Scalable Verification Architecture
The extracted claims are passed to the fact verification stage, where the biggest technical challenge is the trade-off between accuracy and cost. Verifying every claim individually yields high accuracy, but cost increases linearly with long documents. To address this, we introduce two scaling strategies.
First, we apply Grouped Claim Verification, which bundles multiple claims citing the same source into a single input for verification. Experimental results showed that increasing Group Size from 1 to 10 raised accuracy from 0.78 to 0.89 and reduced cost per claim by 85% (Section 6.5). This indicates that grouped input helps in understanding the overall structure of the document and verifying claims, while reducing verification overhead. Second, greater API cost reduction is possible through the Context Retrieval strategy, which selectively provides only the context necessary for fact verification. This method achieved an API cost reduction of approximately 24 times, but a 6% drop in accuracy was observed, indicating a clear trade-off.
This hierarchical pipeline (Batch Extraction $\rightarrow$ Grouped Verification $\rightarrow$ Context Retrieval) is designed to achieve both high-precision verification and cost efficiency even in long reports containing hundreds of claims.
5 Experiment Setup
5.1 Implementation Details.
We use the LLM-as-a-judge method to evaluate report quality. The evaluation model takes as input the task query, the report to be evaluated, expert guidance, and detailed rubrics and evaluation criteria. The evaluation is conducted across 5 dimensions: Request Completeness, Evidence Validity, Structure Consistency, Narration Style, and Ethics Compliance, with each rubric item scored on a 1–10 scale. The model generates a score and a short rationale justifying that score for each item, and the final output is in JSON format consisting of (rationale, score) pairs for each item.
Information Integrity and Information Sufficiency are evaluated through a separate verification pipeline, the Information Verification Module. This module extracts claims from the report and assesses their factuality by comparing them with external evidence obtained via web searches. The verification results are calculated as detailed metrics such as Claim Factuality and Citation Validity.
The evaluation is hierarchical, computing the final quality score by sequentially aggregating the averages of lower-level rubric items to higher levels, using the scores collected above.
Model Req. Comp. Evid. Valid. Struct. Cons. Narr. Style Info. Int. Info. Suff. Ethics Mean
General LLM
qwen3-235b-fast 6.10 6.59 9.04 8.66 3.30 4.40 8.58 6.67
gemini-2.5-pro-fast 6.08 7.02 9.29 8.87 2.92 4.43 8.89 6.79
claude_opus4.1_fast 6.04 6.47 8.96 8.95 4.76 4.81 8.78 6.97
gpt5_fast 5.61 6.63 8.80 8.46 2.18 3.45 8.82 6.28
LLM+Thinking
qwen3-235b-think 6.77 6.53 9.32 8.92 3.28 4.39 8.75 6.85
gemini-2.5-pro-think 6.41 7.66 9.55 9.20 3.86 4.97 9.37 7.20
claude_opus4.1_think 5.93 6.41 9.05 9.04 4.72 5.12 8.99 7.04
gpt5_think 7.69 8.37 9.57 9.32 4.19 5.42 9.37 7.70
LLM+Thinking+Search
qwen3-235b-think_search 6.43 6.31 9.07 8.82 3.60 4.49 8.44 6.74
gemini-2.5-pro-think_search 6.61 7.58 9.50 9.11 4.62 5.02 8.97 7.34
claude_opus4.1_think_search 6.03 6.62 9.04 9.03 5.17 4.80 8.85 7.08
gpt5_think_search 7.88 8.26 9.39 9.04 6.42 6.44 9.38 8.12
Deep Research
webthinker 5.27 5.76 8.48 8.62 7.57 6.92 8.68 7.33
qwen3-235b-deep 4.95 6.00 7.28 8.31 7.13 8.29 8.66 7.23
gemini-2.5-pro_deep 6.08 6.81 8.81 8.78 7.13 7.87 8.62 7.73
claude_opus4.1_deep 6.60 6.91 8.90 8.85 6.79 7.66 8.75 7.78
gpt5_deep 6.26 7.11 9.20 8.93 7.64 7.13 8.99 7.89
Table 2: Baseline models’ expert report evaluation results. In each column, the highest score is marked in bold and the second highest score is underlined.
5.2 Baseline Models
In this study, to compare the expert report generation capability of LLM-based systems, we established four categories of baseline models. Each category is designed to increase in complexity incrementally, ranging from simple language model response generation to specialized research systems that combine search and deep reasoning.
(1) General LLMs This is the basic LLM setting that generates responses in a single pass (inference), serving as a baseline for measuring response quality in language-generation-based models. This category includes qwen3-235b-fast qwen3technicalreport, gemini-2.5-pro-fast google2025gemini, claude_opus4.1_fast anthropic2025claude, and gpt5_fast openai2025gpt5system models.
(2) LLMs with Reasoning This setting uses the reasoning function on the same LLMs as in (1), extending the internal reasoning step before answer generation. This category includes qwen3-235b-think, gemini-2.5-pro-think, claude_opus4.1_think, and gpt5_think models.
(3) LLMs with Reasoning + Web Search This configuration combines the LLMs with the Reasoning setting and web search, evaluating scenarios in which external information is used during reasoning. This category includes qwen3-235b-think_search, gemini-2.5-pro-think_search, claude_opus4.1_think_search, and gpt5_think_search models.
(4) Deep Research Frameworks (Open-Source & Commercial) This is the upper baseline, comprising specialized deep research frameworks and commercial research support systems. We used webthinker webthinker as the open-source framework, and commercial systems include qwen3-235b-deep, gemini-2.5-pro_deep, claude_opus4.1_deep, and gpt5_deep. Detailed configurations and hyperparameters for all baseline models are summarized in Appendix A.
6 Experiments
6.1 Main Results
Summarizing Table 2, all models show generally high performance in metrics related to form, style, and ethics, such as Structure Consistency (Struct. Cons.), Narration Style (Narr. Style), and Ethics Compliance (Ethics). However, they consistently show relatively lower scores on metrics that assess how faithfully they address the requirements in the user’s query and provide persuasive evidence in terms of calculation, methodology, and logical development, such as Request Completeness (Req. Comp.) and Evidence Validity (Evid. Valid.). This indicates that while the latest LLMs are already quite proficient in structuring reports, writing naturally, and adhering to basic ethical standards, their ability to accurately meet user intent at an expert level and support it with sufficiently robust argumentation and evidence structures remains limited. Reasoning models showed consistent improvements across almost all evaluation metrics relative to General LLMs, suggesting that simply expanding the reasoning budget without additional tools improves coverage of requirements, logical development, evidence selection, and narrative consistency. Models combining this with web search show additional gains in information-related metrics. The Deep Research category consistently achieves higher scores in Information Integrity (Info. Int.) and Information Sufficiency (Info. Suff.) than other categories, providing the strongest baseline for the broad collection and organization of external documents.
However, it is characteristic that the Deep Research category consistently records lower scores than reasoning-based single LLMs in the remaining key evaluation axes, especially Req. Comp. and Evid. Valid. These results suggest that the large amount of external material injected through search in the Deep Research configuration may cause the reference point for report development to gradually shift from the query itself to external knowledge, inducing a kind of retrieval-induced drift in which information rather than blurs the problem definition and argumentation structure. Therefore, this demonstrates that simply increasing the volume of search and citations does not guarantee the quality of expert reports.
6.2 Fine-grained Analysis of Evaluation Results
Figure 3: Heatmap visualizations of expert report evaluation results. (a) Criteria-wise scores across detailed evaluation categories. (b) Domain-wise scores averaged over two representative tasks randomly sampled from each domain.
Figure 3(a) decomposes the upper evaluation dimensions into detailed items to summarize the performance patterns of current Deep Research and LLM-based systems in expert reports. In Request Completeness, scores across sub-items are generally low, with scope the lowest, suggesting a lack of clarity about what the report covers and excludes, and under what assumptions and limitations the discussion is developed. In Evidence Validity, the score for the numeric item is slightly lower than that for logical, indicating that current systems make mistakes relatively more frequently in terms of quantitative accuracy, such as calculation processes, used formulas/statistical models, and selection of indicators/units, rather than logical development itself. In Information-related detailed items, ref_diversity and refs_amount appear commonly low, meaning that even if the overall structure and narrative are compliant, the number and diversity of sources actually referenced are limited.
Figure 3(b) shows the domain-specific performance of Deep Research agents as the average of two randomly selected tasks in each domain (domain-specific average scores for all tasks are presented in Appendix LABEL:app:domain-results). While most domains are distributed in a relatively similar range around 7 points, in the case of Physics, most Deep Research agents except gpt5_deep consistently record the lowest scores compared to other domains, and relatively large deviations between Deep Research agents also appear in CS, History, and Linguistics.
6.3 Correlation with Human Evaluation
To verify the validity of the proposed evaluation method and analyze the contribution of each component, we conducted experiments to assess its correlation with human expert evaluation. For 15 tasks across 5 domains, 3 reports randomly selected from those generated by the Deep Research systems used in 6.1 were used as evaluation targets. Each report was evaluated by two independent experts, collecting a total of 90 human evaluation samples. To assess the alignment between the evaluation method and human experts, we used the Pearson correlation coefficient, the Spearman rank correlation coefficient, and pairwise agreement deepbench. Pairwise agreement is a metric that measures the proportion of cases in which our evaluator chose the same priority as the human expert when comparing two reports, indicating how consistently it reproduces human relative judgments across tasks. Evaluations were performed on the same reports using 5 LLM-based evaluation models (GPT-5, GPT-5-mini, Claude Opus 4.1, Claude Sonnet 4.5, Gemini 2.0 Pro), and the correlation between the average of the 5 model evaluations and the human evaluation was analyzed.
Evaluation Method Pearson Spearman Pairwise
$r$ $\rho$ Agr.
Vanilla 0.64 0.61 0.66
+ Dimensions 0.67 0.65 0.80
+ Granular Checklist 0.53 0.46 0.72
+ Expert Guidance 0.75 0.71 0.84
Inter-Human 0.81 0.74 0.79
Table 3: Average human correlation across five AI models with incremental addition of evaluation components. Inter-Human shows inter-annotator agreement.
Table 3 shows how the correlation with human evaluation changes as evaluation components are added step by step. Starting from the Vanilla deepbench setting, which judges overall quality without explicit evaluation criteria, we sequentially expanded by introducing evaluation dimensions (+Dimensions), decomposing them into granular rubrics (+Granular Rubric), and finally adding task-specific expert guidance (+Expert Guidance) to measure the impact of each component. The Vanilla setting showed significant correlation with human experts even without explicit criteria, but this correlation decreased as evaluation dimensions (+Dimensions) were introduced and decomposed into detailed items (+Granular Rubric). Since report evaluation is a high-dimensional task that requires simultaneous consideration of various elements, such as logic, concepts, analysis, and domain knowledge, across long and complex texts, each rubric item comprises broad, interpretive judgment units rather than narrow rules. It is analyzed that the increase in the number of items led to an increase in interpretive variability rather than refinement, resulting in a drop in human correlation. Adding task-specific expert guidance (+Expert Guidance) clarifies the interpretation method for each criterion, restoring consistency and significantly improving correlation across all metrics. Writing longer rationales (+Longer Rationale) further improves correlation by enriching the basis for each item. In conclusion, DEER tends to achieve relatively stable criterion application and high alignment with human experts, even in report evaluation settings that must consider multiple evaluation dimensions simultaneously.
6.4 Inter-evaluator Agreement (IEA)
To verify that the proposed evaluation method yields consistent results across different LLMs, we measured inter-evaluator reliability among the 5 evaluation models used previously IAA; checkeval. We measured the agreement between models using Krippendorff’s $\alpha$ and ICC (Intraclass Correlation Coefficient). High agreement suggests that the proposed method possesses versatility and stability that are not dependent on a specific LLM.
Method Krip. $\alpha$ ICC(2,1) ICC(2,k)
Vanilla 0.46 0.48 0.82
+ Dimensions 0.32 0.37 0.75
+ Granular Checklist 0.17 0.24 0.62
+ Expert Guidance 0.55 0.56 0.87
Table 4: Inter-evaluator reliability across five LLM-based evaluation models. Krip. $\alpha$: Krippendorff’s alpha. Higher values indicate more consistent evaluations.
Table 4 shows how the agreement between evaluation models changes as evaluation components are added step by step. The Vanilla setting showed an appropriate level of inter-model agreement, but agreement decreased significantly when introducing evaluation dimensions (+Dimensions) and decomposing them into detailed items (+Granular Rubric). As explained earlier, this is because rubric items consist of interpretive judgment units, so the increase in the number of items increased the variability in interpretation across models. Adding expert guidance (+Expert Guidance) significantly improved agreement across all metrics, restoring evaluation consistency. In particular, the variation focusing only on the top rubric items achieved the highest agreement. This shows that expert guidance clarifies the interpretation of evaluation criteria, enabling consistent evaluation even between different LLM evaluation models.
6.5 Information Verification Module Evaluation
In this section, we evaluate the proposed Information Verification Module by dividing it into two stages: claim extraction and fact verification.  [^2] Detailed content is covered in Appendix F.
Claim Extraction Analysis
We compared and analyzed 728 claims identified by human experts (Ground Truth) and LLM extraction results using a Paragraph-Level LLM Judge (Table 5). As a result of measuring semantic Recall beyond simple count comparison, GPT-5-mini showed significantly higher performance compared to GPT-4.1. In particular, the GPT-5-mini (Batch 5) setting recorded a Recall of 0.9333 for Type A, B, and C claims requiring external evidence, capturing human expert claims almost perfectly. Also, the number of claims extracted per paragraph (Claims/Para) was the highest at 6.95, confirming that detailed information in the document is extracted densely without omission. On the other hand, GPT-4.1 showed relatively lower Recall (about 0.83) and density (about 4.8), suggesting a tendency for the model to excessively summarize information or omit some details. In conclusion, for precise fact-checking, a strategy of maximizing recall by operating high-performance models like GPT-5-mini with small batch sizes (Batch Size 5, 10) is essential.
Table 5: Comparison of Claim Extraction Performance across Models and Batch Sizes. Recall is measured at the Paragraph-Level using LLM Judge.
Model Batch Size Recall Claims/Para
Ground Truth - - 7.22
GPT-4.1 5 0.9057 5.00
10 0.8897 4.84
20 0.8897 4.65
GPT-5-mini 5 0.9333 6.95
10 0.9264 6.22
20 0.9218 6.05
Claim Verification Performance
In the claim verification stage, we conducted an ablation study on grouping and retrieval settings using GPT-4.1 (Table 6). When applying Grouped Claim Verification without retrieval, the F1 score remained comparable (91 → 90), while the cost per claim decreased by approximately 82%, demonstrating that batching claims provides substantial efficiency gains with minimal performance loss.
Additionally, enabling retrieval further reduced the cost to \$0.004 per claim but lowered the F1 score to 85. Therefore, for scenarios requiring high reliability, grouped verification without a retrieval configuration offers the best balance, while with a retrieval setting, it is preferable when maximizing cost efficiency is the primary objective.
Grouped Retrieval F1 Cost (\$/Claim)
0.91 0.083
20 0.90 0.015
20 0.85 0.004
Table 6: Ablation study on GPT-4.1. Grouping significantly improves cost efficiency with minimal F1 degradation; retrieval further reduces cost at the expense of accuracy.
7 Conclusion
We introduce DEER, a comprehensive benchmark for evaluating Deep Research agents, featuring a multi-dimensional taxonomy and a hybrid evaluation framework that combines LLM-based qualitative assessment with a retrieval-augmented Information Verification Module. Our experiments reveal that while state-of-the-art models excel in structure and style, they still lag behind human experts in critical areas like evidence validity and information sufficiency, although "Deep Research" systems show promising improvements. By demonstrating a high correlation with human judgments, DEER offers a scalable standard for advancing autonomous research agents toward trustworthy, expert-level performance.
Limitations
While our evaluation framework relies on LLM-based judges, which may inherently exhibit biases relative to human experts, our extensive validation shows a high correlation with human judgment, suggesting that these biases are systematic and manageable. Furthermore, although our current benchmark focuses on text-based reports, this specialization enables a deeper, more rigorous analysis of information integrity and logical coherence, laying a solid foundation for future extensions to multimodal research tasks.
Acknowledgments
Benchmark Human rubrics Expert curated Open- ended Non-tech domains LLM judge Claim fact-check Inter- pretability Systematic criteria Avg. rubrics
AcademicBrowse AcademicBrowse
BrowseComp browsecomp
Mind2Web2 mind2web2 $\triangle$ 50
ResearchBench ResearchBench
ResearcherBench ResearcherBench $\triangle$ $\triangle$ 14
LiveDRBench characterizing $\triangle$
SPOT SPOT $\triangle$
ExpertLongBench expertlongbench 16
ReportBench ReportBench $\triangle$
DeepScholar-Bench DeepScholar-Bench $\triangle$ $\triangle$
DeepResearchBench deepbench $\triangle$ $\triangle$ 25
LiveResearchBench LiveResearchBench $\triangle$ $\triangle$
DeepResearchArena DeepResearchArena $\triangle$ $\triangle$
DeepResearchGym deepgym $\triangle$
ResearchRubrics ResearchRubrics $\triangle$ 26
DREX (Ours) 130+
Table 7: Comparison of DREX with representative deep research benchmarks. Here, indicates full support, no support, and $\triangle$ partial support. For Claim fact-check, $\triangle$ means that only a subset of claims (e.g., explicitly cited or gold-labeled ones) are checked rather than all explicit and implicit claims. For Interpretability, $\triangle$ means that the benchmark provides only coarse, dimension-level insight (e.g., a few high-level scores) without shared checklist-level diagnostics across tasks.
Appendix A Comparison with Deep Research Benchmarks
In this section, we compare existing deep research benchmarks with our methodology. Table 7 extends and modifies the comparison table proposed in ResearchRubrics ResearchRubrics, and adds new evaluation axes to summarize how DREX differs from prior work. The ResearchRubrics table uses five axes (whether rubrics are human-written, whether experts curate tasks, whether tasks are open-ended, whether non-technical domains are included, and whether an LLM-as-judge is used). On top of this, we add three axes that we consider essential from the perspective of deep research report evaluation: (1) Claim fact-check, (2) Interpretability, and (3) Systematic criteria. Systematic criteria indicate whether the evaluation criteria form a shared hierarchical ontology across tasks, grounded in external expert norms and guidelines.
The first block of the table consists of benchmarks for evaluating search and browsing agents. AcademicBrowse AcademicBrowse, BrowseComp browsecomp, and Mind2Web2 mind2web2 evaluate, respectively, the ability to browse an academic corpus or the open web to produce short answers to complex queries, the ability to perform agentic search across diverse websites, and the consistency between generated answers and cited sources. These benchmarks focus on “how well the system finds information,” i.e., search and browsing strategies, and thus are one step removed from the deep research report quality evaluation that we study.
The second block covers benchmarks that focus on specific component abilities or conceptualizations of deep research. ResearchBench ResearchBench and ResearcherBench ResearcherBench evaluate, respectively, the ability to generate inspiration and hypotheses from papers, and to provide long-form advice on frontier AI research questions based on expert rubrics. Characterizing Deep Research characterizing and LiveDRBench evaluate the recovery of correct claims in search tasks that require many information units and non-trivial reasoning. SPOT SPOT measures how well a system can detect critical errors in published papers, and ExpertLongBench expertlongbench evaluates the ability to generate long-form expert text without external search. These benchmarks finely probe partial abilities such as hypothesis generation, claim discovery, error detection, and expert writing, but do not directly evaluate the quality of full reports produced by web- or literature-based deep research agents.
The third block targets benchmarks that evaluate the quality of deep research reports themselves. ReportBench ReportBench and DeepScholar-Bench DeepScholar-Bench measure report quality in academic survey and related-work settings using citation coverage, overlap with gold surveys, and verifiability of citations. DeepResearchBench deepbench and LiveResearchBench LiveResearchBench evaluate reports for multi-domain web-based deep research tasks using LLM-generated evaluation criteria (RACE), citation-based fact-checking (FACT), and dimensions such as coverage, depth, presentation, and citation accuracy. DeepResearchArena DeepResearchArena automatically constructs large numbers of research tasks from seminar transcripts, and measures factual alignment and overall quality via coverage/conflict/omission (KSR/KCR/KOR) against evidence keypoints extracted from cited URLs and LLM-generated task-specific rubrics (ACE). DeepResearchGym deepgym provides an offline search environment over a public web corpus and an LLM-as-judge protocol, offering a sandbox for evaluating deep research agents’ reports with respect to alignment with information needs, search faithfulness, and overall quality. ResearchRubrics ResearchRubrics provides 101 deep research prompts and over 2,500 human-written rubrics, and is designed so that an LLM-as-judge assigns per-item scores by following these detailed human criteria.
Among these, we analyze DeepResearchBench, LiveResearchBench, and ResearchRubrics in more detail, as they are closest to DREX in terms of evaluation setting. DeepResearchBench and LiveResearchBench are most similar to our web-based deep research report evaluation setting, but both rely heavily on LLMs to define evaluation criteria and assign scores, and their fact-checking is applied only to a limited subset of claims with explicit citations. ResearchRubrics has the advantage that domain experts write 20–40 rubric items for each deep research prompt, enabling fine-grained capture of what constitutes a “good” answer, but it does not perform source-level claim fact-checking, and the rubric structure differs across prompts, making it challenging to interpret system behavior consistently along shared dimensions and rubrics. In contrast, DREX extracts, as exhaustively as possible, all explicit and implicit claims in a report that can be grounded in evidence documents and performs claim-level fact-checking over this set, and evaluates reports using a shared ontology of 7 high-level areas, 25 criteria, and 130 rubric items derived from expert report evaluation norms and guidelines, augmented with task-specific expert guidance. This allows DREX to avoid designing a new rubric from scratch for each task while still applying a dense set of rubric items to each report, and to interpret system performance along standard dimensions and rubric items across tasks.
Table 8: Comparison of Information Verification Pipelines.
Benchmark Explicit Verif. Implicit Verif. Global Context Key Limitation vs. DREX
DeepResearch Bench
deepbench
Ignores uncited sentences; fails to detect hallucinations in transitional logic.
DeepResearchGym
deepgym
Limited to explicitly cited claims; only calculates citation precision/recall metrics.
ResearcherBench
ResearcherBench
Treats uncited claims simply as “ungrounded” (empty URL) without attempting context-based verification.
ReportBench
ReportBench
$\triangle$ Verifies non-cited claims via external web search voting, ignoring internal document grounding.
DeepScholar-Bench
DeepScholar-Bench
$\triangle$ Relies on physical distance (sliding window $w$); misses long-range semantic dependencies.
DeepResearch Arena
DeepResearchArena
$\triangle$ Evaluates Source $\to$ Report coverage (summarization) rather than Report $\to$ Source verification; misses uncited hallucinations.
LiveDRBench
characterizing
Relies on matching against static Ground Truth claims; lacks dynamic verification of citations against web sources.
SPOT
SPOT
Evaluates internal error detection against static ground truth; penalizes valid but unannotated error predictions (false positives).
[…]
Appendix B Data Construction Details
B.1 Topic Domain Analysis
To configure Deep Research tasks (queries), we first analyzed 5045 in-house user queries collected from Deep Research services to ensure a balanced distribution of tasks across domains. The domain distribution of these queries is presented in Figure 4, from which we finalized 8 topic domains, including the top 7 areas and an "Others" category that groups the rest.
B.2 HLE Subject Mapping
Although actual Deep Research service logs are collected in a research-oriented usage context unlike general QA systems, since the majority of users are not limited to experts in specific academic fields, individual queries are not designed with the premise of generating expert-level academic reports or papers immediately. In addition, they include partial queries that depend on the ongoing session context or practical/pragmatic information needs, so using them as-is for evaluation research tasks may lead to a mismatch with the high-difficulty research tasks we aim for in scope and format.
Accordingly, in this study, while maintaining the 8 topic domain composition derived from actual Deep Research logs, we defined subjects from Humanity’s Last Exam (HLE) hle that semantically correspond to each domain and used HLE items belonging to those subjects as seed tasks for each domain. At this time, only items within the difficulty range currently processable by LLMs among all HLE items were selected through a preliminary performance evaluation, and the selected items were remapped to the previously defined Deep Research topic domains to obtain a higher-difficulty academic task composition while preserving the actual usage distribution. The correspondence between each topic domain of the actual Deep Research service and HLE subjects is summarized in Table 9.
B.3 Conversion from HLE QA to DeepResearch Reports
Since the selected HLE items are initially configured in QA format, they are not suitable for deep report generation tasks as they stand. In this study, experts reviewed the question-answer-rationale of each item to identify the core concepts, theories, and phenomena covered, and then reconstructed them into research-oriented task queries suitable for the Deep Research usage context. At this time, deterministic elements such as specific answers, factual conclusions, or proofs presented in HLE were removed from the prompt, and instead, only high-level writing directions such as analysis scope, comparative perspectives, and issues to be discussed were presented to induce the model to develop the logical structure and conclusion on its own.
In addition, if necessary, the basic structure of the output, such as "Introduction–Related Work–Method–Analysis–Discussion," was briefly specified to make the task more similar to the long-form research report generation expected in actual Deep Research use cases. Through this conversion process, simple answer-type QA items were transformed into a report-generation task set with HLE-level high-quality and high-difficulty problem structures, while following the actual Deep Research usage domain distribution.
Figure 4: Topic domains extracted from real-world Deep Research service logs
In-house Domain Humanity’s Last Exam Subject
Finance & Business Economics
Software Development Computer Science, AI
Science & Technology Mathematics, Physics, Chemistry
Industrial / Hardware Engineering
Education & Jobs Education
Health Biology, Psychology
Others History, Linguistics, Philosophy
Table 9: Mapping between in-house topic domains and Humanity’s Last Exam subjects. Combines "Software Development" and "Software" categories.
Example Research Task (MDP / Value Iteration) Write a research report that explores the conditions under which value iteration in Markov Decision Processes converges geometrically to the optimal value function. Your report should provide precise definitions of the key concepts involved, examine the roles of rewards and discounting in ensuring convergence, and analyze both theoretical and practical factors that influence the algorithm’s behavior. Discuss circumstances where convergence occurs, where it may fail, and highlight open questions or limitations. The report should be structured as a rigorous technical investigation, integrating mathematical reasoning, conceptual explanations, and illustrative examples to support the analysis.
Appendix C Deep Research Report Evaluation Taxonomy Details
C.1 Dimension and Criteria Specification
In this section, we detail the 7 dimensions and 26 detailed criteria of the Deep Research Report Evaluation Taxonomy presented in the main text.
Request Completeness
This dimension assesses how systematically and comprehensively the report addresses the user query with respect to the required elements, analysis scope, and final key message. Request Completeness assesses whether all requirements presented in the query are included without omission and developed with sufficient depth. Scope Boundary assesses whether the covered content, excluded items, assumptions, and limitations are clearly defined and consistently maintained. Value Validity assesses whether the final key message encompasses the entire query, is sufficiently specific and practical, and is logically consistent with the evidence and analysis results presented in the report body.
Evidence Validity
This dimension evaluates how accurate and valid the report’s figures and arguments are in terms of calculation, methodology, and logical development. Numeric Accuracy examines whether calculation processes, used formulas/statistical models, and indicators/units are presented without error, are appropriate for the problem context, and are expressed transparently enough for a third party to reproduce and verify. Logical Support evaluates whether the argument is developed consistently with the topic, necessary background/assumptions/inference steps are specified, and major claims and counter-argument responses are persuasively supported without leaps based on facts, data, and interpretations.
Structure Consistency
This dimension evaluates whether the introduction, body, and conclusion, and the structure of each section of the report, are organized consistently with the topic and scope. Introduction checks whether it concisely and clearly presents the topic, problem, significance, and basic scope. The body checks whether the step-by-step argument is developed without omission or deviation, in accordance with the structure and scope presented in the introduction. Conclusion checks whether it synthesizes the body content to complete the purpose of the introduction without introducing new claims or evidence. Section-Level checks whether each section supports the overall structure through clear organizational principles and appropriate connections.
Narration Style
This dimension assesses whether the report faithfully follows the required format and style and is expressed in a way that the reader can read and understand the content without difficulty. The report form checks whether external requirements, such as document length and section system, conform to professional report practices. Writing Quality checks whether sentences are concise and accurate, and whether terminology and tone are consistent. Paragraph checks whether paragraphs are sufficiently developed around a single point and naturally integrated with structural auxiliary elements. Reader Friendliness assesses whether subheadings and simple explanations/examples are used to guide the reader in following complex content.
Ethics Compliance
This dimension assesses whether the report is written ethically and responsibly regarding sensitive issues, potential harm, and perspective balance. Sensitive Issues checks whether sensitive topics such as politics, race, and gender are addressed with neutral, fair language and a balanced perspective. Safety & Impact checks whether negative impacts, side effects, and potential misuse of proposals/technologies/research results are adequately reviewed, and whether specific method presentations are dangerous. Perspective Balance assesses whether the discussion is balanced by appropriately including related perspectives and opposing views without bias toward a specific position.
Information Sufficiency
This dimension assesses whether the information required to answer the request has been adequately secured and presented in terms of quantity and scope. Source Support evaluates whether verifiable evidence or reliable sources are provided to support claims that require external evidence. Information evaluates whether reliable facts and claims are sufficiently presented. Citations evaluate whether sufficient citations are made at necessary points. References are evaluated to determine whether the number of actually used references reaches an appropriate level.
Information Integrity
This dimension assesses the factuality of the external information used in the report and the reliability and diversity of the citations/sources supporting it. Claim Factuality measures the proportion of verifiable claims that are determined to be factual. Citation Validity measures the proportion of citations that actually support the claim among citations attached to each claim. Reference Accuracy measures the proportion of sources that correctly support the argument among the presented/used references. Reference Quality assesses whether the sources used are reproducible and reliable. Reference Diversity assesses whether the evidence maintains a sufficient level of source diversity without becoming overly concentrated on a few documents.
C.2 Construction Procedure and Evidence Mapping
To systematize and standardize universal and essential elements of expert report evaluation, this study synthesized and normalized evaluation standards across the natural sciences, engineering, and social sciences. To this end, we analyzed authoritative standards and guidelines widely used for writing and evaluating expert reports, such as guidelines for writing/reviewing academic papers and research reports, guidelines for systematic reviews/literature reviews, academic publication norms, and guidelines for writing/evaluating policy/regulatory/market analysis reports and consulting/advisory reports. We extracted the common required elements from these materials and integrated/generalized overlapping or domain-specific items.
To verify the validity of the drafted dimensions and detailed criteria, this study conducted cross-validation with 10 independent experts representing 6 domains (Computer Science, Artificial Intelligence, Biology, Chemistry, Mathematics, Psychology). Two experts per domain reviewed the suitability of each item, focusing on factors considered necessary in light of domain practices, such as whether each dimension and criterion is actually a meaningful and necessary evaluation standard in that domain, whether the definition and scope of application are excessively ambiguous, and whether it unnecessarily overlaps with other criteria. Based on this, they made binary (pass/fail) evaluations for each criterion. The final evaluation framework included only the criteria that passed this cross-validation process.
The following presents the mapping of which external standards were referenced for each evaluation dimension.
Appendix D Foundations of Deep Research Report Evaluation Taxonomy
Request Completeness
Request Completeness aligns with guideline-level completeness checks that evaluate whether all required components of a report are present according to prescribed inclusion mandates across authoritative standards Bastian2021PRISMA; Boutron2010CONSORT; ISO25010; WWC_Handbook; Percie2020ARRIVE; GRI_Standards; STROBE_2007. Scope Boundary reflects boundary-setting requirements that assess whether a report explicitly specifies its operative limits—such as eligibility conditions, assumptions, and constraints—within the structural fields defined in major evaluative frameworks Bastian2021PRISMA; NIST_AI_RMF; AHA2024Standards; AEA2024Data; OSF_Prereg; NASA_SE_Handbook; SEC_Reg_SK; SPIRIT_2013. Value Validity corresponds to coherence and utility standards that evaluate whether the report’s core conclusions not only align logically with the evidence base but also deliver sufficient specificity and practical feasibility to comprehensively address the user’s inquiry Guyatt2013GRADE; APA_JARS2025; NBER2024; IES_SEER; OECD_Policy_Eval.
Evidence Validity
Numeric Accuracy reflects evaluative requirements that check whether quantitative statements in a report correspond to verifiable computations, prespecified analytic procedures, and reproducible numerical evidence as established in major methodological frameworks Guyatt2013GRADE; JCGM_GUM; ACM_Badging; IUPAC_Green; IUPAP_Red; Econometric_Soc; AERA_APA_NCME_Standards; CODATA_2022; OECD_Quality; IFRS_Standards; ISO30414; IPCC_2019. Logical Support aligns with reasoning-assessment criteria that evaluate whether inferential steps are explicitly grounded in evidence, free of unsupported leaps, and consistent with theoretical or mathematical rigor EMS_Code2025; APA_GoodPractices; AERA2006Standards; AMS_Handbook; Fed_Rules_Evidence; COREQ_2007.
Structure Consistency
Introduction corresponds to structural-orientation requirements that assess whether a report’s opening section presents its scope and analytic pathway in accordance with prescribed organizational fields in established reporting and specification frameworks Bastian2021PRISMA; Boutron2010CONSORT; ISO29148_2018; IFRS_Integrated; CARE_2013. Body aligns with structural-progression criteria that evaluate whether the main sections develop the promised analytical sequence without omission or drift relative to the ordered components defined in recognized guideline structures Bastian2021PRISMA; Boutron2010CONSORT; ISO29148_2018; CF_Conventions; CARE_2013. Conclusion reflects closure-consistency requirements that examine whether final statements integrate preceding evidence without introducing unsupported expansions, consistent with the conclusion-governance rules embedded in major evaluative standards Bastian2021PRISMA; Boutron2010CONSORT; CARE_2013. Section-Level corresponds to intra- and inter-section coherence checks that assess alignment, ordering, and non-duplication of content in accordance with structured specification and reporting templates in authoritative frameworks Bastian2021PRISMA; ISO29148_2018; IFRS_Integrated.
Narration & Style
Report Form corresponds to format-governance requirements that evaluate whether a report adheres to prescribed structural conventions for scientific communication as codified in established editorial and reporting standards Bastian2021PRISMA; Boutron2010CONSORT; ICMJE2025; Leipzig_Rules; SIAM_Style; IPA_Handbook; APSA_Style; ASA_Style. Writing Quality aligns with language-precision criteria that assess clarity, specificity, and terminological consistency according to recognized guidelines for accurate and unbiased scholarly expression Boutron2010CONSORT; ICMJE2025; Chicago_Style; Blacks_Law; HBR_Guide. Paragraph reflects cohesion-assessment rules that examine whether individual paragraphs follow a coherent internal logic—anchoring topic statements, supporting evidence, and transitional structure—in line with authoritative reporting and specification frameworks Bastian2021PRISMA; ISO29148_2018; The_Bluebook; ASA_Style. Reader Friendliness corresponds to accessibility-oriented requirements that evaluate whether the narrative facilitates comprehension through appropriate signaling, explanatory devices, and presentation practices as outlined in major editorial and reporting guidelines Bastian2021PRISMA; ICMJE2025; HBR_Guide; GRI_Standards.
Ethics & Compliance
Sensitive Issues corresponds to ethical-screening requirements that evaluate whether sensitive domains are addressed with neutrality, respect, and responsible language in accordance with established editorial and reporting ethics standards ICMJE2025; Bastian2021PRISMA; OHA_Principles; OAH_Standards; BPS_Code; Belmont_Report; APSA_Ethics. Safety & Impact aligns with harm-assessment provisions that examine whether potential adverse effects, misuse risks, or disproportionate impacts are identified and mitigated under recognized guidelines for responsible scientific communication ICMJE2025; Guyatt2013GRADE; NRC_Prudent; Helsinki_Decl; ESA_Code. Perspective Balance reflects fairness-evaluation criteria that assess whether multiple viewpoints and counterpositions are represented without undue bias following principles of impartiality articulated in authoritative publication and reporting frameworks ICMJE2025; Bastian2021PRISMA; BPA_GoodPractice; AAP_Code; CFA_Code; ABA_Model_Rules; AAPOR_Code.
Information Sufficiency
Source Support aligns with evidence-provision requirements that evaluate whether claims are accompanied by verifiable supporting sources in accordance with established evidentiary and reporting guidelines Bastian2021PRISMA; FORCE11_Data; NBER2024; ISSCR2025; GIPS_2020. Information corresponds to content-adequacy criteria that assess whether a report supplies all necessary factual and contextual material needed to justify its reasoning under recognized completeness and transparency standards Bastian2021PRISMA; ISO25010; EQUATOR_Network; LSA_Ethics2024; JOC_Guidelines; USGS_FSP; GRI_Standards. Citations align with source-attribution rules that examine whether in-text citations are provided at appropriate argumentative locations following authoritative norms for evidentiary traceability ICMJE2025; Wilkinson2016FAIR; IEEE2025Ethics; ACS2021Ethical; The_Bluebook. References are evaluated to determine whether the number of actually used references is appropriate and whether they collectively form a well-documented, traceable, and accessible evidence base StanfordAI2025; FORCE11_Data; DA_RT_Principles.
Information Integrity
Claim Factuality reflects evidence-verification provisions that assess whether factual assertions in a report correspond to verifiable sources and documented evidence traces within established evaluative frameworks Wilkinson2016FAIR; APS2023Conduct; NIST_AI_RMF; PDG_Review; NIST_DLMF; SEC_Reg_SK. Citation Validity corresponds to source-justification checks that evaluate whether cited references substantively support the claims they accompany according to recognized standards for evidential accountability Wilkinson2016FAIR; ASA_Ethics; AHA2024Standards; Austin_Principles. Reference Accuracy aligns with source-quality and provenance requirements that examine whether referenced materials reliably and transparently support the arguments for which they are cited, in line with authoritative data-governance and reporting criteria Wilkinson2016FAIR; NBER2024; FORCE11_Data; APA_JARS2025; IFRS_Standards. Reference Quality evaluates whether the set of sources consists of reproducible, trustworthy, and methodologically sound information sources, ensuring that the report’s evidentiary foundation is grounded in reliable materials ACS2021Ethical; IEEE2025Ethics; SEP_Policy; Helsinki_Decl; Belmont_Report. Reference Diversity evaluates whether the evidence base avoids excessive concentration on a small number of sources and instead maintains a balanced level of source diversity, reducing the risk of narrow, biased, or selectively framed arguments EQUATOR_Network; LSA_Ethics2024; AAPOR_Code.
To establish a robust and universally applicable Deep Research Report Evaluation Taxonomy, this framework comprehensively integrates authoritative standards and guidelines from a wide spectrum of disciplines, ranging from quantitative sciences and engineering to the humanities, social sciences, and professional fields such as law and finance (see Table 10). By synthesizing the core principles shared across these diverse domains, this taxonomy prioritizes criteria that are universally recognized as essential for high-quality intellectual work. Consequently, the resulting evaluation model ensures that the generated reports not only adhere to domain-specific best practices but also satisfy the fundamental requirements of validity, clarity, and professional integrity demanded by the global research and practitioner communities.
Domain (80 listings) Reference
Artificial Intelligence (8) NIST_AI_RMF; StanfordAI2025; ACL2024Guidelines; NeurIPS2025Guidelines; ICML2025Instructions; ICLR2024Guidelines; JMLR2024AuthorGuide; Springer2025MLJournal
Biology (4) Bastian2021PRISMA; Boutron2010CONSORT; ISSCR2025; Percie2020ARRIVE
Business (4) GRI_Standards; HBR_Guide; ISO30414; IFRS_Integrated
Chemistry (4) ACS2021Ethical; IUPAC_Green; JOC_Guidelines; NRC_Prudent
Computer Science (4) ISO25010; ACM_Badging; ISO29148_2018; Wilkinson2016FAIR
Earth & Env. Science (4) IPCC_2019; USGS_FSP; CF_Conventions; ESA_Code
Economics (4) NBER2024; AEA2024Data; Econometric_Soc; OECD_Quality
Education (4) WWC_Handbook; AERA2006Standards; IES_SEER; AERA_APA_NCME_Standards
Engineering (4) IEEE_Std2025; IEEE2025Ethics; NASA_SE_Handbook; ISO29148_2018
Finance (4) IFRS_Standards; CFA_Code; GIPS_2020; SEC_Reg_SK
History (4) AHA2024Standards; Chicago_Style; OHA_Principles; OAH_Standards
Law (4) The_Bluebook; Blacks_Law; Fed_Rules_Evidence; ABA_Model_Rules
Linguistics (4) LSA_Ethics2024; Leipzig_Rules; Austin_Principles; IPA_Handbook
Mathematics (4) EMS_Code2025; AMS_Handbook; SIAM_Style; NIST_DLMF
Medicine (4) Helsinki_Decl; STROBE_2007; CARE_2013; SPIRIT_2013
Philosophy (4) APA_GoodPractices; SEP_Policy; BPA_GoodPractice; AAP_Code
Physics (4) APS2023Conduct; IUPAP_Red; PDG_Review; CODATA_2022
Political Science (4) DA_RT_Principles; APSA_Style; OECD_Policy_Eval; APSA_Ethics
Psychology (4) APA_JARS2025; OSF_Prereg; BPS_Code; AERA_APA_NCME_Standards
Sociology (4) AAPOR_Code; COREQ_2007; Belmont_Report; ASA_Style
Table 10: References by Domain used in Deep Research Report Evaluation Taxonomy. (Total 80 listings from 78 unique sources; 2 interdisciplinary standards are cross-listed).
Appendix E Rubric Structure
In this appendix, we summarize the hierarchical structure and scoring method of the rubric used to evaluate the Deep Research report.
Dimension / Criterion Element / Rubric description
2. Evidence Validity └ 2.2 Logical Support 2.2.5 Element All major claims follow logically from the previously presented facts, data, interpretations, and assumptions, without skipped steps or unsupported leaps. [2.2.5.1 Coverage] All claims that require logical support are explicitly linked to the relevant facts, data, and interpretations (including the key evidence specified in the Expert Evaluation Guidance, EG), and no central claim is left without the necessary supporting evidence. [2.2.5.2 Quality] Each major claim is explained through clear, well-structured reasoning that shows how the underlying facts, data, and interpretations lead to that conclusion, making the logical connection explicit and easy for an expert reader to follow.
Table 11: Example Coverage/Quality rubric factors for a reasoning-focused element under 2. Evidence Validity / 2.2 Logical Support.
E.1 Hierarchical Levels
We have the following 4-level hierarchical structure to systematically evaluate report quality in multiple dimensions.
Level 1: Evaluation Dimensions (7)
The 7 upper dimensions presented in Table 1 represent different evaluation perspectives of report quality.
Level 2: Detailed Criteria (26)
Detailed criteria that decompose each dimension into finer evaluation units; 2–5 detailed criteria are defined under each dimension. This forms an intermediate-level unit that specifies which aspects to examine for each dimension.
Level 3: Evaluation Elements (40)
Evaluation elements that concretize each detailed criterion into units actually observable in the report. One or more elements are mapped under one detailed criterion, and are defined as evaluation units directly corresponding to report content, such as "Inclusion of requested items," "Specification of scope/limitations," and "Specificity of final message."
Level 4: Rubric Factors (130)
Rubric factors corresponding to the fourth level of the rubric are defined to evaluate each evaluation element (Level 3) from two perspectives. One element is subdivided into multiple factors from the perspectives of Coverage (C) and Quality (Q), and these factors serve as the minimum unit for actual scoring.
Thus, with 7 dimensions, 26 detailed criteria, and 40 elements as the framework, each element is divided into Coverage/Quality factors to constitute a total of 130 specific check items. Table 11 shows an example of one element and its corresponding Coverage/Quality factors in the 2. Evidence Validity / 2.2 Logical Support dimension.
E.2 Coverage and Quality Factors
The rubric factors corresponding to Level 4 of the rubric evaluate each evaluation element (Level 3) from two perspectives: Coverage(C) and Quality(Q). The Coverage factor examines whether the element has been covered without omission as required throughout the document, i.e., how faithfully it includes content scattered in multiple locations or composed of multiple detailed requirements. On the other hand, the Quality factor evaluates whether the description possesses sufficient depth, logic, and rigor based on professional report standards, assuming the parts already written, i.e., "how well it is written." By separating Coverage and Quality in this way, we can independently measure (1) what is missing and (2) what the level of the included content is in a long report. Both factors use a 1–10 point scale, and detailed criteria for scoring are presented in Table 12.
Score Range Coverage Quality
9–10 (Perfect) Fully meets all requirements; No omissions; No revision needed Excellent quality in all relevant aspects; No revision needed — Top-tier international journal level, or high-end professional report meeting or exceeding standards in specific technical/industrial contexts
7–8 (Excellent) Meets almost all requirements; Only 1–2 minor omissions, minimal impact High quality; Meets most academic/professional standards, only minor improvements possible — Solid peer-reviewed journal, excellent doctoral research, high-quality industry report level
5–6 (Good) Meets more than half; Meets most key requirements, minor elements missing Meets essential professional standards; Clear structure and competent analysis but room for improvement — Well-written master’s thesis or standard professional report level
3–4 (Inadequate) Partially meets; Several key omissions Noticeable flaws in several aspects; Significant revision needed — Undergraduate thesis or entry-level professional report level
1–2 (Poor) Most requirements are missing or treated only superficially Fails to meet basic professional standards; Lacks depth, rigor, accuracy — Below undergraduate level; Unsuitable for publication or professional use
Table 12: Interpretation of 1–10 score ranges for Coverage (C) and Quality (Q) factors.
E.3 Scoring and Aggregation
The score aggregation flow in this evaluation framework consists of factor $\rightarrow$ element $\rightarrow$ criterion $\rightarrow$ dimension. First, rubric factors (the lowest items) are evaluated with a 1–10 score (or N/A). Coverage (C) and Quality (Q) are calculated separately up to the factor/element level, then integrated at the element level, and the resulting element score is raised to the criterion and dimension levels. N/A items are excluded from all average calculations.
Let $r$ be the report, and type $T\in\{C,Q\}$ represent Coverage and Quality, respectively. Let the score of each rubric factor $f$ be $s_{r,f}\in\{1,\dots,10\}$ (or N/A), and the set of non-N/A factors of type $T$ belonging to element $e$ be $F_{e,r}^{T}$. Let $E_{c,r}$ be the set of elements belonging to criterion $c$, and $\mathcal{C}_{d,r}$ be the set of criteria belonging to dimension $d$. Then, the factor $\rightarrow$ element $\rightarrow$ criterion $\rightarrow$ dimension aggregation is defined as follows:
The Coverage/Quality score (element level) of report $r$ for element $e$ is
$S_{r}^{T}(e)=\frac{1}{\lvert F_{e,r}^{T}\rvert}\sum_{f\in F_{e,r}^{T}}s_{r,f},\quad T\in\{C,Q\},$ (3)
and if $F_{e,r}^{T}=\varnothing$, $S_{r}^{T}(e)$ is set to N/A.
The integrated element score $S_{r}(e)$ of element $e$ is set as the average of defined values among C/Q. Let the set of types defined in element $e$ be $\mathcal{T}_{e}=\{T\in\{C,Q\}\mid S_{r}^{T}(e)\neq\text{N/A}\}$.
$S_{r}(e)=\frac{1}{\lvert\mathcal{T}_{e}\rvert}\sum_{T\in\mathcal{T}_{e}}S_{r}^{T}(e),$ (4)
If $\mathcal{T}_{e}=\varnothing$, $S_{r}(e)$ is set to N/A.
The score of criterion $c$ is defined as the average of element scores belonging to that criterion.
$S_{r}(c)=\frac{1}{\lvert E_{c,r}\rvert}\sum_{e\in E_{c,r}}S_{r}(e).$ (5)
If $E_{c,r}=\varnothing$, $S_{r}(c)$ is set to N/A.
The score of dimension $d$ is defined as the average of criterion scores belonging to that dimension.
$S_{r}(d)=\frac{1}{\lvert\mathcal{C}_{d,r}\rvert}\sum_{c\in\mathcal{C}_{d,r}}S_{r}(c).$ (6)
If $\mathcal{C}_{d,r}=\varnothing$, $S_{r}(d)$ is set to N/A.
In summary, we calculate $S_{r}^{C}(e),S_{r}^{Q}(e)$ by averaging C/Q factors in each element, integrate them to create the element score $S_{r}(e)$, then average the element scores by criterion to calculate $S_{r}(c)$, and finally average the criterion scores to calculate the dimension score $S_{r}(d)$.
Appendix F Human-based Information Verification Protocol
F.1 Overview
DREX’s Information Verification Protocol is based on a stepwise information verification procedure performed by human evaluators. This section details the actual two-step procedure (Claim Extraction, Factual Accuracy Evaluation) performed by human evaluators.
F.2 Step 1: Claim Extraction and Classification
Human procedure.
Evaluators segmented the report into sentences (in ‘Lx.Sy‘ format), reviewed each sentence individually to identify sentences containing claims, and extracted only the core claims. Pronouns were replaced with explicit references according to the context, and if a single sentence contained multiple claims, they were separated. All claims were classified into 6 types (A–F): Explicit Citation (A), Implicit – Same Section (B), Implicit – Previous Section (C), Structural Recap (D), No Citation Required (E), and Unknown Source (F). For A–C type claims, the corresponding citation or evidence position was recorded together. This process was performed on 2 randomly selected reports (total 728 claims), thereby constructing a Ground Truth for evaluating the recall of the extraction model.
Table 13 shows the definitions of the 6 claim types, and Table 14 shows examples of extracted sentences and classification results.
Table 13: Information Verification ModuleA–F Claim Type Definitions and Examples. Types A–C are targets for external evidence verification, while Types D–E–F are classified as internal information or unverifiable areas.
Type Definition and Example
A Cited Claim: A claim explicitly including a citation marker within the sentence. Example: “Multi-junction solar cells achieve efficiencies above 45% [1].”
B Uncited – Same Section / Paragraph: When the evidence citation exists in a previous sentence within the same section (or paragraph). Example: “This efficiency improvement is due to the layered structure.” (Evidence → L2.S1)
C Uncited – Previous Section / Paragraph: When the citation evidence exists in a previous section or paragraph. Example: “These findings confirm the results of earlier solar-cell studies.” (Evidence → L1.S3)
D Uncited – Structural Recap: A claim corresponding to a restatement of content in the document structure, such as introduction, conclusion, or summary. Example: “In conclusion, this paper reviewed recent advances in solar technology.”
E Uncited – No Citation Required: The author’s direct results, general knowledge, or a claim not requiring citation. Example: “Photosynthesis converts light energy into chemical energy.”
F No Citation – Unknown Source: A claim requiring external evidence but for which no source is presented. Example: “These panels can last for 50 years without degradation.”
Table 14: Example of human-annotated claims and corresponding LLM classification results.
Pos. Sentence / Extracted Claim Class (label)
L1.S3 The development of multi-junction solar cells has achieved efficiencies above 45% in laboratories [1]. A (Explicit citation)
L2.S1 This dramatic increase is due to the layering of different semiconductor materials. B (Implicit citation → linked to L1.S3)
L2.S3 The enhanced efficiency of these cells will reduce the land area required for solar farms. C (Implicit cross-section)
L2.S4 These new panels are durable enough to withstand a Category 4 hurricane. F (No known source)
F.3 Step 2: Claim Verification
Human procedure.
For 100 claims randomly selected from types A–C, two human evaluators independently assessed their factuality. The evaluation followed a single-criterion protocol determining whether the cited source explicitly supports (Supported) or lacks information/is irrelevant to (Not Supported) the content of the claim.
Annotator Qualifications & Adjudication.
The evaluators consisted of 2 individuals holding a master’s degree or equivalent experience in the report’s domain. The two evaluators made judgments independently, and for items where disagreement occurred, the final label unanimously agreed upon through discussion was established as the Ground Truth. Through this process, personal bias was excluded, and the objectivity of the evaluation was secured. The human verification results for these 100 claims were used as Ground Truth for the model performance evaluation in Section 6.5.
Detailed Verification Rubric.
Fact verification was strictly performed according to the following detailed criteria:
  • Supported: When the cited document clearly and directly includes the core facts (figures, causality, definitions, etc.) of the claim. The implied meaning in context must match the intent of the claim.
  • Not Supported: When the basis for the claim cannot be found in the cited document, or the document is irrelevant to the topic. Cases where the document link is invalid (404) or the content cannot be verified are also included.
Inter-human Agreement.
To verify the reliability of this protocol, we measured the agreement (Cohen’s Kappa) between the two evaluators. The analysis result showed that a high level of agreement (Substantial Agreement) of $\kappa=0.80$ was achieved in the Claim Support judgment. This suggests that the proposed verification criteria are objective and reproducible.
Appendix G LLM-based Information Verification Implementation
G.1 Overview
The Information Verification Moduleis designed to automate the human verification protocol described above. The LLM analyzes the report sentence by sentence to extract and classify claims, and if necessary, retrieves external documents to verify their factuality. In this process, algorithms such as Batch Extraction, Back-tracking, and Relevant Context Filtering were applied to achieve both cost efficiency and accuracy.
G.2 Claim Extraction and Classification
LLM adaptation.
The Information Verification Moduleautomates the procedure of human evaluators. After tokenizing the report into sentences, the LLM evaluates the "claim likelihood" of each sentence and predicts A–F types based on the 6-Type definitions defined in Table 13. For Type B and C claims, evidence sentences are subsequently searched using Back-tracking.
Batch Extraction Strategy
To mitigate the "Lost-in-the-Middle" phenomenon that occurs when processing long contexts and to increase cost efficiency, this study introduced a Batch Extraction strategy. After dividing the entire report $D$ into sentence units $S=\{s_{1},s_{2},\dots,s_{m}\}$, they are processed in batches of a fixed size $B$ (in this study, $B=20$). Each batch processing provides the full report context ($D$) at the beginning of the prompt, but instructs the model to extract claims only for the sentences belonging to the current batch ($s_{i},\dots,s_{i+B-1}$).
Batch Extraction Prompt Structure System: You are an expert fact-checker and claim extractor. User: # Full Report Context {Full_Report_Text} # Target Sentences to Extract Claims From L10.S1: … L10.S20: … Extract claims only from the target sentences above. Use the full report context for coreference resolution.
Example JSON Output
 "claims": [{
  "position": "L10.S1",
  "index": 1,
  "claim_text": "Solar efficiency reached
  45% [1].",
  "claim_class": "A",
  "direct_citation": "[1]",
  "evidence_position": null
This method induces the model to focus on local sentences while remaining aware of the entire context, achieving human-level claim-extraction performance.
Classification Performance.
The results of performing type classification on 728 claims using GPT-5-mini are shown in Table 15. Explicit Citation (A) recorded a very high F1 score of 0.95, being identified almost perfectly. The distinction between Implicit Citation (BC) and Internal Information (DE) was somewhat challenging, but it recorded meaningful classification performance with F1 scores of 0.52 and 0.53, respectively. In particular, for the Unknown Source (F) type, which is the most difficult to identify, it achieved a precision of 0.35 and a recall of 0.25, indicating the ability to detect potential hallucination sentences. The overall accuracy was 0.68, which is a decent level considering the context of complex academic reports.
Table 15: Claim Classification Performance (GPT-5-mini). Classes are merged into A (Explicit), BC (Implicit), DE (Internal/Structural), and F (Unknown).
Class Precision Recall F1-Score Support
A (Explicit) 0.92 1.00 0.95 272
BC (Implicit) 0.42 0.68 0.52 163
DE (Internal) 0.76 0.41 0.53 242
F (Unknown) 0.35 0.25 0.30 51
Overall 0.68 0.68 0.67 728
Comparison with Sliding Window Baseline.
To validate the effectiveness of the LLM’s targeted evidence prediction, we compared it with a “Sliding Window” baseline on the subset of correctly classified B/C claims ($N=131$). The sliding window method collects all citations within a window of size $k$ centered on the claim. As shown in Table 16, the proposed LLM method achieves the highest Jaccard Index (0.7070) and Precision (0.7109), outperforming the sliding window baselines ($k=5,10,15$). While increasing the window size ($k$) improves Recall (up to 0.93), it significantly degrades Precision and Jaccard due to the inclusion of irrelevant citations. This result demonstrates that the LLM’s “Evidence Position” prediction provides a precise pointer to the supporting evidence, which is crucial for efficient verification.
Table 16: Baseline comparison on correctly predicted B/C claims ($N=131$). The GPT-5-mini method achieves the best balance of precision and accuracy (Jaccard).
Method Jaccard Precision Recall
Backtracking (Ours) 0.7070 0.7109 0.7383
Sliding ($k=5$) 0.6822 0.6822 0.8022
Sliding ($k=10$) 0.6247 0.6247 0.8791
Sliding ($k=15$) 0.5627 0.5627 0.9341
G.3 Claim Verification
Semantic Back-tracking for Citation Recovery
The Information Verification Moduleuses a Backtracking algorithm to find evidence for claims without explicit citations (Types B and C). Type B (Same Section) and Type C (Previous Section) claims often share citations from previous sentences in context. The algorithm traces the ‘evidence_position‘ (location ID of the reference target sentence, e.g., L1.S3) recorded in the claim’s metadata and adds the citations held by that target sentence to the citations of the current claim. This restores omitted citation relationships and allows the corresponding source to be reviewed together in the subsequent verification step.
Formally, for a set of claims $\mathcal{C}=\{c_{1},c_{2},\dots,c_{n}\}$, each claim $c_{i}$ has a position $p_{i}$, type $t_{i}$, explicit citations $R_{i}$, and a reference position $ref_{i}$ (if $t_{i}\in\{B,C\}$). The Back-tracking function $f_{backtrack}(c_{i})$ is defined as follows:
$f_{backtrack}(c_{i})=\begin{cases}R_{j}&\text{if }t_{i}\in\{B,C\}\text{ and }\exists c_{j}\text{ s.t. }p_{j}=ref_{i}\\ \emptyset&\text{otherwise}\end{cases}$
Finally, the citation set used for verification becomes $R^{\prime}_{i}=R_{i}\cup f_{backtrack}(c_{i})$.
Relevant Context Filtering (Retrieval-Augmented Verification)
Using the entire report or long retrieved documents as input in the verification step is costly and can induce hallucinations due to unnecessary information (Noise). To solve this, we apply Relevant Context Filtering. For the verification target claim $q$ and the retrieved document $D_{retrieved}$, the document is divided into chunks $K=\{k_{1},k_{2},\dots,k_{r}\}$ (Size $\approx$ 1000 tokens). Then, the relevance score $Sim(q,k_{j})$ between each chunk and the claim is calculated, and only the top $N$ (Top-K) chunks are selected and used as input for the verification model.
In this study, the BM25 robertson2009probabilistic algorithm was used as the default Selector: The selected chunk set $K_{selected}=\{k\in K\mid rank(Sim(q,k))\leq N\}$ is combined while maintaining the original document order to form the final context $C_{final}$.
$C_{final}=\text{Concatenate}(\text{SortByPosition}(K_{selected}))$
Through this process, token costs can be reduced by more than 80% while maintaining or improving verification accuracy.
LLM Verification Logic.
The Information Verification Module’s automatic verifier is prompt-engineered to determine whether the given context supports the claim, identical to the human protocol above. The LLM infers whether the claim’s core content and numerical information match the source to make a final judgment.
Example.
Claim: “Multi-junction solar cells achieve efficiencies above 45% in lab settings [1].”
Reference [1]: Reports a 46.2% lab efficiency under concentrated light.
Evaluation: The reference explicitly states 46.2% efficiency, supporting the claim of "above 45%".
$\Rightarrow$ Final Result: Supported
Fair Use and Direct Quote Detection
To ensure compliance with copyright and research ethics, the appropriateness of direct quotes is evaluated. First, the similarity between sentences in the report and the original text (Reference) is calculated using RapidFuzz and SequenceMatcher. If the similarity ($\tau$) is 0.9 or higher, it is considered a direct quote. Sentences detected as direct quotes are primarily checked for quotation marks (" ") or block quote formatting using regular expressions. If the format is insufficient or additional judgment is needed, the LLM is presented with the sentence and the original text to comprehensively evaluate compliance with "Fair Use Guidelines" (citation indication, minimal quotation, etc.). The final score is calculated by deducting points from a maximum of 10 based on the non-compliant ratio among all sentences.
G.4 Evaluation Metrics
The Information Verification Modulecalculates the following detailed metrics to evaluate Integrity and Sufficiency.
Integrity Metrics
  • External Claim Accuracy: The proportion of claims verified as factual among claims requiring external evidence (Type A, B, C).
[…]
  • Citation Accuracy: The proportion of citations that correctly support the corresponding claim among all citations.
    $\text{Cit. Acc}=\frac{|\text{Supported Citations}|}{|\text{Total Citations}|}$
  • Reference Accuracy (Supported per Shown): The proportion of references that actually contributed to content verification (Supported) among unique references shown in the report.
    $\text{Ref. Acc}=\frac{|\text{Supported Unique References}|}{|\text{Total Unique References Shown}|}$
  • Reference Reproducibility: The proportion of references that were accessible during the actual verification process and for which the webpage in markdown format could be successfully retrieved using the Jina API (not Error).
    $\text{Reproducibility}=1-\frac{|\text{Error References}|}{|\text{Used References}|}$
  • Information Reliability: The proportion of references that are both reliable sources and support the content among the used references.
    $\text{Reliability}=\frac{|\text{Reliable \& Supported References}|}{|\text{Used References}|}$
  • Reference Diversity (Citations CV): Uses the Coefficient of Variation of citation counts to measure dependency on specific references. (Lower is better; inverted for scoring)
    $\text{CV}=\frac{\sigma(\text{Citation Counts})}{\mu(\text{Citation Counts})}$
Sufficiency Metrics
  • Externally Verifiable Claims Ratio: The proportion of claims verifiable with external evidence (Type A, B, C) among all claims.
    $\text{Verifiable Ratio}=\frac{|\text{Claims (A, B, C)}|}{|\text{Total Claims}|}$
  • Information Quantity: The total number of claims verified as factual (Supported). (Raw Count before normalization)
    $\text{Info. Qty}=|\text{Accurate Verifiable Claims}|$
  • Citation Quantity: The total number of valid citations (Supported Citation) supporting claims.
    $\text{Cit. Qty}=|\text{Supported Citations}|$
  • Reference Quantity: The total number of valid references (Supported Reference) supporting claims.
    $\text{Ref. Qty}=|\text{Supported References}|$

### v2

@@
-DEER: A Comprehensive and Reliable Benchmark for Deep-Research Expert Reports
+DEER: A Benchmark for Evaluating Deep Research Agents on Expert Report Generation
 Abstract
-As large language models (LLMs) advance, deep research systems can generate expert-level reports via multi-step reasoning and evidence-based synthesis, but evaluating such reports remains challenging.
-Existing benchmarks often lack systematic criteria for expert reporting, evaluations that rely heavily on LLM judges can fail to capture issues that require expert judgment, and source verification typically covers only a limited subset of explicitly cited statements rather than report-wide factual reliability.
-We introduce DEER, a benchmark for evaluating expert-level deep research reports.
-DEER comprises 50 report-writing tasks spanning 13 domains and an expert-grounded evaluation taxonomy (7 dimensions, 25 sub-dimension) operationalized into 130 fine-grained rubric items.
-DEER further provides task-specific expert guidance to help LLM judges assess expert-level report quality more consistently.
-Complementing rubric-based assessment, we propose a document-level fact-checking architecture that extracts and verifies all claims across the entire report, including both cited and uncited ones, and quantifies external-evidence quality.
-DEER correlates closely with human expert judgments and yields interpretable diagnostics of system strengths and weaknesses.
-DEER: A Comprehensive and Reliable Benchmark
-for Deep-Research Expert Reports
+As large language models advance, deep research systems capable of generating expert-level reports through multi-step reasoning and evidence-based synthesis are emerging.
+However, evaluating such reports remains challenging.
+Existing benchmarks often lack systematic evaluation criteria, rely heavily on LLM-based judges that may miss issues requiring expert judgment, and verify only a limited subset of explicitly cited statements rather than report-wide factual reliability.
+To address these limitations, we introduce DEER, a benchmark for evaluating expert-level deep research reports.
+DEER comprises 50 report-writing tasks spanning 13 domains, along with an expert-grounded evaluation taxonomy with seven dimensions and 25 subdimensions, operationalized into 101 fine-grained rubric items.
+To improve evaluation consistency, DEER provides task-specific Expert Evaluation Guidance to support LLM-based judging.
+Complementing rubric-based assessment, we propose a document-level fact-checking architecture that verifies both cited and uncited claims and quantifies the quality and reliability of the supporting evidence.
+Experimental results show that DEER exhibits strong correlation with human expert judgments and yields interpretable diagnostics of system strengths and weaknesses.
+DEER: A Benchmark for Evaluating Deep Research Agents
+on Expert Report Generation
 Janghoon Han1, Heegyu Kim1, Changho Lee1, Dahm Lee1, Min Hyung Park1, Hosung Song1, Stanley Jungkyu Choi1, Moontae Lee1,2, Honglak Lee1,3 1LG AI Research, 2University of Illinois Chicago, 3University of Michigan, Ann Arbor Correspondence: hanjh9439@gmail.com
 1 Introduction
-Driven by the rapid advancement of large language models (LLMs), automated deep research systems are emerging as a core technology in both academia and industry openai2025deepresearch; google2025gemini; anthropic2025claude; qwen3technicalreport; webthinker; manusearch; webweaver.
-Unlike simple information retrieval, deep research systems tackle complex research queries by decomposing them into multiple steps and dynamically seeking additional information based on intermediate results.
-This enables them to integrate information from diverse sources and analyze different perspectives in a balanced manner, producing reliable, evidence-based research reports deepresearch_survey; deepresearch_survey2; characterizing.
-These advanced capabilities enable deep research systems to achieve strong performance even on challenging benchmark tasks.
-Figure 1: Deep Research System Performance Comparison. Shows the performance by type for 5 major models in the proposed benchmark.
-Early evaluation of deep research systems relied on complex reasoning benchmarks gpqa; hle, which indirectly assessed information-gathering, hypothesis testing, and multi-step reasoning through performance on these tasks.
-Subsequently, deep web search QA benchmarks were developed to directly assess systems’ web browsing and information retrieval abilities—core capabilities of deep research—by measuring multi-step search, information integration, and answer derivation browsecomp; frames; gaia; browsecomp_plus; mind2web2.
-More recently, deep research report benchmarks have emerged to evaluate the quality of generated reports from multiple perspectives, going beyond simple short-answer evaluation webthinker; deepbench; deepconsult; deepgym.
-Despite these advancements, current methodologies for evaluating deep research systems still face key limitations in assessing expert reports.
-First, while prior work has proposed high-level dimensions for report evaluation, these frameworks have yet to comprehensively specify the diverse elements required for expert-level reports.
-In particular, many frameworks deepbench; DeepScholar-Bench; LiveResearchBench; DeepResearchArena rely on LLMs—rather than human experts—to instantiate detailed sub-dimension and concrete evaluation items, which can lead to inconsistencies and reduced reliability.
-Second, even with a fixed rubric, evaluations that rely on LLM judges can fail to capture issues that require expert judgment.
-Third, current approaches to source verification are typically limited to sections with explicit citation markers, leaving report-wide claim reliability insufficiently examined.
+Figure 1: Deep Research System Performance Comparison. Results for five systems on the proposed benchmark.
+Driven by rapid advances in large language models (LLMs), automated deep research systems are emerging as a core technology in both academia and industry OpenAI (2025); Google (2025); Anthropic (2025); Yang et al.
+(2025); Li et al.
+(2025b); Huang et al.
+(2025); Li et al.
+(2025c).
+Unlike traditional information retrieval systems, these systems address complex research queries by decomposing them into multiple steps and dynamically seeking additional information based on intermediate results.
+Through this process, they integrate information from diverse sources and synthesize multiple perspectives to produce reliable, evidence-based research reports Xu and Peng (2025); Zhang et al.
+(2025); Java et al.
+(2025).
+As a result, deep research systems demonstrate strong performance even on challenging benchmark tasks.
+Early evaluations of deep research systems relied primarily on complex reasoning benchmarks Rein et al.
+(2023); Phan et al.
+(2025), which indirectly assessed information-gathering, hypothesis testing, and multi-step reasoning through task performance.
+Subsequently, deep web search QA benchmarks were introduced to more directly measure systems’ web browsing and information retrieval abilities—core capabilities of deep research—by evaluating multi-step search, information integration, and answer derivation Wei et al.
+(2025a); Krishna et al.
+(2025); Mialon et al.
+(2023); Chen et al.
+(2025); Gou et al.
+(2025).
+More recently, deep research report benchmarks have emerged to evaluate the quality of generated reports from multiple perspectives, moving beyond simple short-answer-based evaluation Li et al.
+(2025b); Du et al.
+(2025); Consult (2025); Coelho et al.
+(2025).
+Despite these advancements, existing methodologies for evaluating deep research systems continue to face important limitations when applied to expert-level reports.
+First, although prior work has proposed high-level dimensions for report evaluation, these frameworks do not comprehensively specify the diverse elements required for expert reporting.
+In particular, many approaches Du et al.
+(2025); Patel et al.
+(2025); Wang et al.
+(2025); Wan et al.
+(2025) rely on LLMs rather than human experts to define detailed subdimensions and concrete evaluation items, which can introduce inconsistencies and reduce evaluation reliability.
+Second, even when a fixed rubric is provided, evaluations that rely on LLM judges may fail to identify issues that require expert judgment.
+Third, current approaches to source verification are typically restricted to sections with explicit citation markers, leaving the factual reliability of claims across the full report insufficiently examined.
 Together, these limitations hinder comprehensive and reliable evaluation of deep research systems.
-To address these limitations, we propose the Deep rEsearch Expert Report benchmark (DEER), consisting of 50 report-generation tasks spanning 13 domains, for comprehensive and reliable evaluation.
-DEER analyzes a wide range of expert report assessment criteria and consolidates them through expert consensus into a Deep Research Report Evaluation Taxonomy with 7 dimensions and 25 sub-dimension.
-Based on this taxonomy, DEER evaluates each report along two complementary components: (i) report-quality dimensions that require holistic document-level judgment and (ii) external-information dimensions that can be assessed at the claim level via fact verification.
-For report quality, we operationalize the taxonomy into a fixed set of 130 fine-grained rubric items and supplement it with report-specific Expert Evaluation Guidance authored by domain experts to improve the consistency and validity of LLM-judge-based scoring.
-For external information, we introduce a fact-verification module that checks all cited and uncited claims across the report and produces quantitative measures of external-evidence quality.
-Finally, DEER combines these metric-based signals with rubric-based LLM evaluations to yield an integrated, multi-dimensional assessment of each expert report.
-Figure 1 presents an overview of dimension-level performance for major deep research systems evaluated on DEER.
-The main contributions of this study are as follows:
-  • We present DEER, a systematic and interpretable benchmark for deep research, built on a hierarchical evaluation taxonomy.
-  • We instantiate this taxonomy into a fixed set of 130 fine-grained rubric items and supplement it with report-specific Expert Evaluation Guidance, improving the consistency and reliability of LLM-judge-based evaluation.
-  • We propose a fact verification architecture that backtracks inter-claim dependency relations to retrieve and verify relevant citations, enabling more complete report-wide evaluation of all claims.
+To address these limitations, we propose DEER (the DEep research Expert Report benchmark), which provides comprehensive and reliable evaluation of deep research reports through 50 report-generation tasks spanning 13 domains.
+DEER analyzes a broad range of expert report assessment criteria and consolidates them via expert consensus into a Deep Research Report Evaluation Taxonomy comprising 7 dimensions and 25 subdimensions.
+Based on this taxonomy, DEER evaluates each report across two complementary components: (i) report-quality dimensions that require holistic, document-level judgment, and (ii) external-information dimensions that can be assessed at the claim level by fact verification.
+For report quality, we operationalize the taxonomy into a fixed set of 101 fine-grained rubric items and supplement these with report-specific Expert Evaluation Guidance authored by domain experts to improve the consistency and validity of LLM-based scoring.
+For external information, we introduce a fact-verification module that examines both cited and uncited claims across the report and produces quantitative measures of evidence quality.
+Finally, DEER integrates rubric-based LLM evaluations with the resulting metric-based signals to yield a unified, multi-dimensional assessment of each expert report.
+Figure 1 presents an overview of dimension-level performance for major deep research systems evaluated using DEER.
+The main contributions of this work are as follows:
+  • We present DEER, a systematic and interpretable benchmark for evaluating deep research reports, grounded in a hierarchical evaluation taxonomy.
+  • We operationalize this taxonomy into a fixed set of 101 fine-grained rubric items and provide report-specific Expert Evaluation Guidance, improving the consistency and reliability of LLM-based evaluation.
+  • We propose a report-level fact verification architecture that backtracks inter-claim dependencies to retrieve and verify relevant citations, enabling more complete evaluation of factual reliability across full reports.
+Figure 2: Overview of the DEER evaluation framework.
+(a) Research question and expert guidance generation from real-world deep research queries.
+(b) Construction of the Deep Research Evaluation Taxonomy consisting of 7 dimensions, 25 sub-dimensions, and 101 granular rubrics.
+(c) The DEER evaluation pipeline, integrating expert-guided LLM-as-a-judge scoring with claim extraction and information verification to assess deep research reports.
 2 Related Works
-2.1 LLM-based Evaluation Automation
-With the proliferation of Large Language Models, various approaches have been proposed to automate human-centric evaluation tasks using LLMs.
-Most notably, the LLM-as-a-Judge method, which directly uses LLMs as evaluators, has been widely researched.
-JudgeLM zhu2023judgelm and G-Eval liu2023g demonstrate that models can score generated responses based on multiple criteria such as accuracy, consistency, and logic.
-In addition, the scope of automated LLM evaluation is expanding to pairwise comparison-based evaluations such as Chatbot Arena chiang2024chatbot, CheckList-style robustness and bias checks checkeval, and various automated scoring frameworks.
-However, these approaches often struggle to control bias or hallucinations in the LLM’s own judgments and do not adequately address evaluations that require complex evidence structures, such as long-form reports.
-2.2 Deep Research and High-Difficulty QA Evaluation
-Recently, Deep Research benchmarks have emerged that go beyond simple knowledge recall to evaluate complex reasoning, web browsing, and information integration capabilities.
-High-difficulty QA benchmarks such as HLE hle, GPQA gpqa, and FRAMES frames reveal the limitations of existing models by requiring long-context reasoning, expert-level thinking, and information synthesis skills.
-Furthermore, ManuSearch manusearch and Mind2Web-2 mind2web2 employ multi-agent web browsing and agent-as-a-judge architectures to evaluate models’ ability to replicate real-world research processes.
-Additionally, DeepResearchBench deepbench and WebThinker Eval webthinker present Deep Research report-centric evaluations that assess literature-based report generation and multi-document reasoning, classifying them as measuring "research capability" beyond single-answer QA.
-However, while these benchmarks evaluate correctness, they cannot precisely verify the validity of evidence, citations, and arguments within the report, focusing more on whether the correct answer is reached than on factuality.
-2.3 Claim Verification
-In automated claim verification, various approaches have been proposed to mitigate LLM hallucinations.
-TrustGPT huang2023trustgpt attempted to improve factuality based on RAG, and InFact rothermel-etal-2024-infact established a structural verification system by decomposing claim verification into a 6-step pipeline.
-Additionally, MAPLE zeng2024maple, FactDetect jafari2024robustclaimverificationfact, and ClaimCheck putta-etal-2025-claimcheck implemented sophisticated claim-level fact-checking by combining claim type classification, evidence extraction, and verification automation.
-Furthermore, research analyzing citation quality in detail, such as ALiiCE xu2024aliiceevaluatingpositionalfinegrained and CiteEval xu2025citeevalprincipledrivencitationevaluation, which evaluate citation accuracy and evidence alignment in academic texts, is increasing.
-However, most approaches focus on verifying explicit citations or single claims and fail to address issues such as implicit evidence, contextual citation inheritance, and Fair Use that appear in long-form reports.
+With the proliferation of high-performing LLMs, LLM-as-a-Judge approaches—using LLMs as evaluators—have been widely proposed and studied Zheng et al.
+(2023); Liu et al.
+(2023b); Chiang et al.
+(2024); Kim et al.
+(2024a, b).
+In this line of work, early evaluations for deep research primarily focused on how well models solve expert-level, high-difficulty questions Phan et al.
+(2025); Rein et al.
+(2023).
+Subsequently, to directly assess a core property of deep research—accessing and leveraging external information from the open web—benchmarks were proposed to measure models’ ability to browse/search the web, synthesize the required information, and construct answers grounded in that evidence Gou et al.
+(2025); Wei et al.
+(2025a); Huang et al.
+(2025).
+More recently, beyond simple short-answer evaluation, studies have sought to evaluate the ability to generate long-form reports that require expert-level analysis, reasoning, and interpretation by integrating information from multiple sources Consult (2025); Coelho et al.
+(2025); Du et al.
+(2025).
+Despite this progress, existing report-evaluation methods are not based on expert-defined, fine-grained criteria.
+Some approaches rely on only a few coarse axes, leaving judgments largely to the evaluator LLM’s implicit standards Consult (2025); Coelho et al.
+(2025).
+Even when rubrics are subdivided, their specification and application can still depend substantially on the evaluator LLM Du et al.
+(2025), making it unclear whether the resulting scores correctly reflect report quality or align with expert assessment.
+In addition, source verification is often omitted Consult (2025); Coelho et al.
+(2025) or limited to a subset of cited claims Du et al.
+(2025), which may be insufficient to assess report-level factuality.
+To address these limitations, we propose an evaluation framework that (i) scores long-form research reports along multiple dimensions using expert-systematized, fine-grained criteria, and (ii) performs systematic source verification for claims throughout the report.
+A detailed comparison with prior work is provided in Appendix A.
 3 Data Construction
-To construct an evaluation dataset that reflects real-world usage, we analyzed 5,045 user queries collected from a deep research system and estimated the resulting topic distribution.
-For topic classification, we used the taxonomy of weborganizer, following prior work deepbench.
-We use Humanity’s Last Exam (HLE) hle as a source of seed questions because its expert-written, high-difficulty, multi-disciplinary questions align with the expert-level topics addressed in deep research reports.
-To match our target topic distribution, we mapped our topic domains to the 13 HLE subject domains and sampled 50 seed questions accordingly.
-Because HLE items are posed in a QA format, they require reformulation before they can serve as deep research report-generation tasks.
-We therefore asked domain experts (with at least a master’s degree or equivalent expertise in the relevant field) to review each item’s question, answer, and rationale to identify the underlying concepts, theories, and phenomena, and to reformulate the items as report- or paper-style topics for report generation.
-Each reformulated prompt was drafted by one domain expert and cross-reviewed by another expert from the same field, with iterative revisions as needed.
-During reformulation, we removed answer-revealing elements (e.g., factual conclusions, proofs, or specific answers) and instead provided only high-level writing directions, such as the intended scope of analysis and comparative perspectives, so that the model must develop the reasoning and narrative on its own.
-This reformulation shifts the task from reproducing a short answer to generating a report that requires expert-level analysis, reasoning, and interpretation.
-Detailed content is presented in Appendix B.
+To construct an evaluation dataset that reflects real-world usage, we analyze 5,842 in-house user queries collected from an internal deep research system and derive a topic distribution.
+For topic classification, we adopt the taxonomy of Wettig et al.
+(2025), consistent with prior work Du et al.
+(2025).
+We use Humanity’s Last Exam (HLE) Phan et al.
+(2025) as the source of seed questions because it provides expert-written, high-difficulty, multidisciplinary questions that align well with the expert-level topics addressed in deep research reports.
+To match our target topic distribution, we map our topics to the 13 HLE subject domains and sample 50 seed questions accordingly.
+Because HLE items are posed in a QA format, they need to be reformulated before they can be used as deep research report-generation queries.
+We therefore ask domain experts (each holding at least a master’s degree or possessing equivalent expertise in the relevant field) to review the original questions, answers, and rationales to identify the underlying concepts, theories, and phenomena, and to reformulate each item as a report- or paper-style prompt.
+Each reformulated prompt is drafted by one domain expert and cross-reviewed by another expert from the same field, with iterative revisions conducted as needed.
+During this process, we remove answer-revealing elements (e.g., factual conclusions, proofs, or specific answers) and retain only high-level writing directions, such as the intended scope of analysis and comparative perspectives.
+As a result, models are required to develop the reasoning and narrative independently.
+This reformulation shifts the task from producing a short answer to generating a report that requires expert-level analysis, reasoning, and interpretation.
+Detailed examples and procedures are provided in Appendix B.
 4 Approach
 4.1 Overview
-Figure 2: Overview of the DEER evaluation framework.
-(a) Research question and expert guidance generation from real-world deep research queries.
-(b) Construction of the Deep Research Evaluation Taxonomy consisting of 7 dimensions, 26 criteria, and 130 granular rubrics.
-(c) The DEER evaluation pipeline, integrating expert-guided LLM-as-a-judge scoring with claim extraction and information verification to assess deep research reports.
-To systematically evaluate deep research reports, we establish an evaluation framework grounded in two core aspects of deep research: external evidence acquisition and report-level synthesis characterizing; deepresearch_survey; deepresearch_survey2.
-Based on these aspects, we construct a Deep Research Report Evaluation Taxonomy with 7 major dimensions (§ 4.2), which we group into 5 report-quality dimensions and 2 external-information dimensions.
-Report-quality dimensions primarily assess synthesis and presentation via holistic, document-level judgment, whereas external-information dimensions assess external evidence acquisition and use via claim-level verification.
-Based on these differences in evaluation characteristics, we propose a hybrid architecture that integrates methodologies tailored to each component, as illustrated in Figure 2.
-Expert-Guided Report Quality Assessment (§ 4.3) uses an LLM-as-a-judge approach, combining a fixed set of granular evaluation items with report-specific Expert Evaluation Guidance authored by domain experts, to assess report quality against expert-designed criteria.
-Information Verification (§ 4.4) performs metric-based evaluation by automatically extracting claims and citations and verifying claims against external sources, producing quantitative metrics of external evidence sufficiency and integrity.
-Scores from the rubric items and verification metrics are aggregated into 7 dimension scores, which are then combined into an overall report score.
-Major Dimensions Sub-criteria
-Request Completeness Request Completeness, Scope Boundary, Value Validity
-Evidence Validity Numeric Accuracy, Logical Support
-Structure Consistency Structure Intro, Structure Body, Structure Conclusion, Structure Section
-Narration Style Report Form, Writing Quality, Paragraph, Reader Friendliness
-Ethics Compliance Sensitive Issues, Safety Impact, Perspective Balance
-Information Sufficiency Source Support, Information, Citations, References
-Information Integrity Claim Factuality, Citation Validity, Reference Accuracy, Reference Quality, Reference Diversity
-Table 1: Deep Research Report Evaluation Taxonomy: 7 Major Dimensions and 25 Sub-criteria
+To systematically evaluate deep research reports, we establish an evaluation framework grounded in two core aspects of deep research: external evidence acquisition and report-level synthesis Java et al.
+(2025); Xu and Peng (2025); Zhang et al.
+(2025).
+Based on these aspects, we construct a Deep Research Report Evaluation Taxonomy comprising seven major dimensions (§ 4.2), grouped into five report-quality dimensions and two external-information dimensions.
+The report-quality dimensions focus on synthesis and presentation through holistic, document-level judgment, whereas the external-information dimensions assess how external evidence is acquired and used via claim-level verification.
+Given these differences in evaluation characteristics, we propose a hybrid evaluation architecture that integrates methodologies tailored to each component, as illustrated in Figure 2.
+Expert-Guided Report Quality Assessment (§ 4.3) adopts an LLM-as-a-judge approach, combining a fixed set of granular rubric items with report-specific Expert Evaluation Guidance authored by domain experts to assess report quality against expert-designed criteria.
+Information Verification (§ 4.4) performs metric-based evaluation by automatically extracting claims and citations, then verifying claims against external sources, yielding quantitative measures of the sufficiency and integrity of external evidence.
+Scores from the rubric-based assessment and information-verification metrics are aggregated into seven dimension-level scores, which are then combined into an overall report score.
 4.2 Deep Research Report Evaluation Taxonomy
 In the absence of systematic evaluation criteria for deep research reports, we construct an evaluation taxonomy by synthesizing expert report assessment standards from multiple fields.
-We draw on internationally established standards spanning 20 domains of expertise, including systematic research reporting guidelines, technical and professional report writing and evaluation guidelines, and academic publishing norms, for a total of 80 standards.
+We draw on 80 established standards across 20 domains of expertise, including systematic research reporting guidelines, technical and professional report-writing and evaluation guidelines, and academic publishing norms.
 A panel of experts with experience in deep research system development and academic reviewing iteratively analyzes and consolidates these standards to derive report-quality dimensions.
-We then introduce complementary external-information dimensions that reflect deep research–specific characteristics of external evidence use and verification, resulting in 7 major dimensions and 25 detailed criteria in total (Table 1).
+We then introduce complementary external-information dimensions that reflect deep research-specific characteristics of external evidence use and verification, resulting in a taxonomy comprising seven major dimensions and 25 detailed sub-dimensions in total (Table 9).
 We validate the taxonomy through independent review by 10 experts from diverse domains.
-This hierarchical taxonomy organizes diverse quality elements into structured evaluation dimensions and sub-dimension, enabling systematic, multi-dimensional evaluation.
-It facilitates diagnosis of report strengths and weaknesses and guides systematic improvement.
+This hierarchical taxonomy organizes diverse quality elements into structured evaluation dimensions and subdimensions, enabling systematic, multi-dimensional assessment of deep research reports.
+It supports diagnosis of report strengths and weaknesses and provides a foundation for targeted improvement.
 Full descriptions, validation procedures, and the mapping from source standards to our criteria are provided in Appendix C.
 4.3 Expert-Guided Report Quality Assessment
-Existing evaluation methods have two limitations.
-First, existing methods apply broad, high-level criteria to long-form reports, giving LLM judges substantial discretion over what to focus on.
-As a result, judges may (i) examine only parts of a report, (ii) consider only a subset of the many sub-aspects implied by each criterion, and (iii) even focus on non-critical or superficial sub-aspects, increasing variance across judges and reducing evaluation consistency webthinker; deepconsult; deepgym.
-Second, even with a clear evaluation focus, assessing correctness and completeness often requires domain knowledge, which LLM judges may lack.
-As a result, they may miss subtle but important issues such as non-obvious logical leaps, domain-specific misinterpretations, and fine-grained inaccuracies deepbench.
-This study addresses these issues by (1) making the evaluation well-specified through a fixed set of fine-grained rubric items and (2) providing task-specific Expert Evaluation Guidance that includes domain-specific context and reference points, helping LLM judges identify hard-to-detect errors and omissions.
+Existing evaluation methods exhibit two key limitations when applied to long-form expert reports.
+First, they typically apply broad, high-level criteria, granting LLM judges substantial discretion over what aspects of a report to attend to.
+As a result, judges may (i) examine only parts of a report, (ii) consider only a subset of the many sub-aspects implied by each criterion, and (iii) focus on non-critical or superficial elements, leading to increased variance across judges and reduced evaluation consistency Li et al.
+(2025b); Consult (2025); Coelho et al.
+(2025).
+Second, even when the evaluation focus is clearly defined, assessing correctness and completeness often requires domain expertise that LLM judges may lack.
+Consequently, they may miss subtle but important issues, including non-obvious logical leaps, domain-specific misinterpretations, and fine-grained inaccuracies Du et al.
+(2025).
+To address these limitations, this study (1) makes evaluation criteria explicit through a fixed set of fine-grained rubric items and (2) provides task-specific Expert Evaluation Guidance that includes domain-specific context and reference points, enabling LLM judges to identify hard-to-detect errors and omissions.
 Granular Rubric Design.
-Broad evaluation criteria can lead to inconsistent LLM-judge-based scoring.
-Recent work addresses this by decomposing broad criteria into granular rubrics checkeval; rocketeval; expertlongbench.
-For deep research report evaluation, deepbench follows a similar direction by using an LLM to generate task-specific evaluation criteria.
-While this narrows the evaluation focus, it raises two concerns: whether the generated criteria are sufficiently complete and capture what matters most in expert reports, and whether dynamically changing criteria enable interpretable, comparable scores that support systematic diagnosis of system strengths and weaknesses across tasks.
+Broad evaluation criteria can lead to inconsistent LLM-based scoring.
+Recent work addresses this issue by decomposing high-level criteria into more granular rubrics Lee et al.
+(2025); Wei et al.
+(2025b); Ruan et al.
+(2025).
+For deep research report evaluation, Du et al.
+(2025) follows a related approach by using an LLM to generate task-specific evaluation criteria.
+While this strategy narrows the evaluation focus, it raises two important concerns: whether the generated criteria are sufficiently comprehensive and capture what matters most in expert reports, and whether dynamically varying criteria enable interpretable and comparable scores that support systematic diagnosis of system strengths and weaknesses across tasks.
 This study uses a fixed, expert-designed rubric to ensure reliable and interpretable evaluation.
-Building on the taxonomy in Table 1, a panel of experts operationalizes the 25 sub-dimensions into 40 evaluation criteria and translates each criterion into concrete, checkable rubric items.
-Rubric items are organized into two aspects: coverage (whether the required components are present at all relevant points in the report and addressed in full whenever they occur) and quality (the degree to which the targeted components are executed to a high standard in the report).
-This process yields 130 scorable rubric items in total.
-[^1] The 40 criteria were validated through independent review by 10 experts following the procedure in § 4.2.
+Building on the taxonomy in Table 9, a panel of experts operationalizes the 25 sub-dimensions into 46 evaluation criteria and translates each criterion into concrete, checkable rubric items.
+The items are organized into two aspects: coverage, whether required components are present and fully addressed wherever they occur in the report; and quality, the degree to which the targeted components are executed to a high standard.
+This process yields 101 scorable rubric items in total.
+[^1] The 46 criteria were validated through independent review by 10 experts following the procedure in § 4.2.
 The rubric is applied identically to all deep research reports across 50 tasks, enabling diagnosis of the system’s overall strengths and weaknesses with richer, more interpretable signals (see Appendix E for the rubric structure and examples).
 Expert Evaluation Guidance.
-Even with a well-specified, granular rubric, scoring individual items can still require judgment informed by deep domain knowledge, especially for specialized and complex tasks—for example, to assess whether the report’s reasoning is sound, its calculations are correct, and its coverage and depth are adequate for the topic.
-In such cases, evaluators may miss subtle domain-specific errors or omissions that human experts would catch.
-To mitigate this, we provide task-specific expert evaluation guidance that enumerates the required content elements and expert expectations for each task, presented as concrete, checkable statements that evaluators can verify.
-For each task, an expert with at least a master’s degree in the relevant field was assigned to detail the key concepts, theories, data, and analytical perspectives that the report must cover from an expert perspective.
-The expert first modified the original query to naturally include the essential content of the topic (Section 3), and then wrote the Expert Evaluation Guidance based on the modified query, specifying the content elements that must be included.
-This approach is based on the logic that if a report responds appropriately to the modified query, it should naturally contain the key content of the topic, so the omission of key content can be clearly judged as a quality issue.
-Each written Guidance was verified through cross-review by another expert in the same field.
-See Appendix B for detailed methodology.
+Even with a well-specified, fine-grained rubric, evaluating expert-level reports often requires substantial domain knowledge.
+In such settings, non-expert evaluators—including LLM judges—may fail to detect subtle domain-specific errors or omissions that subject-matter experts would identify.
+To mitigate this risk, we introduce task-specific Expert Evaluation Guidance, which provides domain-grounded context and reference points to support consistent and informed evaluation.
+Expert Evaluation Guidance enumerates the required content elements and expert expectations for each task as concrete, verifiable statements that can be checked directly against the report.
+The guidance for each task is produced using the same expert drafting and cross-review procedure as the query reformulation process described in Section 3.
+Specifically, a domain expert reformulates each HLE item into a report-style prompt with explicit requirements (Section 3).
+The expert then derives the guidance by identifying the substantive content an adequate expert report should cover under the task prompt.
+Each required element is expressed as a concrete, verifiable statement that can be checked against the report.
+The resulting guidance is subsequently cross-reviewed by another expert from the same field to identify omissions, redundancies, or ambiguities, and revised as needed.
+Through this process, Expert Evaluation Guidance complements the fixed rubric by anchoring evaluation in task-specific expert knowledge while maintaining consistency and interpretability across tasks.
+Appendix B provides detailed guidance construction procedures and illustrative examples.
+Model Req. Ful. Analyt. Sound. Struct. Coh. Format & Style Info. Int. Info. Suff. Ethics Mean
+General LLMs
+Qwen3-235B (fast) 4.51 5.02 6.09 7.49 1.24 4.20 7.19 5.11
+Gemini 2.5 Flash (fast) 4.64 5.33 6.55 7.85 1.30 3.99 7.52 5.31
+Claude Opus 4.5 (fast) 4.94 5.48 6.54 7.99 2.29 4.50 7.78 5.65
+GPT-5 (fast) 4.11 4.75 5.84 7.21 1.05 3.13 7.30 4.77
+LLMs+Reasoning
+Qwen3-235B (think) 5.00 5.33 6.64 7.88 1.12 3.90 7.38 5.32
+Gemini 2.5 Pro (think) 4.88 5.81 6.99 8.09 2.23 4.40 7.73 5.73
+Claude Opus 4.5 (think) 4.96 5.48 6.68 8.10 2.27 4.22 7.73 5.63
+GPT-5 (think) 5.57 6.18 7.00 8.06 2.11 4.16 8.08 5.88
+LLMs+Reasoning+WebSearch
+Qwen3-235B (think+search) 4.05 4.34 5.68 6.83 5.22 5.45 7.06 5.52
+Claude Opus 4.5 (think+search) 4.52 5.13 5.99 7.41 7.03 7.62 7.37 6.44
+GPT-5 (think+search) 5.57 6.08 6.97 8.15 5.63 6.17 8.11 6.67
+Deep Research
+WebThinker 4.11 4.64 5.51 7.35 6.21 6.40 7.13 5.91
+Qwen3-235B (deep) 4.13 4.69 4.85 7.06 6.55 7.90 7.43 6.09
+Gemini 2.5 Pro (deep) 4.71 5.37 6.25 7.59 6.01 7.61 7.39 6.42
+Claude Opus 4.5 (deep) 4.53 5.22 5.69 7.22 6.04 5.66 7.57 5.99
+OpenAI (deep) 4.67 5.29 6.28 7.66 7.14 6.89 7.48 6.49
+Table 1: Baseline models’ expert report evaluation results (updated). In each column, the highest score is marked in bold and the second highest score is underlined.
 4.4 Information Verification
-To automate fact verification in expert reports, we must overcome the limitations of existing systems that focus solely on explicit citations.
+To automate information verification in expert reports, we must overcome the limitations of existing systems that focus solely on explicit citations.
 In reality, many sentences implicitly inherit preceding evidence or construct new claims through contextual connections.
@@
 Batch Processing for Claim Extraction
-When processing long documents as a single input, LLMs suffer from the Lost-in-the-Middle problem, which leads to the omission of important claims liu2023lostmiddlelanguagemodels.
+When processing long documents as a single input, LLMs suffer from the Lost-in-the-Middle Liu et al.
+(2023a) problem, which leads to the omission of important claims.
 To address this, we apply a Batch Extraction strategy that maintains the context of the entire document $D$ while parallel-processing only the target sentences divided into small batches $B_{j}$:
@@
 First, we apply Grouped Claim Verification, which bundles multiple claims citing the same source into a single input for verification.
-Experimental results showed that increasing Group Size from 1 to 10 raised accuracy from 0.78 to 0.89 and reduced cost per claim by 85% (Section 6.5).
-This indicates that grouped input helps in understanding the overall structure of the document and verifying claims, while reducing verification overhead.
+Experimental results showed that increasing Group Size from 1 to 20 raised F1 from 87.25 to 91.61 and reduced cost per 1K claims by 72%.
+This indicates that grouped input improves understanding the overall structure of the document and verifying claims, while reducing verification overhead.
 Second, greater API cost reduction is possible through the Context Retrieval strategy, which selectively provides only the context necessary for fact verification.
-This method achieved an API cost reduction of approximately 24 times, but a 6% drop in accuracy was observed, indicating a clear trade-off.
+This method achieved an API cost reduction of approximately 72%, but a 3.4% drop in F1 was observed, indicating a clear trade-off.
 This hierarchical pipeline (Batch Extraction $\rightarrow$ Grouped Verification $\rightarrow$ Context Retrieval) is designed to achieve both high-precision verification and cost efficiency even in long reports containing hundreds of claims.
 5 Experiment Setup
-5.1 Implementation Details.
-We use the LLM-as-a-judge method to evaluate report quality.
-The evaluation model takes as input the task query, the report to be evaluated, expert guidance, and detailed rubrics and evaluation criteria.
-The evaluation is conducted across 5 dimensions: Request Completeness, Evidence Validity, Structure Consistency, Narration Style, and Ethics Compliance, with each rubric item scored on a 1–10 scale.
-The model generates a score and a short rationale justifying that score for each item, and the final output is in JSON format consisting of (rationale, score) pairs for each item.
-Information Integrity and Information Sufficiency are evaluated through a separate verification pipeline, the Information Verification Module.
-This module extracts claims from the report and assesses their factuality by comparing them with external evidence obtained via web searches.
-The verification results are calculated as detailed metrics such as Claim Factuality and Citation Validity.
-The evaluation is hierarchical, computing the final quality score by sequentially aggregating the averages of lower-level rubric items to higher levels, using the scores collected above.
-Model Req. Comp. Evid. Valid. Struct. Cons. Narr. Style Info. Int. Info. Suff. Ethics Mean
-General LLM
-qwen3-235b-fast 6.10 6.59 9.04 8.66 3.30 4.40 8.58 6.67
-gemini-2.5-pro-fast 6.08 7.02 9.29 8.87 2.92 4.43 8.89 6.79
-claude_opus4.1_fast 6.04 6.47 8.96 8.95 4.76 4.81 8.78 6.97
-gpt5_fast 5.61 6.63 8.80 8.46 2.18 3.45 8.82 6.28
-LLM+Thinking
-qwen3-235b-think 6.77 6.53 9.32 8.92 3.28 4.39 8.75 6.85
-gemini-2.5-pro-think 6.41 7.66 9.55 9.20 3.86 4.97 9.37 7.20
-claude_opus4.1_think 5.93 6.41 9.05 9.04 4.72 5.12 8.99 7.04
-gpt5_think 7.69 8.37 9.57 9.32 4.19 5.42 9.37 7.70
-LLM+Thinking+Search
-qwen3-235b-think_search 6.43 6.31 9.07 8.82 3.60 4.49 8.44 6.74
-gemini-2.5-pro-think_search 6.61 7.58 9.50 9.11 4.62 5.02 8.97 7.34
-claude_opus4.1_think_search 6.03 6.62 9.04 9.03 5.17 4.80 8.85 7.08
-gpt5_think_search 7.88 8.26 9.39 9.04 6.42 6.44 9.38 8.12
-Deep Research
-webthinker 5.27 5.76 8.48 8.62 7.57 6.92 8.68 7.33
-qwen3-235b-deep 4.95 6.00 7.28 8.31 7.13 8.29 8.66 7.23
-gemini-2.5-pro_deep 6.08 6.81 8.81 8.78 7.13 7.87 8.62 7.73
-claude_opus4.1_deep 6.60 6.91 8.90 8.85 6.79 7.66 8.75 7.78
-gpt5_deep 6.26 7.11 9.20 8.93 7.64 7.13 8.99 7.89
-Table 2: Baseline models’ expert report evaluation results. In each column, the highest score is marked in bold and the second highest score is underlined.
-5.2 Baseline Models
-In this study, to compare the expert report generation capability of LLM-based systems, we established four categories of baseline models.
-Each category is designed to increase in complexity incrementally, ranging from simple language model response generation to specialized research systems that combine search and deep reasoning.
-(1) General LLMs This is the basic LLM setting that generates responses in a single pass (inference), serving as a baseline for measuring response quality in language-generation-based models.
-This category includes qwen3-235b-fast qwen3technicalreport, gemini-2.5-pro-fast google2025gemini, claude_opus4.1_fast anthropic2025claude, and gpt5_fast openai2025gpt5system models.
-(2) LLMs with Reasoning This setting uses the reasoning function on the same LLMs as in (1), extending the internal reasoning step before answer generation.
-This category includes qwen3-235b-think, gemini-2.5-pro-think, claude_opus4.1_think, and gpt5_think models.
-(3) LLMs with Reasoning + Web Search This configuration combines the LLMs with the Reasoning setting and web search, evaluating scenarios in which external information is used during reasoning.
-This category includes qwen3-235b-think_search, gemini-2.5-pro-think_search, claude_opus4.1_think_search, and gpt5_think_search models.
-(4) Deep Research Frameworks (Open-Source & Commercial) This is the upper baseline, comprising specialized deep research frameworks and commercial research support systems.
-We used webthinker webthinker as the open-source framework, and commercial systems include qwen3-235b-deep, gemini-2.5-pro_deep, claude_opus4.1_deep, and gpt5_deep.
-Detailed configurations and hyperparameters for all baseline models are summarized in Appendix A.
-6 Experiments
-6.1 Main Results
-Summarizing Table 2, all models show generally high performance in metrics related to form, style, and ethics, such as Structure Consistency (Struct.
-Cons.), Narration Style (Narr.
-Style), and Ethics Compliance (Ethics).
-However, they consistently show relatively lower scores on metrics that assess how faithfully they address the requirements in the user’s query and provide persuasive evidence in terms of calculation, methodology, and logical development, such as Request Completeness (Req.
-Comp.) and Evidence Validity (Evid.
-Valid.).
-This indicates that while the latest LLMs are already quite proficient in structuring reports, writing naturally, and adhering to basic ethical standards, their ability to accurately meet user intent at an expert level and support it with sufficiently robust argumentation and evidence structures remains limited.
-Reasoning models showed consistent improvements across almost all evaluation metrics relative to General LLMs, suggesting that simply expanding the reasoning budget without additional tools improves coverage of requirements, logical development, evidence selection, and narrative consistency.
-Models combining this with web search show additional gains in information-related metrics.
-The Deep Research category consistently achieves higher scores in Information Integrity (Info.
-Int.) and Information Sufficiency (Info.
-Suff.) than other categories, providing the strongest baseline for the broad collection and organization of external documents.
-However, it is characteristic that the Deep Research category consistently records lower scores than reasoning-based single LLMs in the remaining key evaluation axes, especially Req.
-Comp.
-and Evid.
-Valid.
-These results suggest that the large amount of external material injected through search in the Deep Research configuration may cause the reference point for report development to gradually shift from the query itself to external knowledge, inducing a kind of retrieval-induced drift in which information rather than blurs the problem definition and argumentation structure.
-Therefore, this demonstrates that simply increasing the volume of search and citations does not guarantee the quality of expert reports.
-6.2 Fine-grained Analysis of Evaluation Results
+5.1 Implementation Details
+We evaluate report quality using an LLM-as-a-judge approach across five dimensions—Request Fulfillment, Analytical Soundness, Structural Coherence, Format & Style, and Ethics Compliance.
+The LLM judge receives the task query, the report being evaluated, task-specific Expert Evaluation Guidance, and a fixed set of fine-grained rubric items.
+It assigns a 1–10 score and a brief rationale to each item and returns the results as a JSON object mapping each item to its score and rationale.
+In parallel, Information Integrity and Information Sufficiency are assessed by the Information Verification Module, which extracts claims from the report and verifies them against external evidence retrieved via web search, producing metrics such as Claim Factuality and Citation Support.
+Details are provided in Appendix G.4.
+Finally, we compute the overall quality score hierarchically by aggregating scores bottom-up from rubric items to higher-level categories.
+Details of the scoring procedure are provided in Appendix E.3.
 Figure 3: Heatmap visualizations of expert report evaluation results.
@@
 (b) Domain-wise scores averaged over two representative tasks randomly sampled from each domain.
-Figure 3(a) decomposes the upper evaluation dimensions into detailed items to summarize the performance patterns of current Deep Research and LLM-based systems in expert reports.
-In Request Completeness, scores across sub-items are generally low, with scope the lowest, suggesting a lack of clarity about what the report covers and excludes, and under what assumptions and limitations the discussion is developed.
-In Evidence Validity, the score for the numeric item is slightly lower than that for logical, indicating that current systems make mistakes relatively more frequently in terms of quantitative accuracy, such as calculation processes, used formulas/statistical models, and selection of indicators/units, rather than logical development itself.
-In Information-related detailed items, ref_diversity and refs_amount appear commonly low, meaning that even if the overall structure and narrative are compliant, the number and diversity of sources actually referenced are limited.
-Figure 3(b) shows the domain-specific performance of Deep Research agents as the average of two randomly selected tasks in each domain (domain-specific average scores for all tasks are presented in Appendix LABEL:app:domain-results).
-While most domains are distributed in a relatively similar range around 7 points, in the case of Physics, most Deep Research agents except gpt5_deep consistently record the lowest scores compared to other domains, and relatively large deviations between Deep Research agents also appear in CS, History, and Linguistics.
+5.2 Baseline Models
+To compare expert report generation performance, we consider four baseline families: General LLMs (fast), LLMs+Reasoning (think), LLMs+Reasoning+WebSearch (think+search), and Deep Research (deep).
+Each family is instantiated using multiple model families/backbones (Qwen, Gemini, Claude, and GPT), and they differ in whether they (i) use reasoning, (ii) use web search, and (iii) employ research-system orchestration.
+Detailed model configurations are provided in Appendix H.
+6 Experiments
+6.1 Main Results
+Summarizing Table 1, most models score highly on structure, style, and ethics (Structural Coherence, Format & Style, Ethics), but lag on requirement fulfillment and analytical soundness (Req.
+Fulfillment, Analytical Soundness).
+This suggests that report-writing quality is largely mature, whereas expert-level intent alignment and evidence-based justification remain limited.
+Across baseline families, adding reasoning in think improves overall performance over fast.
+Moreover, external-information metrics—Information Integrity and Information Sufficiency—see further gains with think+search and deep.
+An interesting finding is that reasoning models without web search (think) outperform think+search and deep on report-quality metrics excluding information-related scores.
+This suggests that integrating diverse external information can blur the problem definition and argument structure.
+Accordingly, simply adding search and external sources may not guarantee improved report-writing quality.
+6.2 Fine-grained Analysis
+Figure 3(a) breaks down the performance patterns of Deep Research systems on expert report generation into fine-grained dimensions.
+The sub-dimension scores for Request Fulfillment, Analytical Soundness, and Structural Coherence are generally low, with scope particularly low within Request Fulfillment.
+This indicates that Deep Research systems often fail to clearly specify the report scope (what is covered vs.
+excluded) and the assumptions/constraints used to develop the report.
+In addition, ref_amount is low in Information Sufficiency, suggesting that systems tend to rely on a small number of references rather than leveraging a broad set of sources.
+Figure 3(b) shows domain-level performance of Deep Research agents.
+For each domain, we report results on two tasks randomly selected from that domain.
+Performance tends to drop uniformly across agents on relatively difficult tasks, and Physics consistently records the lowest scores compared to other domains.
 6.3 Correlation with Human Evaluation
-To verify the validity of the proposed evaluation method and analyze the contribution of each component, we conducted experiments to assess its correlation with human expert evaluation.
-For 15 tasks across 5 domains, 3 reports randomly selected from those generated by the Deep Research systems used in 6.1 were used as evaluation targets.
-Each report was evaluated by two independent experts, collecting a total of 90 human evaluation samples.
-To assess the alignment between the evaluation method and human experts, we used the Pearson correlation coefficient, the Spearman rank correlation coefficient, and pairwise agreement deepbench.
-Pairwise agreement is a metric that measures the proportion of cases in which our evaluator chose the same priority as the human expert when comparing two reports, indicating how consistently it reproduces human relative judgments across tasks.
-Evaluations were performed on the same reports using 5 LLM-based evaluation models (GPT-5, GPT-5-mini, Claude Opus 4.1, Claude Sonnet 4.5, Gemini 2.0 Pro), and the correlation between the average of the 5 model evaluations and the human evaluation was analyzed.
 Evaluation Method Pearson Spearman Pairwise
 $r$ $\rho$ Agr.
-Vanilla 0.64 0.61 0.66
-+ Dimensions 0.67 0.65 0.80
-+ Granular Checklist 0.53 0.46 0.72
-+ Expert Guidance 0.75 0.71 0.84
+Vanilla 0.64(0.16) 0.61(0.17) 0.66(0.14)
++ Dimensions 0.67(0.10) 0.65(0.14) 0.80(0.07)
++ Granular Rubrics 0.62(0.14) 0.59(0.17) 0.78(0.08)
++ Expert Guidance 0.75(0.07) 0.71(0.06) 0.84(0.03)
 Inter-Human 0.81 0.74 0.79
-Table 3: Average human correlation across five AI models with incremental addition of evaluation components. Inter-Human shows inter-annotator agreement.
-Table 3 shows how the correlation with human evaluation changes as evaluation components are added step by step.
-Starting from the Vanilla deepbench setting, which judges overall quality without explicit evaluation criteria, we sequentially expanded by introducing evaluation dimensions (+Dimensions), decomposing them into granular rubrics (+Granular Rubric), and finally adding task-specific expert guidance (+Expert Guidance) to measure the impact of each component.
-The Vanilla setting showed significant correlation with human experts even without explicit criteria, but this correlation decreased as evaluation dimensions (+Dimensions) were introduced and decomposed into detailed items (+Granular Rubric).
-Since report evaluation is a high-dimensional task that requires simultaneous consideration of various elements, such as logic, concepts, analysis, and domain knowledge, across long and complex texts, each rubric item comprises broad, interpretive judgment units rather than narrow rules.
-It is analyzed that the increase in the number of items led to an increase in interpretive variability rather than refinement, resulting in a drop in human correlation.
-Adding task-specific expert guidance (+Expert Guidance) clarifies the interpretation method for each criterion, restoring consistency and significantly improving correlation across all metrics.
-Writing longer rationales (+Longer Rationale) further improves correlation by enriching the basis for each item.
-In conclusion, DEER tends to achieve relatively stable criterion application and high alignment with human experts, even in report evaluation settings that must consider multiple evaluation dimensions simultaneously.
-6.4 Inter-evaluator Agreement (IEA)
-To verify that the proposed evaluation method yields consistent results across different LLMs, we measured inter-evaluator reliability among the 5 evaluation models used previously IAA; checkeval.
-We measured the agreement between models using Krippendorff’s $\alpha$ and ICC (Intraclass Correlation Coefficient).
-High agreement suggests that the proposed method possesses versatility and stability that are not dependent on a specific LLM.
+Table 2: Average human correlation across five AI models with incremental addition of evaluation components (reported as mean(std)). Inter-Human shows inter-annotator agreement.
+To validate the proposed evaluation method and assess the contribution of each component, we compared LLM-based evaluations against human expert judgments.
+For each of the 45 reports, we collected two independent ratings from domain experts whose expertise aligned with the report’s topic (90 ratings total).
+We computed Pearson’s r, Spearman’s ρ, and LLM–human pairwise agreement for each of the five LLM-based evaluator models, and report the average of each metric across models.
+Additional details on the experimental setup and human evaluation protocol are provided in Appendix I.
+Table 2 shows how alignment with human judgments changes as evaluation components are added step by step.
+Vanilla, which assesses overall quality holistically, and +Dimensions, which introduces high-level evaluation dimensions, both achieve relatively high correlation with human judgments.
+However, Pearson’s r and Spearman’s ρ drop at the +Granular Rubrics stage where the rubric is further decomposed into many fine-grained items.
+In contrast, task-specific +Expert Guidance helps evaluators apply these rubric items by surfacing domain-relevant cues that non-experts may overlook, thereby achieving the highest correlation with human judgments.
+6.4 Inter-evaluator Reliability
 Method Krip. $\alpha$ ICC(2,1) ICC(2,k)
@@
 + Dimensions 0.32 0.37 0.75
-+ Granular Checklist 0.17 0.24 0.62
++ Granular Rubrics 0.33 0.38 0.76
 + Expert Guidance 0.55 0.56 0.87
 Table 3: Inter-evaluator reliability across five LLM-based evaluation models. Krip. $\alpha$: Krippendorff’s alpha. Higher values indicate more consistent evaluations.
-Table 4 shows how the agreement between evaluation models changes as evaluation components are added step by step.
-The Vanilla setting showed an appropriate level of inter-model agreement, but agreement decreased significantly when introducing evaluation dimensions (+Dimensions) and decomposing them into detailed items (+Granular Rubric).
-As explained earlier, this is because rubric items consist of interpretive judgment units, so the increase in the number of items increased the variability in interpretation across models.
-Adding expert guidance (+Expert Guidance) significantly improved agreement across all metrics, restoring evaluation consistency.
-In particular, the variation focusing only on the top rubric items achieved the highest agreement.
-This shows that expert guidance clarifies the interpretation of evaluation criteria, enabling consistent evaluation even between different LLM evaluation models.
+To verify that the proposed evaluation method yields consistent results across different LLM evaluators, we measure inter-evaluator reliability Artstein (2017); Lee et al.
+(2025).
+Specifically, we compute Krippendorff’s $\alpha$ and the intraclass correlation coefficient (ICC) using the scores from the five LLM judges described in Section 6.3.As shown in Table 3, Vanilla achieves moderate inter-evaluator reliability, but reliability drops substantially under +Dimensions.
++Granular Rubrics yields a slight recovery, though still below Vanilla.
++Expert Guidance attains the highest reliability across metrics.
+This suggests that expert guidance clarifies what to look for under each criterion, enabling more consistent judgments across different LLM judges.
 6.5 Information Verification Module Evaluation
-In this section, we evaluate the proposed Information Verification Module by dividing it into two stages: claim extraction and fact verification.  [^2] Detailed content is covered in Appendix F.
+In this section, we evaluate the proposed Information Verification module by dividing it into two stages: claim extraction and claim verification.  [^2] Detailed content is covered in Appendix F.
 Claim Extraction Analysis
-We compared and analyzed 728 claims identified by human experts (Ground Truth) and LLM extraction results using a Paragraph-Level LLM Judge (Table 5).
-As a result of measuring semantic Recall beyond simple count comparison, GPT-5-mini showed significantly higher performance compared to GPT-4.1.
-In particular, the GPT-5-mini (Batch 5) setting recorded a Recall of 0.9333 for Type A, B, and C claims requiring external evidence, capturing human expert claims almost perfectly.
-Also, the number of claims extracted per paragraph (Claims/Para) was the highest at 6.95, confirming that detailed information in the document is extracted densely without omission.
-On the other hand, GPT-4.1 showed relatively lower Recall (about 0.83) and density (about 4.8), suggesting a tendency for the model to excessively summarize information or omit some details.
-In conclusion, for precise fact-checking, a strategy of maximizing recall by operating high-performance models like GPT-5-mini with small batch sizes (Batch Size 5, 10) is essential.
-Table 5: Comparison of Claim Extraction Performance across Models and Batch Sizes. Recall is measured at the Paragraph-Level using LLM Judge.
-Model Batch Size Recall Claims/Para
-Ground Truth - - 7.22
-GPT-4.1 5 0.9057 5.00
-10 0.8897 4.84
-20 0.8897 4.65
-GPT-5-mini 5 0.9333 6.95
-10 0.9264 6.22
-20 0.9218 6.05
+Table 4 compares claim extraction performance across models and batch sizes using paragraph-level semantic recall measured by an LLM Judge.
+GPT-5-mini achieves the highest recall (92.17%) and classification F1 (68.80) at batch size 10, while incurring substantially lower cost than GPT-5.
+Although GPT-5 extracts more claims per paragraph, its recall and F1 saturate, indicating that denser extraction does not improve coverage.
+Increasing the batch size consistently reduces cost while incurring only moderate degradation in recall, making GPT-5-mini with a batch size of 20 a cost-effective configuration for large-scale claim extraction.
+Model Batch Density Recall Cls. F1 Cost (\$)
+Ground Truth - 7.22 - - -
+GPT-5 10 6.27 89.42 64.65 0.92
+20 6.04 89.29 66.07 0.59
+GPT-5-mini 10 5.54 92.17 68.80 0.16
+20 5.19 90.66 67.52 0.10
+Table 4: Comparison of claim-level extraction Density (claims per paragraph), recall, and classification F1 across models (low effort) and batch sizes.
 Claim Verification Performance
-In the claim verification stage, we conducted an ablation study on grouping and retrieval settings using GPT-4.1 (Table 6).
-When applying Grouped Claim Verification without retrieval, the F1 score remained comparable (91 → 90), while the cost per claim decreased by approximately 82%, demonstrating that batching claims provides substantial efficiency gains with minimal performance loss.
-Additionally, enabling retrieval further reduced the cost to \$0.004 per claim but lowered the F1 score to 85.
-Therefore, for scenarios requiring high reliability, grouped verification without a retrieval configuration offers the best balance, while with a retrieval setting, it is preferable when maximizing cost efficiency is the primary objective.
-Grouped Retrieval F1 Cost (\$/Claim)
-0.91 0.083
-20 0.90 0.015
-20 0.85 0.004
-Table 6: Ablation study on GPT-4.1. Grouping significantly improves cost efficiency with minimal F1 degradation; retrieval further reduces cost at the expense of accuracy.
+Grouped Retrieval F1 Cost (\$/1k)
+87.25 12.84
+10 90.91 3.65
+10 83.10 0.95
+20 91.61 3.46
+20 87.25 0.95
+Table 5: Ablation study on GPT-5-mini (low effort) comparing grouped verification and retrieval.
+Table 5 presents an ablation study on grouped verification and retrieval using GPT-5-mini.
+Grouping multiple claims substantially reduces cost while maintaining strong verification accuracy.
+Without retrieval, increasing the group size to 20 achieves the highest F1 score (91.61) at a low cost.
+Retrieval further reduces the cost to \$0.95 per 1k claims but introduces a clear drop in accuracy, highlighting a trade-off between verification fidelity and efficiency.
+Overall, grouped verification without retrieval provides the best accuracy–cost balance, while retrieval-augmented settings are suitable for budget-constrained scenarios.
 7 Conclusion
-We introduce DEER, a comprehensive benchmark for evaluating Deep Research agents, featuring a multi-dimensional taxonomy and a hybrid evaluation framework that combines LLM-based qualitative assessment with a retrieval-augmented Information Verification Module.
-Our experiments reveal that while state-of-the-art models excel in structure and style, they still lag behind human experts in critical areas like evidence validity and information sufficiency, although "Deep Research" systems show promising improvements.
-By demonstrating a high correlation with human judgments, DEER offers a scalable standard for advancing autonomous research agents toward trustworthy, expert-level performance.
+We propose DEER, a benchmark and evaluation framework for systematically assessing deep research reports.
+DEER builds a hierarchical taxonomy, instantiates it as 101 fixed rubrics, and provides task-specific expert guidance to improve the reliability of LLM-based judging.
+Beyond rubric scoring, DEER evaluates information use by tracing all claims back to their external information sources to verify evidence and quantifying evidence quality, enabling a more complete report-level assessment.
+Experiments show that deep research systems perform well on structure/style and information use, but remain limited in meeting expert-level requirements and producing analytically sound analyses.
+With its taxonomy, fixed rubrics, and quantitative metrics, DEER supports fine-grained diagnosis and systematic improvement beyond mere performance assessment.
 Limitations
@@
 Furthermore, although our current benchmark focuses on text-based reports, this specialization enables a deeper, more rigorous analysis of information integrity and logical coherence, laying a solid foundation for future extensions to multimodal research tasks.
-Acknowledgments
-Benchmark Human rubrics Expert curated Open- ended Non-tech domains LLM judge Claim fact-check Inter- pretability Systematic criteria Avg. rubrics
-AcademicBrowse AcademicBrowse
-BrowseComp browsecomp
-Mind2Web2 mind2web2 $\triangle$ 50
-ResearchBench ResearchBench
-ResearcherBench ResearcherBench $\triangle$ $\triangle$ 14
-LiveDRBench characterizing $\triangle$
-SPOT SPOT $\triangle$
-ExpertLongBench expertlongbench 16
-ReportBench ReportBench $\triangle$
-DeepScholar-Bench DeepScholar-Bench $\triangle$ $\triangle$
-DeepResearchBench deepbench $\triangle$ $\triangle$ 25
-LiveResearchBench LiveResearchBench $\triangle$ $\triangle$
-DeepResearchArena DeepResearchArena $\triangle$ $\triangle$
-DeepResearchGym deepgym $\triangle$
-ResearchRubrics ResearchRubrics $\triangle$ 26
-DREX (Ours) 130+
-Table 7: Comparison of DREX with representative deep research benchmarks.
-Here, indicates full support, no support, and $\triangle$ partial support.
-For Claim fact-check, $\triangle$ means that only a subset of claims (e.g., explicitly cited or gold-labeled ones) are checked rather than all explicit and implicit claims.
-For Interpretability, $\triangle$ means that the benchmark provides only coarse, dimension-level insight (e.g., a few high-level scores) without shared checklist-level diagnostics across tasks.
 Appendix A Comparison with Deep Research Benchmarks
 In this section, we compare existing deep research benchmarks with our methodology.
-Table 7 extends and modifies the comparison table proposed in ResearchRubrics ResearchRubrics, and adds new evaluation axes to summarize how DREX differs from prior work.
+Table 6 extends and modifies the comparison table proposed in ResearchRubrics Sharma et al.
+(2025), and adds new evaluation axes to summarize how DEER differs from prior work.
 The ResearchRubrics table uses five axes (whether rubrics are human-written, whether experts curate tasks, whether tasks are open-ended, whether non-technical domains are included, and whether an LLM-as-judge is used).
@@
 The first block of the table consists of benchmarks for evaluating search and browsing agents.
-AcademicBrowse AcademicBrowse, BrowseComp browsecomp, and Mind2Web2 mind2web2 evaluate, respectively, the ability to browse an academic corpus or the open web to produce short answers to complex queries, the ability to perform agentic search across diverse websites, and the consistency between generated answers and cited sources.
+AcademicBrowse Zhou et al.
+(2025), BrowseComp Wei et al.
+(2025a), and Mind2Web2 Gou et al.
+(2025) evaluate, respectively, the ability to browse an academic corpus or the open web to produce short answers to complex queries, the ability to perform agentic search across diverse websites, and the consistency between generated answers and cited sources.
 These benchmarks focus on “how well the system finds information,” i.e., search and browsing strategies, and thus are one step removed from the deep research report quality evaluation that we study.
 The second block covers benchmarks that focus on specific component abilities or conceptualizations of deep research.
-ResearchBench ResearchBench and ResearcherBench ResearcherBench evaluate, respectively, the ability to generate inspiration and hypotheses from papers, and to provide long-form advice on frontier AI research questions based on expert rubrics.
-Characterizing Deep Research characterizing and LiveDRBench evaluate the recovery of correct claims in search tasks that require many information units and non-trivial reasoning.
-SPOT SPOT measures how well a system can detect critical errors in published papers, and ExpertLongBench expertlongbench evaluates the ability to generate long-form expert text without external search.
+ResearchBench Liu et al.
+(2025) and ResearcherBench Xu et al.
+(2025) evaluate, respectively, the ability to generate inspiration and hypotheses from papers, and to provide long-form advice on frontier AI research questions based on expert rubrics.
+Characterizing Deep Research Java et al.
+(2025) and LiveDRBench evaluate the recovery of correct claims in search tasks that require many information units and non-trivial reasoning.
+SPOT Son et al.
+(2025) measures how well a system can detect critical errors in published papers, and ExpertLongBench Ruan et al.
+(2025) evaluates the ability to generate long-form expert text without external search.
 These benchmarks finely probe partial abilities such as hypothesis generation, claim discovery, error detection, and expert writing, but do not directly evaluate the quality of full reports produced by web- or literature-based deep research agents.
 The third block targets benchmarks that evaluate the quality of deep research reports themselves.
-ReportBench ReportBench and DeepScholar-Bench DeepScholar-Bench measure report quality in academic survey and related-work settings using citation coverage, overlap with gold surveys, and verifiability of citations.
-DeepResearchBench deepbench and LiveResearchBench LiveResearchBench evaluate reports for multi-domain web-based deep research tasks using LLM-generated evaluation criteria (RACE), citation-based fact-checking (FACT), and dimensions such as coverage, depth, presentation, and citation accuracy.
-DeepResearchArena DeepResearchArena automatically constructs large numbers of research tasks from seminar transcripts, and measures factual alignment and overall quality via coverage/conflict/omission (KSR/KCR/KOR) against evidence keypoints extracted from cited URLs and LLM-generated task-specific rubrics (ACE).
-DeepResearchGym deepgym provides an offline search environment over a public web corpus and an LLM-as-judge protocol, offering a sandbox for evaluating deep research agents’ reports with respect to alignment with information needs, search faithfulness, and overall quality.
-ResearchRubrics ResearchRubrics provides 101 deep research prompts and over 2,500 human-written rubrics, and is designed so that an LLM-as-judge assigns per-item scores by following these detailed human criteria.
-Among these, we analyze DeepResearchBench, LiveResearchBench, and ResearchRubrics in more detail, as they are closest to DREX in terms of evaluation setting.
+ReportBench Li et al.
+(2025a) and DeepScholar-Bench Patel et al.
+(2025) measure report quality in academic survey and related-work settings using citation coverage, overlap with gold surveys, and verifiability of citations.
+DeepResearchBench Du et al.
+(2025) and LiveResearchBench Wang et al.
+(2025) evaluate reports for multi-domain web-based deep research tasks using LLM-generated evaluation criteria (RACE), citation-based fact-checking (FACT), and dimensions such as coverage, depth, presentation, and citation accuracy.
+DeepResearchArena Wan et al.
+(2025) automatically constructs large numbers of research tasks from seminar transcripts, and measures factual alignment and overall quality via coverage/conflict/omission (KSR/KCR/KOR) against evidence keypoints extracted from cited URLs and LLM-generated task-specific rubrics (ACE).
+DeepResearchGym Coelho et al.
+(2025) provides an offline search environment over a public web corpus and an LLM-as-judge protocol, offering a sandbox for evaluating deep research agents’ reports with respect to alignment with information needs, search faithfulness, and overall quality.
+ResearchRubrics Sharma et al.
+(2025) provides 101 deep research prompts and over 2,500 human-written rubrics, and is designed so that an LLM-as-judge assigns per-item scores by following these detailed human criteria.
+Among these, we analyze DeepResearchBench, LiveResearchBench, and ResearchRubrics in more detail, as they are closest to DEER in terms of evaluation setting.
 DeepResearchBench and LiveResearchBench are most similar to our web-based deep research report evaluation setting, but both rely heavily on LLMs to define evaluation criteria and assign scores, and their fact-checking is applied only to a limited subset of claims with explicit citations.
 ResearchRubrics has the advantage that domain experts write 20–40 rubric items for each deep research prompt, enabling fine-grained capture of what constitutes a “good” answer, but it does not perform source-level claim fact-checking, and the rubric structure differs across prompts, making it challenging to interpret system behavior consistently along shared dimensions and rubrics.
-In contrast, DREX extracts, as exhaustively as possible, all explicit and implicit claims in a report that can be grounded in evidence documents and performs claim-level fact-checking over this set, and evaluates reports using a shared ontology of 7 high-level areas, 25 criteria, and 130 rubric items derived from expert report evaluation norms and guidelines, augmented with task-specific expert guidance.
-This allows DREX to avoid designing a new rubric from scratch for each task while still applying a dense set of rubric items to each report, and to interpret system performance along standard dimensions and rubric items across tasks.
-Table 8: Comparison of Information Verification Pipelines.
-Benchmark Explicit Verif. Implicit Verif. Global Context Key Limitation vs. DREX
+In contrast, DEER extracts, as exhaustively as possible, all explicit and implicit claims in a report that can be grounded in evidence documents and performs claim-level fact-checking over this set, and evaluates reports using a shared ontology of 7 high-level areas, 25 criteria, and 101 rubric items derived from expert report evaluation norms and guidelines, augmented with task-specific expert guidance.
+This allows DEER to avoid designing a new rubric from scratch for each task while still applying a dense set of rubric items to each report, and to interpret system performance along standard dimensions and rubric items across tasks.
+Benchmark Explicit Verif. Implicit Verif. Global Context Key Limitation vs. DEER
+LiveDRBench
+Java et al. (2025)
+Relies on matching against static Ground Truth claims; lacks dynamic verification of citations against web sources.
 DeepResearch Bench
-deepbench
+Du et al. (2025)
 Ignores uncited sentences; fails to detect hallucinations in transitional logic.
 DeepResearchGym
-deepgym
+Coelho et al. (2025)
 Limited to explicitly cited claims; only calculates citation precision/recall metrics.
 ResearcherBench
-ResearcherBench
+Xu et al. (2025)
 Treats uncited claims simply as “ungrounded” (empty URL) without attempting context-based verification.
+DeepResearch Arena
+Wan et al. (2025)
+Evaluates Source $\to$ Report coverage (summarization) rather than Report $\to$ Source verification; misses uncited hallucinations.
+LiveResearchBench
+Wang et al. (2025)
+Checklist- and judge-based evaluation depends on task-specific annotations and LLM judgment; implicit claims and cross-sentence dependencies are not systematically enumerated or verified at the claim level.
+SPOT
+Son et al. (2025)
+$\triangle$ Evaluates internal error detection against static ground truth; penalizes valid but unannotated error predictions (false positives).
 ReportBench
-ReportBench
-$\triangle$ Verifies non-cited claims via external web search voting, ignoring internal document grounding.
+Li et al. (2025a)
+Verifies non-cited claims via external web search voting, ignoring internal document grounding.
 DeepScholar-Bench
-DeepScholar-Bench
-$\triangle$ Relies on physical distance (sliding window $w$); misses long-range semantic dependencies.
-DeepResearch Arena
-DeepResearchArena
-$\triangle$ Evaluates Source $\to$ Report coverage (summarization) rather than Report $\to$ Source verification; misses uncited hallucinations.
-LiveDRBench
-characterizing
-Relies on matching against static Ground Truth claims; lacks dynamic verification of citations against web sources.
-SPOT
-SPOT
-Evaluates internal error detection against static ground truth; penalizes valid but unannotated error predictions (false positives).
+Patel et al. (2025)
+Relies on physical distance (sliding window $w$); misses long-range semantic dependencies.
 […]
+Table 7: Comparison of Information Verification Pipelines.
 Appendix B Data Construction Details
 B.1 Topic Domain Analysis
-To configure Deep Research tasks (queries), we first analyzed 5045 in-house user queries collected from Deep Research services to ensure a balanced distribution of tasks across domains.
-The domain distribution of these queries is presented in Figure 4, from which we finalized 8 topic domains, including the top 7 areas and an "Others" category that groups the rest.
+Figure 4: Topic domains extracted from real-world Deep Research service logs
+To construct deep research tasks, we analyzed 5,842 in-house user queries collected from our deep research system to estimate real-world domain demand and to guide the benchmark’s target domain distribution.
+Figure 4 shows the resulting distribution.
+Based on this analysis, we finalized 11 topic domains: the 10 most frequent domains and an Others category that aggregates all remaining domains.
 B.2 HLE Subject Mapping
-Although actual Deep Research service logs are collected in a research-oriented usage context unlike general QA systems, since the majority of users are not limited to experts in specific academic fields, individual queries are not designed with the premise of generating expert-level academic reports or papers immediately.
-In addition, they include partial queries that depend on the ongoing session context or practical/pragmatic information needs, so using them as-is for evaluation research tasks may lead to a mismatch with the high-difficulty research tasks we aim for in scope and format.
-Accordingly, in this study, while maintaining the 8 topic domain composition derived from actual Deep Research logs, we defined subjects from Humanity’s Last Exam (HLE) hle that semantically correspond to each domain and used HLE items belonging to those subjects as seed tasks for each domain.
-At this time, only items within the difficulty range currently processable by LLMs among all HLE items were selected through a preliminary performance evaluation, and the selected items were remapped to the previously defined Deep Research topic domains to obtain a higher-difficulty academic task composition while preserving the actual usage distribution.
-The correspondence between each topic domain of the actual Deep Research service and HLE subjects is summarized in Table 9.
-B.3 Conversion from HLE QA to DeepResearch Reports
-Since the selected HLE items are initially configured in QA format, they are not suitable for deep report generation tasks as they stand.
-In this study, experts reviewed the question-answer-rationale of each item to identify the core concepts, theories, and phenomena covered, and then reconstructed them into research-oriented task queries suitable for the Deep Research usage context.
-At this time, deterministic elements such as specific answers, factual conclusions, or proofs presented in HLE were removed from the prompt, and instead, only high-level writing directions such as analysis scope, comparative perspectives, and issues to be discussed were presented to induce the model to develop the logical structure and conclusion on its own.
-In addition, if necessary, the basic structure of the output, such as "Introduction–Related Work–Method–Analysis–Discussion," was briefly specified to make the task more similar to the long-form research report generation expected in actual Deep Research use cases.
-Through this conversion process, simple answer-type QA items were transformed into a report-generation task set with HLE-level high-quality and high-difficulty problem structures, while following the actual Deep Research usage domain distribution.
-Figure 4: Topic domains extracted from real-world Deep Research service logs
 In-house Domain Humanity’s Last Exam Subject
@@
 Table 8: Mapping between in-house topic domains and Humanity’s Last Exam subjects. Combines "Software Development" and "Software" categories.
+Although our deep research service logs reflect research-oriented usage rather than general-purpose QA, most users are not domain experts, and their queries are typically not formulated as prompts for expert-level academic reports or papers.
+Moreover, the logs contain many context-dependent fragments (e.g., follow-up queries within an ongoing session) and pragmatic information needs.
+Therefore, using these queries directly as evaluation tasks would likely mismatch the high-difficulty, expert-level report-generation setting we target in both scope and format.
+Accordingly, we used Humanity’s Last Exam (HLE) Phan et al.
+(2025) as a source of expert-written, high-difficulty seed items that align with expert-level report generation.
+To preserve the 11-topic domain composition derived from actual deep research logs, we mapped each domain to one of HLE’s 13 subject domains and sampled HLE items from the corresponding subjects as domain-specific seeds.
+At this time, we filtered candidates via a preliminary performance evaluation, excluding items that were excessively difficult for LLM-based evaluation and retaining only those within an appropriate difficulty range.
+As a result, we selected a total of 50 seed items across HLE’s 13 subject domains.
+The correspondence between deep research topic domains and HLE subject domains is summarized in Table 8.
+B.3 Conversion from HLE QA to Deep Research Reports
+Because the selected HLE items are presented in a QA format, they are not directly suitable for deep research report-generation tasks in their original form.
+Accordingly, for each item, a domain expert reviewed the question, answer, and rationale to identify the underlying concepts, theories, and phenomena, and then reformulated it into a research-oriented task query appropriate for the deep research setting.
+During this reformulation, we removed answer-revealing elements from the prompt, such as specific answers, factual conclusions, and proofs, so that the model must derive the reasoning and conclusions on its own.
+When necessary, we also included writing guidance that constrains report development—such as the intended scope of analysis, comparative perspectives, and key issues to be addressed—to prevent uncontrolled drift and to enable more fine-grained evaluation of whether required elements are covered.
+Overall, this conversion reconstructs short-answer QA items into long-form report-generation tasks that require expert-level reasoning and exposition.
+Each task query was drafted by one domain expert and cross-reviewed by another expert from the same field.
+Cross-review was repeated in multiple rounds as needed to check whether the reformulated task (i) is not overly narrow, (ii) requires expert-level domain expertise, (iii) specifies a sufficiently concrete research scope and direction, and (iv) does not contain excessive hints that could steer the model toward the answer or conclusion; the task was revised accordingly.
+Experts were individuals with a master’s degree in the relevant field or equivalent domain expertise.
+An example research task is provided in Box 5.
 Example Research Task (MDP / Value Iteration) Write a research report that explores the conditions under which value iteration in Markov Decision Processes converges geometrically to the optimal value function.
@@
 The report should be structured as a rigorous technical investigation, integrating mathematical reasoning, conceptual explanations, and illustrative examples to support the analysis.
+Figure 5: Task Query Example
+B.4 Construction of Expert Evaluation Guidance
+For each task, we constructed Expert Evaluation Guidance to specify what an expert report for the given prompt must cover.
+The Guidance includes only the mandatory elements required by the topic (excluding optional content or stylistic preferences) and describes each element in as concrete and verifiable a form as possible so that compliance can be judged.
+The required elements are derived naturally from the task prompt and its writing requirements, and, when applicable, the Guidance also reflects key concepts implied by the underlying HLE item.
+In addition, because the prompt’s writing requirements (writing-direction instructions/additional requested constraints) are intended to keep the report from deviating from the intended direction and to enable fine-grained evaluation, they are also included as mandatory elements in the Guidance.
+For all 50 tasks, Expert Evaluation Guidance was written in the same expert workflow simultaneously with query reformulation, so that each task’s requirements and evaluation criteria are mutually aligned.
+As in query reformulation, each Guidance was drafted by one domain expert and cross-reviewed by another expert from the same field.
+During cross-review, we checked and revised whether (i) the topic’s mandatory content elements were included at a sufficiently detailed level and written in an evaluable way, (ii) optional content or stylistic preferences were not included as mandatory evaluation elements, and (iii) requirements specified in the task prompt’s writing requirements (writing-direction instructions/additional requested constraints) were reflected without omission in the Guidance.
+An example Expert Evaluation Guidance is provided in Box 6.
+Example Expert Evaluation Guidance (MDP / Value Iteration) 1. Formal MDP and Value Iteration Setup
+The report precisely defines the components of a finite-state MDP (state space, action space, transition probabilities, reward function, discount factor $\gamma$), states the Bellman optimality equation, and presents the value iteration update rule with consistent notation (e.g., $V_{k+1}=\mathcal{T}V_{k}$), ensuring all subsequent analysis refers unambiguously to this formal framework.
+2. Contraction Mapping and Norm Specification
+[…]
+3. Banach Fixed-Point Theorem Application
+The report invokes the Banach fixed-point theorem to establish existence, uniqueness, and geometric convergence of the optimal value function, clearly stating the theorem’s conditions (completeness of space, contraction property) and demonstrating how they are satisfied by the MDP and value iteration operator.
+4. Reward Boundedness and Sign Independence
+The report analyzes the role of reward boundedness in ensuring geometric convergence, demonstrating that bounded rewards (regardless of sign or symmetry) preserve contraction under the Bellman operator, and proving that unbounded rewards can violate the contraction condition; it explicitly states that the sign or symmetry of the reward interval (e.g., symmetric around zero or non-negative) does not affect convergence as long as boundedness and discounting are maintained.
+5. Geometric Convergence Rate Derivation
+[…]
+6. Edge Case Analysis for Reward Boundedness
+The report analyzes representative edge cases for reward boundedness, including symmetric vs.
+asymmetric intervals, zero rewards, and unbounded or improperly scaled rewards, using theoretical reasoning or controlled simulations to test whether geometric convergence holds or breaks down under each condition.
+7. Impact of Reward Sign and Symmetry
+The report evaluates whether the sign or symmetry of the reward interval (e.g., symmetric around zero vs.
+non-negative) affects convergence, proving that geometric convergence depends only on the discount factor and boundedness, not on sign, unless reward shaping alters the effective $\gamma$ or violates boundedness.
+8. Numerical Stability and Termination Criteria
+The report analyzes practical convergence behavior, including how finite precision arithmetic, reward scaling, and choice of stopping criterion (e.g., $V_{k+1}-V_{k}||_{\infty}<\epsilon$) interact with theoretical guarantees, and demonstrates at least one case where numerical error or poor scaling leads to premature termination or slow apparent convergence.
+9. Counterexamples for Non-Convergence
+The report constructs at least one verifiable scenario where value iteration fails to converge geometrically—such as when $\gamma=1$, rewards are unbounded, or the contraction condition is violated—and explains how this informs the boundary of the reward range for guaranteed convergence.
+10. Discount Factor and Reward Interaction
+The report examines how the interplay between $\gamma$ and reward magnitude influences the effective contraction rate, showing that while $\gamma$ controls the rate, reward scaling affects the constant factor in the convergence bound, and that rescaling rewards (e.g., dividing by $R_{\max}$) can normalize behavior across different domains.
+11. Illustrative Examples with Reproducible Design
+The report includes illustrative examples that feature explicit rules for MDP construction (number of states/actions, transition sparsity, reward assignment), initialization of $V_{0}$, value of $\gamma$, stopping threshold, and random seed control.
+These examples should be described with enough detail to make the reasoning transparent and verifiable.
+12. Limitations and Generalizations
+The report discusses limitations of the geometric convergence guarantee, including infinite state spaces, continuous actions, non-stationary environments, or non-linear function approximation, and clarifies whether the reward boundedness condition extends to policy iteration, Q-learning, or other dynamic programming variants, grounding each claim in earlier analysis.
+Figure 6: Example of Expert Evaluation Guidance
 Appendix C Deep Research Report Evaluation Taxonomy Details
+Dimensions Sub-dimensions
+Request Fulfillment Completeness, Scope, Helpfulness
+Analytical Soundness Quantification, Reasoning
+Structural Coherence Introduction, Body, Conclusion, Section
+Format & Style Report Format, Writing Quality, Paragraph Quality, Readability
+Ethics & Compliance Sensitive Handling, Safety & Impact, Perspective Balance
+Information Sufficiency Evidence Coverage, Claim Amount, Citation Amount, Reference Amount
+Information Integrity Claim Factuality, Citation Support, Reference Support, Reference Quality, Reference Diversity
+Table 9: Deep Research Report Evaluation Taxonomy: 7 Major Dimensions and 25 Sub-dimensions
 C.1 Dimension and Criteria Specification
-In this section, we detail the 7 dimensions and 26 detailed criteria of the Deep Research Report Evaluation Taxonomy presented in the main text.
-Request Completeness
-This dimension assesses how systematically and comprehensively the report addresses the user query with respect to the required elements, analysis scope, and final key message.
-Request Completeness assesses whether all requirements presented in the query are included without omission and developed with sufficient depth.
-Scope Boundary assesses whether the covered content, excluded items, assumptions, and limitations are clearly defined and consistently maintained.
-Value Validity assesses whether the final key message encompasses the entire query, is sufficiently specific and practical, and is logically consistent with the evidence and analysis results presented in the report body.
-Evidence Validity
-This dimension evaluates how accurate and valid the report’s figures and arguments are in terms of calculation, methodology, and logical development.
-Numeric Accuracy examines whether calculation processes, used formulas/statistical models, and indicators/units are presented without error, are appropriate for the problem context, and are expressed transparently enough for a third party to reproduce and verify.
-Logical Support evaluates whether the argument is developed consistently with the topic, necessary background/assumptions/inference steps are specified, and major claims and counter-argument responses are persuasively supported without leaps based on facts, data, and interpretations.
-Structure Consistency
-This dimension evaluates whether the introduction, body, and conclusion, and the structure of each section of the report, are organized consistently with the topic and scope.
+In this section, we detail the 7 dimensions and 25 sub-dimensions of the Deep Research Report Evaluation Taxonomy presented in the main text.
+Request Fulfillment
+Request Fulfillment evaluates whether the report meets the user’s request at a professional standard.
+It is assessed along three sub-dimensions: Completeness, Scope, and Helpfulness.
+Completeness evaluates whether all elements explicitly stated or implicitly required by the query are addressed without omission and with sufficient depth.
+Scope evaluates whether, in addressing these elements, the report clearly specifies what is included and excluded, as well as its assumptions, constraints, and limitations, and maintains these consistently throughout.
+Helpfulness evaluates whether the report materially advances the user’s goal by providing information that is sufficiently specific, practical, and actionable for direct use.
+Analytical Soundness
+Analytical Soundness evaluates how accurate and valid the report’s figures and arguments are in terms of calculation, methodology, and logical development.
+Quantification examines whether calculation processes, used formulas/statistical models, and indicators/units are presented without error, are appropriate for the problem context, and are expressed transparently enough for a third party to reproduce and verify.
+Reasoning evaluates whether the argument is developed consistently with the topic, necessary background/assumptions/inference steps are specified, and major claims and counter-argument responses are persuasively supported without leaps based on facts, data, and interpretations.
+Structural Coherence
+Structural Coherence evaluates whether the introduction, body, and conclusion, and the structure of each section of the report, are organized consistently with the topic and scope.
 Introduction checks whether it concisely and clearly presents the topic, problem, significance, and basic scope.
@@
 Conclusion checks whether it synthesizes the body content to complete the purpose of the introduction without introducing new claims or evidence.
-Section-Level checks whether each section supports the overall structure through clear organizational principles and appropriate connections.
-Narration Style
-This dimension assesses whether the report faithfully follows the required format and style and is expressed in a way that the reader can read and understand the content without difficulty.
-The report form checks whether external requirements, such as document length and section system, conform to professional report practices.
+Section checks whether each section supports the overall structure through clear organizational principles and appropriate connections.
+Format & Style
+Format & Style assesses whether the report faithfully follows the required format and style and is expressed in a way that the reader can read and understand the content without difficulty.
+Report Format checks whether external requirements, such as document length and section system, conform to professional report practices.
 Writing Quality checks whether sentences are concise and accurate, and whether terminology and tone are consistent.
-Paragraph checks whether paragraphs are sufficiently developed around a single point and naturally integrated with structural auxiliary elements.
-Reader Friendliness assesses whether subheadings and simple explanations/examples are used to guide the reader in following complex content.
-Ethics Compliance
-This dimension assesses whether the report is written ethically and responsibly regarding sensitive issues, potential harm, and perspective balance.
-Sensitive Issues checks whether sensitive topics such as politics, race, and gender are addressed with neutral, fair language and a balanced perspective.
+Paragraph Quality checks whether paragraphs are sufficiently developed around a single point and naturally integrated with structural auxiliary elements.
+Readability assesses whether subheadings and simple explanations/examples are used to guide the reader in following complex content.
+Ethics & Compliance
+Ethics & Compliance assesses whether the report is written ethically and responsibly regarding sensitive issues, potential harm, and perspective balance.
+Sensitive Handling checks whether sensitive topics such as politics, race, and gender are addressed with neutral, fair language and a balanced perspective.
 Safety & Impact checks whether negative impacts, side effects, and potential misuse of proposals/technologies/research results are adequately reviewed, and whether specific method presentations are dangerous.
@@
 Information Sufficiency
-This dimension assesses whether the information required to answer the request has been adequately secured and presented in terms of quantity and scope.
-Source Support evaluates whether verifiable evidence or reliable sources are provided to support claims that require external evidence.
-Information evaluates whether reliable facts and claims are sufficiently presented.
-Citations evaluate whether sufficient citations are made at necessary points.
-References are evaluated to determine whether the number of actually used references reaches an appropriate level.
+Information Sufficiency assesses whether the information required to answer the request has been adequately secured and presented in terms of quantity and scope.
+Evidence Coverage evaluates whether verifiable evidence or reliable sources are provided to support claims that require external evidence.
+Claim Amount evaluates whether reliable facts and claims are sufficiently presented.
+Citation Amount evaluates whether sufficient citations are made at necessary points.
+Reference Amount is evaluated to determine whether the number of actually used references reaches an appropriate level.
 Information Integrity
-This dimension assesses the factuality of the external information used in the report and the reliability and diversity of the citations/sources supporting it.
+Information Integrity assesses the factuality of the external information used in the report and the reliability and diversity of the citations/sources supporting it.
 Claim Factuality measures the proportion of verifiable claims that are determined to be factual.
-Citation Validity measures the proportion of citations that actually support the claim among citations attached to each claim.
-Reference Accuracy measures the proportion of sources that correctly support the argument among the presented/used references.
+Citation Support measures the proportion of citations that actually support the claim among citations attached to each claim.
+Reference Support measures the proportion of sources that correctly support the argument among the presented/used references.
 Reference Quality assesses whether the sources used are reproducible and reliable.
@@
 We extracted the common required elements from these materials and integrated/generalized overlapping or domain-specific items.
-To verify the validity of the drafted dimensions and detailed criteria, this study conducted cross-validation with 10 independent experts representing 6 domains (Computer Science, Artificial Intelligence, Biology, Chemistry, Mathematics, Psychology).
+To verify the validity of the drafted dimensions and sub-dimensions, this study conducted cross-validation with 10 independent experts representing 6 domains (Computer Science, Artificial Intelligence, Biology, Chemistry, Mathematics, Psychology).
 Two experts per domain reviewed the suitability of each item, focusing on factors considered necessary in light of domain practices, such as whether each dimension and criterion is actually a meaningful and necessary evaluation standard in that domain, whether the definition and scope of application are excessively ambiguous, and whether it unnecessarily overlaps with other criteria.
@@
 Appendix D Foundations of Deep Research Report Evaluation Taxonomy
-Request Completeness
-Request Completeness aligns with guideline-level completeness checks that evaluate whether all required components of a report are present according to prescribed inclusion mandates across authoritative standards Bastian2021PRISMA; Boutron2010CONSORT; ISO25010; WWC_Handbook; Percie2020ARRIVE; GRI_Standards; STROBE_2007.
-Scope Boundary reflects boundary-setting requirements that assess whether a report explicitly specifies its operative limits—such as eligibility conditions, assumptions, and constraints—within the structural fields defined in major evaluative frameworks Bastian2021PRISMA; NIST_AI_RMF; AHA2024Standards; AEA2024Data; OSF_Prereg; NASA_SE_Handbook; SEC_Reg_SK; SPIRIT_2013.
-Value Validity corresponds to coherence and utility standards that evaluate whether the report’s core conclusions not only align logically with the evidence base but also deliver sufficient specificity and practical feasibility to comprehensively address the user’s inquiry Guyatt2013GRADE; APA_JARS2025; NBER2024; IES_SEER; OECD_Policy_Eval.
-Evidence Validity
-Numeric Accuracy reflects evaluative requirements that check whether quantitative statements in a report correspond to verifiable computations, prespecified analytic procedures, and reproducible numerical evidence as established in major methodological frameworks Guyatt2013GRADE; JCGM_GUM; ACM_Badging; IUPAC_Green; IUPAP_Red; Econometric_Soc; AERA_APA_NCME_Standards; CODATA_2022; OECD_Quality; IFRS_Standards; ISO30414; IPCC_2019.
-Logical Support aligns with reasoning-assessment criteria that evaluate whether inferential steps are explicitly grounded in evidence, free of unsupported leaps, and consistent with theoretical or mathematical rigor EMS_Code2025; APA_GoodPractices; AERA2006Standards; AMS_Handbook; Fed_Rules_Evidence; COREQ_2007.
-Structure Consistency
-Introduction corresponds to structural-orientation requirements that assess whether a report’s opening section presents its scope and analytic pathway in accordance with prescribed organizational fields in established reporting and specification frameworks Bastian2021PRISMA; Boutron2010CONSORT; ISO29148_2018; IFRS_Integrated; CARE_2013.
-Body aligns with structural-progression criteria that evaluate whether the main sections develop the promised analytical sequence without omission or drift relative to the ordered components defined in recognized guideline structures Bastian2021PRISMA; Boutron2010CONSORT; ISO29148_2018; CF_Conventions; CARE_2013.
-Conclusion reflects closure-consistency requirements that examine whether final statements integrate preceding evidence without introducing unsupported expansions, consistent with the conclusion-governance rules embedded in major evaluative standards Bastian2021PRISMA; Boutron2010CONSORT; CARE_2013.
-Section-Level corresponds to intra- and inter-section coherence checks that assess alignment, ordering, and non-duplication of content in accordance with structured specification and reporting templates in authoritative frameworks Bastian2021PRISMA; ISO29148_2018; IFRS_Integrated.
-Narration & Style
-Report Form corresponds to format-governance requirements that evaluate whether a report adheres to prescribed structural conventions for scientific communication as codified in established editorial and reporting standards Bastian2021PRISMA; Boutron2010CONSORT; ICMJE2025; Leipzig_Rules; SIAM_Style; IPA_Handbook; APSA_Style; ASA_Style.
-Writing Quality aligns with language-precision criteria that assess clarity, specificity, and terminological consistency according to recognized guidelines for accurate and unbiased scholarly expression Boutron2010CONSORT; ICMJE2025; Chicago_Style; Blacks_Law; HBR_Guide.
-Paragraph reflects cohesion-assessment rules that examine whether individual paragraphs follow a coherent internal logic—anchoring topic statements, supporting evidence, and transitional structure—in line with authoritative reporting and specification frameworks Bastian2021PRISMA; ISO29148_2018; The_Bluebook; ASA_Style.
-Reader Friendliness corresponds to accessibility-oriented requirements that evaluate whether the narrative facilitates comprehension through appropriate signaling, explanatory devices, and presentation practices as outlined in major editorial and reporting guidelines Bastian2021PRISMA; ICMJE2025; HBR_Guide; GRI_Standards.
+Request Fulfillment
+Completeness aligns with guideline-level completeness checks that evaluate whether all required components of a report are present according to prescribed inclusion mandates across authoritative standards Bastian and Moher (2021); Boutron et al.
+(2010); International Organization for Standardization (2011); What Works Clearinghouse (2022); Percie du Sert et al.
+(2020); Global Reporting Initiative (2023); von Elm and others (2007).
+Scope reflects boundary-setting requirements that assess whether a report explicitly specifies its operative limits—such as eligibility conditions, assumptions, and constraints—within the structural fields defined in major evaluative frameworks Bastian and Moher (2021); National Institute of Standards and Technology (2023); American Historical Association (2024); American Economic Association (2024); Center for Open Science (2024); NASA (2016); U.S.
+Securities and Exchange Commission (2024); Chan and others (2013).
+Helpfulness corresponds to coherence and utility standards that evaluate whether the report’s core conclusions not only align logically with the evidence base but also deliver sufficient specificity and practical feasibility to comprehensively address the user’s inquiry Guyatt et al.
+(2013); American Psychological Association (2025); Chang et al.
+(2024); Institute of Education Sciences (2022); OECD (2020).
+Analytical Soundness
+Quantification reflects evaluative requirements that check whether quantitative statements in a report correspond to verifiable computations, prespecified analytic procedures, and reproducible numerical evidence as established in major methodological frameworks Guyatt et al.
+(2013); Joint Committee for Guides in Metrology (2008); Association for Computing Machinery (2025); International Union of Pure and Applied Chemistry (2007); International Union of Pure and Applied Physics (2010); The Econometric Society (2024); American Educational Research Association et al.
+(2014); Mohr and others (2024); OECD (2011); International Accounting Standards Board (2024); International Organization for Standardization (2018); Intergovernmental Panel on Climate Change (2019).
+Reasoning aligns with reasoning-assessment criteria that evaluate whether inferential steps are explicitly grounded in evidence, free of unsupported leaps, and consistent with theoretical or mathematical rigor European Mathematical Society (2025); American Philosophical Association (2024); American Educational Research Association (2006); American Mathematical Society (2022); U.S.
+Congress (2024); Tong et al.
+(2007).
+Structural Coherence
+Introduction corresponds to structural-orientation requirements that assess whether a report’s opening section presents its scope and analytic pathway in accordance with prescribed organizational fields in established reporting and specification frameworks Bastian and Moher (2021); Boutron et al.
+(2010); IEEE Computer Society (2018); IFRS Foundation (2021); Gagnier and others (2013).
+Body aligns with structural-progression criteria that evaluate whether the main sections develop the promised analytical sequence without omission or drift relative to the ordered components defined in recognized guideline structures Bastian and Moher (2021); Boutron et al.
+(2010); IEEE Computer Society (2018); Eaton and others (2024); Gagnier and others (2013).
+Conclusion reflects closure-consistency requirements that examine whether final statements integrate preceding evidence without introducing unsupported expansions, consistent with the conclusion-governance rules embedded in major evaluative standards Bastian and Moher (2021); Boutron et al.
+(2010); Gagnier and others (2013).
+Section-Level corresponds to intra- and inter-section coherence checks that assess alignment, ordering, and non-duplication of content in accordance with structured specification and reporting templates in authoritative frameworks Bastian and Moher (2021); IEEE Computer Society (2018); IFRS Foundation (2021).
+Format & Style
+Report Format corresponds to format-governance requirements that evaluate whether a report adheres to prescribed structural conventions for scientific communication as codified in established editorial and reporting standards Bastian and Moher (2021); Boutron et al.
+(2010); International Committee of Medical Journal Editors (2025); Comrie, B.
+and Haspelmath, M.
+and Bickel, B.
+(2015); Society for Industrial and Applied Mathematics (2024); International Phonetic Association (1999); American Political Science Association (2018); American Sociological Association (2022).
+Writing Quality aligns with language-precision criteria that assess clarity, specificity, and terminological consistency according to recognized guidelines for accurate and unbiased scholarly expression Boutron et al.
+(2010); International Committee of Medical Journal Editors (2025); University of Chicago Press (2024); Garner (2019, 2013).
+Paragraph Quality reflects cohesion-assessment rules that examine whether individual paragraphs follow a coherent internal logic—anchoring topic statements, supporting evidence, and transitional structure—in line with authoritative reporting and specification frameworks Bastian and Moher (2021); IEEE Computer Society (2018); Harvard Law Review Association (2020); American Sociological Association (2022).
+Readability corresponds to accessibility-oriented requirements that evaluate whether the narrative facilitates comprehension through appropriate signaling, explanatory devices, and presentation practices as outlined in major editorial and reporting guidelines Bastian and Moher (2021); International Committee of Medical Journal Editors (2025); Garner (2013); Global Reporting Initiative (2023).
 Ethics & Compliance
-Sensitive Issues corresponds to ethical-screening requirements that evaluate whether sensitive domains are addressed with neutrality, respect, and responsible language in accordance with established editorial and reporting ethics standards ICMJE2025; Bastian2021PRISMA; OHA_Principles; OAH_Standards; BPS_Code; Belmont_Report; APSA_Ethics.
-Safety & Impact aligns with harm-assessment provisions that examine whether potential adverse effects, misuse risks, or disproportionate impacts are identified and mitigated under recognized guidelines for responsible scientific communication ICMJE2025; Guyatt2013GRADE; NRC_Prudent; Helsinki_Decl; ESA_Code.
-Perspective Balance reflects fairness-evaluation criteria that assess whether multiple viewpoints and counterpositions are represented without undue bias following principles of impartiality articulated in authoritative publication and reporting frameworks ICMJE2025; Bastian2021PRISMA; BPA_GoodPractice; AAP_Code; CFA_Code; ABA_Model_Rules; AAPOR_Code.
+Sensitive Handling corresponds to ethical-screening requirements that evaluate whether sensitive domains are addressed with neutrality, respect, and responsible language in accordance with established editorial and reporting ethics standards International Committee of Medical Journal Editors (2025); Bastian and Moher (2021); Oral History Association (2024); Organization of American Historians (2018); British Psychological Society (2021); National Commission for the Protection of Human Subjects of Biomedical and Behavioral Research (1979); American Political Science Association (2012a).
+Safety & Impact aligns with harm-assessment provisions that examine whether potential adverse effects, misuse risks, or disproportionate impacts are identified and mitigated under recognized guidelines for responsible scientific communication International Committee of Medical Journal Editors (2025); Guyatt et al.
+(2013); National Research Council (2011); World Medical Association (2013); Ecological Society of America (2021).
+Perspective Balance reflects fairness-evaluation criteria that assess whether multiple viewpoints and counterpositions are represented without undue bias following principles of impartiality articulated in authoritative publication and reporting frameworks International Committee of Medical Journal Editors (2025); Bastian and Moher (2021); British Philosophical Association (2024); Australasian Association of Philosophy (2023); CFA Institute (2024); American Bar Association (2023); American Association for Public Opinion Research (2021).
 Information Sufficiency
-Source Support aligns with evidence-provision requirements that evaluate whether claims are accompanied by verifiable supporting sources in accordance with established evidentiary and reporting guidelines Bastian2021PRISMA; FORCE11_Data; NBER2024; ISSCR2025; GIPS_2020.
-Information corresponds to content-adequacy criteria that assess whether a report supplies all necessary factual and contextual material needed to justify its reasoning under recognized completeness and transparency standards Bastian2021PRISMA; ISO25010; EQUATOR_Network; LSA_Ethics2024; JOC_Guidelines; USGS_FSP; GRI_Standards.
-Citations align with source-attribution rules that examine whether in-text citations are provided at appropriate argumentative locations following authoritative norms for evidentiary traceability ICMJE2025; Wilkinson2016FAIR; IEEE2025Ethics; ACS2021Ethical; The_Bluebook.
-References are evaluated to determine whether the number of actually used references is appropriate and whether they collectively form a well-documented, traceable, and accessible evidence base StanfordAI2025; FORCE11_Data; DA_RT_Principles.
+Evidence Coverage aligns with evidence-provision requirements that evaluate whether claims are accompanied by verifiable supporting sources in accordance with established evidentiary and reporting guidelines Bastian and Moher (2021); Data Citation Synthesis Group (2014); Chang et al.
+(2024); International Society for Stem Cell Research (2025); CFA Institute (2020).
+Claim Amount corresponds to content-adequacy criteria that assess whether a report supplies all necessary factual and contextual material needed to justify its reasoning under recognized completeness and transparency standards Bastian and Moher (2021); International Organization for Standardization (2011); EQUATOR Network (2025); Linguistic Society of America (2024); The Journal of Organic Chemistry (2025); U.S.
+Geological Survey (2024); Global Reporting Initiative (2023).
+Citation Amount align with source-attribution rules that examine whether in-text citations are provided at appropriate argumentative locations following authoritative norms for evidentiary traceability International Committee of Medical Journal Editors (2025); Wilkinson et al.
+(2016); IEEE (2025); American Chemical Society (2021); Harvard Law Review Association (2020).
+Reference Amount are evaluated to determine whether the number of actually used references is appropriate and whether they collectively form a well-documented, traceable, and accessible evidence base Stanford HAI (2025); Data Citation Synthesis Group (2014); American Political Science Association (2012b).
 Information Integrity
-Claim Factuality reflects evidence-verification provisions that assess whether factual assertions in a report correspond to verifiable sources and documented evidence traces within established evaluative frameworks Wilkinson2016FAIR; APS2023Conduct; NIST_AI_RMF; PDG_Review; NIST_DLMF; SEC_Reg_SK.
-Citation Validity corresponds to source-justification checks that evaluate whether cited references substantively support the claims they accompany according to recognized standards for evidential accountability Wilkinson2016FAIR; ASA_Ethics; AHA2024Standards; Austin_Principles.
-Reference Accuracy aligns with source-quality and provenance requirements that examine whether referenced materials reliably and transparently support the arguments for which they are cited, in line with authoritative data-governance and reporting criteria Wilkinson2016FAIR; NBER2024; FORCE11_Data; APA_JARS2025; IFRS_Standards.
-Reference Quality evaluates whether the set of sources consists of reproducible, trustworthy, and methodologically sound information sources, ensuring that the report’s evidentiary foundation is grounded in reliable materials ACS2021Ethical; IEEE2025Ethics; SEP_Policy; Helsinki_Decl; Belmont_Report.
-Reference Diversity evaluates whether the evidence base avoids excessive concentration on a small number of sources and instead maintains a balanced level of source diversity, reducing the risk of narrow, biased, or selectively framed arguments EQUATOR_Network; LSA_Ethics2024; AAPOR_Code.
+Claim Factuality reflects evidence-verification provisions that assess whether factual assertions in a report correspond to verifiable sources and documented evidence traces within established evaluative frameworks Wilkinson et al.
+(2016); American Physical Society (2023); National Institute of Standards and Technology (2023); Navas and others (2024); National Institute of Standards and Technology (2024); U.S.
+Securities and Exchange Commission (2024).
+Citation Support corresponds to source-justification checks that evaluate whether cited references substantively support the claims they accompany according to recognized standards for evidential accountability Wilkinson et al.
+(2016); American Sociological Association (2018); American Historical Association (2024); Berez-Kroeker and others (2018).
+Reference Support aligns with source-quality and provenance requirements that examine whether referenced materials reliably and transparently support the arguments for which they are cited, in line with authoritative data-governance and reporting criteria Wilkinson et al.
+(2016); Chang et al.
+(2024); Data Citation Synthesis Group (2014); American Psychological Association (2025); International Accounting Standards Board (2024).
+Reference Quality evaluates whether the set of sources consists of reproducible, trustworthy, and methodologically sound information sources, ensuring that the report’s evidentiary foundation is grounded in reliable materials American Chemical Society (2021); IEEE (2025); Stanford Encyclopedia of Philosophy (2025); World Medical Association (2013); National Commission for the Protection of Human Subjects of Biomedical and Behavioral Research (1979).
+Reference Diversity evaluates whether the evidence base avoids excessive concentration on a small number of sources and instead maintains a balanced level of source diversity, reducing the risk of narrow, biased, or selectively framed arguments EQUATOR Network (2025); Linguistic Society of America (2024); American Association for Public Opinion Research (2021).
 To establish a robust and universally applicable Deep Research Report Evaluation Taxonomy, this framework comprehensively integrates authoritative standards and guidelines from a wide spectrum of disciplines, ranging from quantitative sciences and engineering to the humanities, social sciences, and professional fields such as law and finance (see Table 10).
@@
 Consequently, the resulting evaluation model ensures that the generated reports not only adhere to domain-specific best practices but also satisfy the fundamental requirements of validity, clarity, and professional integrity demanded by the global research and practitioner communities.
-Domain (80 listings) Reference
-Artificial Intelligence (8) NIST_AI_RMF; StanfordAI2025; ACL2024Guidelines; NeurIPS2025Guidelines; ICML2025Instructions; ICLR2024Guidelines; JMLR2024AuthorGuide; Springer2025MLJournal
-Biology (4) Bastian2021PRISMA; Boutron2010CONSORT; ISSCR2025; Percie2020ARRIVE
-Business (4) GRI_Standards; HBR_Guide; ISO30414; IFRS_Integrated
-Chemistry (4) ACS2021Ethical; IUPAC_Green; JOC_Guidelines; NRC_Prudent
-Computer Science (4) ISO25010; ACM_Badging; ISO29148_2018; Wilkinson2016FAIR
-Earth & Env. Science (4) IPCC_2019; USGS_FSP; CF_Conventions; ESA_Code
-Economics (4) NBER2024; AEA2024Data; Econometric_Soc; OECD_Quality
-Education (4) WWC_Handbook; AERA2006Standards; IES_SEER; AERA_APA_NCME_Standards
-Engineering (4) IEEE_Std2025; IEEE2025Ethics; NASA_SE_Handbook; ISO29148_2018
-Finance (4) IFRS_Standards; CFA_Code; GIPS_2020; SEC_Reg_SK
-History (4) AHA2024Standards; Chicago_Style; OHA_Principles; OAH_Standards
-Law (4) The_Bluebook; Blacks_Law; Fed_Rules_Evidence; ABA_Model_Rules
-Linguistics (4) LSA_Ethics2024; Leipzig_Rules; Austin_Principles; IPA_Handbook
-Mathematics (4) EMS_Code2025; AMS_Handbook; SIAM_Style; NIST_DLMF
-Medicine (4) Helsinki_Decl; STROBE_2007; CARE_2013; SPIRIT_2013
-Philosophy (4) APA_GoodPractices; SEP_Policy; BPA_GoodPractice; AAP_Code
-Physics (4) APS2023Conduct; IUPAP_Red; PDG_Review; CODATA_2022
-Political Science (4) DA_RT_Principles; APSA_Style; OECD_Policy_Eval; APSA_Ethics
-Psychology (4) APA_JARS2025; OSF_Prereg; BPS_Code; AERA_APA_NCME_Standards
-Sociology (4) AAPOR_Code; COREQ_2007; Belmont_Report; ASA_Style
+Domain (80) Reference
+AI (8) National Institute of Standards and Technology,2023; Stanford HAI,2025; Association for Computational Linguistics,2024; NeurIPS Foundation,2025; International Conference on Machine Learning,2025; International Conference on Learning Representations,2024; Journal of Machine Learning Research,2024; Springer Nature,2025
+Biology (4) Bastian and Moher,2021; Boutron et al.,2010; International Society for Stem Cell Research,2025; Percie du Sert et al.,2020
+Business (4) Global Reporting Initiative,2023; Garner,2013; International Organization for Standardization,2018; IFRS Foundation,2021
+Chemistry (4) American Chemical Society,2021; International Union of Pure and Applied Chemistry,2007; The Journal of Organic Chemistry,2025; National Research Council,2011
+Computer Science (4) International Organization for Standardization,2011; Association for Computing Machinery,2025; IEEE Computer Society,2018; Wilkinson et al.,2016
+Earth & Env. Science (4) Intergovernmental Panel on Climate Change,2019; U.S. Geological Survey,2024; Eaton and others,2024; Ecological Society of America,2021
+Economics (4) Chang et al.,2024; American Economic Association,2024; The Econometric Society,2024; OECD,2011
+Education (4) What Works Clearinghouse,2022; American Educational Research Association,2006; Institute of Education Sciences,2022; American Educational Research Association et al.,2014
+Engineering (4) IEEE Computer Society,2025; IEEE,2025; NASA,2016; IEEE Computer Society,2018
+Finance (4) International Accounting Standards Board,2024; CFA Institute,2024, 2020; U.S. Securities and Exchange Commission,2024
+History (4) American Historical Association,2024; University of Chicago Press,2024; Oral History Association,2024; Organization of American Historians,2018
+Law (4) Harvard Law Review Association,2020; Garner,2019; U.S. Congress,2024; American Bar Association,2023
+Linguistics (4) Linguistic Society of America,2024; Comrie, B. and Haspelmath, M. and Bickel, B.,2015; Berez-Kroeker and others,2018; International Phonetic Association,1999
+Mathematics (4) European Mathematical Society,2025; American Mathematical Society,2022; Society for Industrial and Applied Mathematics,2024; National Institute of Standards and Technology,2024
+Medicine (4) World Medical Association,2013; von Elm and others,2007; Gagnier and others,2013; Chan and others,2013
+Philosophy (4) American Philosophical Association,2024; Stanford Encyclopedia of Philosophy,2025; British Philosophical Association,2024; Australasian Association of Philosophy,2023
+Physics (4) American Physical Society,2023; International Union of Pure and Applied Physics,2010; Navas and others,2024; Mohr and others,2024
+Political Science (4) American Political Science Association,2012b, 2018; OECD,2020; American Political Science Association,2012a
+Psychology (4) American Psychological Association,2025; Center for Open Science,2024; British Psychological Society,2021; American Educational Research Association et al.,2014
+Sociology (4) American Association for Public Opinion Research,2021; Tong et al.,2007; National Commission for the Protection of Human Subjects of Biomedical and Behavioral Research,1979; American Sociological Association,2022
 Table 10: References by Domain used in Deep Research Report Evaluation Taxonomy. (Total 80 listings from 78 unique sources; 2 interdisciplinary standards are cross-listed).
@@
 In this appendix, we summarize the hierarchical structure and scoring method of the rubric used to evaluate the Deep Research report.
-Dimension / Criterion Element / Rubric description
+Dim / Sub-dim Criterion / Rubric items Description
 2.
-Evidence Validity └ 2.2 Logical Support 2.2.5 Element All major claims follow logically from the previously presented facts, data, interpretations, and assumptions, without skipped steps or unsupported leaps.
+Analytical Soundness └ 2.2 Reasoning 2.2.5 Criterion All major claims follow logically from the previously presented facts, data, interpretations, and assumptions, without skipped steps or unsupported leaps.
 [2.2.5.1 Coverage] All claims that require logical support are explicitly linked to the relevant facts, data, and interpretations (including the key evidence specified in the Expert Evaluation Guidance, EG), and no central claim is left without the necessary supporting evidence.
 [2.2.5.2 Quality] Each major claim is explained through clear, well-structured reasoning that shows how the underlying facts, data, and interpretations lead to that conclusion, making the logical connection explicit and easy for an expert reader to follow.
-Table 11: Example Coverage/Quality rubric factors for a reasoning-focused element under 2. Evidence Validity / 2.2 Logical Support.
+Table 11: Example Coverage/Quality rubric factors for a reasoning-focused element under 2. Analytical Soundness / 2.2 Reasoning.
 E.1 Hierarchical Levels
@@
 The 7 upper dimensions presented in Table 9 represent different evaluation perspectives of report quality.
-Level 2: Detailed Criteria (26)
-Detailed criteria that decompose each dimension into finer evaluation units; 2–5 detailed criteria are defined under each dimension.
+Level 2: sub-dimensions (25)
+Detailed criteria that decompose each dimension into finer evaluation units; 2–5 sub-dimensions are defined under each dimension.
 This forms an intermediate-level unit that specifies which aspects to examine for each dimension.
@@
 One element is subdivided into multiple factors from the perspectives of Coverage (C) and Quality (Q), and these factors serve as the minimum unit for actual scoring.
-Thus, with 7 dimensions, 26 detailed criteria, and 40 elements as the framework, each element is divided into Coverage/Quality factors to constitute a total of 130 specific check items.
+Thus, with 7 dimensions, 25 sub-dimensions, and 46 elements as the framework, each element is divided into Coverage/Quality factors to constitute a total of 101 specific check items.
 Table 11 shows an example of one element and its corresponding Coverage/Quality factors in the 2.
-Evidence Validity / 2.2 Logical Support dimension.
+Analytical Soundness / 2.2 Reasoning dimension.
 E.2 Coverage and Quality Factors
@@
 The Coverage factor examines whether the element has been covered without omission as required throughout the document, i.e., how faithfully it includes content scattered in multiple locations or composed of multiple detailed requirements.
-On the other hand, the Quality factor evaluates whether the description possesses sufficient depth, logic, and rigor based on professional report standards, assuming the parts already written, i.e., "how well it is written." By separating Coverage and Quality in this way, we can independently measure (1) what is missing and (2) what the level of the included content is in a long report.
-Both factors use a 1–10 point scale, and detailed criteria for scoring are presented in Table 12.
+On the other hand, the Quality factor evaluates whether the description possesses sufficient depth, logic, and rigor based on professional report standards, assuming the parts already written, i.e., “how well it is written.” By separating Coverage and Quality in this way, we can independently measure (1) what is missing and (2) what the level of the included content is in a long report.
+In our rubric, Level 4 consists of 101 factor-level items in total, comprising 66 Coverage items and 35 Quality items.
+Both factors use a 1–10 point scale, and sub-dimensions for scoring are presented in Table 12.
 Score Range Coverage Quality
@@
 F.1 Overview
-DREX’s Information Verification Protocol is based on a stepwise information verification procedure performed by human evaluators.
+DEER’s Information Verification Protocol is based on a stepwise information verification procedure performed by human evaluators.
 This section details the actual two-step procedure (Claim Extraction, Factual Accuracy Evaluation) performed by human evaluators.
@@
 Table 13 shows the definitions of the 6 claim types, and Table 14 shows examples of extracted sentences and classification results.
-Table 13: Information Verification ModuleA–F Claim Type Definitions and Examples.
-Types A–C are targets for external evidence verification, while Types D–E–F are classified as internal information or unverifiable areas.
 Type Definition and Example
@@
 F No Citation – Unknown Source: A claim requiring external evidence but for which no source is presented. Example: “These panels can last for 50 years without degradation.”
-Table 14: Example of human-annotated claims and corresponding LLM classification results.
+Table 13: Information Verification Module A–F Claim Type Definitions and Examples.
+Types A–C are targets for external evidence verification, while Types D–E–F are classified as internal information or unverifiable areas.
 Pos. Sentence / Extracted Claim Class (label)
@@
 L2.S4 These new panels are durable enough to withstand a Category 4 hurricane. F (No known source)
+Table 14: Example of human-annotated claims and corresponding LLM classification results.
 F.3 Step 2: Claim Verification
@@
 Detailed Verification Rubric.
-Fact verification was strictly performed according to the following detailed criteria:
+Fact verification was strictly performed according to the following sub-dimensions:
   • Supported: When the cited document clearly and directly includes the core facts (figures, causality, definitions, etc.) of the claim.
@@
   • Not Supported: When the basis for the claim cannot be found in the cited document, or the document is irrelevant to the topic.
-Cases where the document link is invalid (404) or the content cannot be verified are also included.
+  • Error: When the verification process fails due to accessibility issues (e.g., HTTP 4xx/5xx errors, Paywall, Captcha) or processing errors, preventing content verification.
+Source Reliability Check
+Independently of the content verification, we also evaluate the trustworthiness of the source domain itself.
+  • Reliable: Trustworthy sources such as academic journals, official statistics, and authoritative institutions.
+  • Unreliable: Sources with low credibility, such as personal blogs, social media posts, or unverified community forums.
 Inter-human Agreement.
@@
 Batch Extraction Strategy
-To mitigate the "Lost-in-the-Middle" phenomenon that occurs when processing long contexts and to increase cost efficiency, this study introduced a Batch Extraction strategy.
+To mitigate the “Lost-in-the-Middle” phenomenon that occurs when processing long contexts and to increase cost efficiency, this study introduced a Batch Extraction strategy.
 After dividing the entire report $D$ into sentence units $S=\{s_{1},s_{2},\dots,s_{m}\}$, they are processed in batches of a fixed size $B$ (in this study, $B=20$).
@@
 This method induces the model to focus on local sentences while remaining aware of the entire context, achieving human-level claim-extraction performance.
-Classification Performance.
-The results of performing type classification on 728 claims using GPT-5-mini are shown in Table 15.
-Explicit Citation (A) recorded a very high F1 score of 0.95, being identified almost perfectly.
-The distinction between Implicit Citation (BC) and Internal Information (DE) was somewhat challenging, but it recorded meaningful classification performance with F1 scores of 0.52 and 0.53, respectively.
-In particular, for the Unknown Source (F) type, which is the most difficult to identify, it achieved a precision of 0.35 and a recall of 0.25, indicating the ability to detect potential hallucination sentences.
-The overall accuracy was 0.68, which is a decent level considering the context of complex academic reports.
-Table 15: Claim Classification Performance (GPT-5-mini). Classes are merged into A (Explicit), BC (Implicit), DE (Internal/Structural), and F (Unknown).
-Class Precision Recall F1-Score Support
-A (Explicit) 0.92 1.00 0.95 272
-BC (Implicit) 0.42 0.68 0.52 163
-DE (Internal) 0.76 0.41 0.53 242
-F (Unknown) 0.35 0.25 0.30 51
-Overall 0.68 0.68 0.67 728
+Claim Extraction Evaluation Setup
+To rigorously evaluate the claim extraction performance, we employed an LLM-based Judge (GPT-5).
+Standard lexical metrics (e.g., ROUGE Lin (2004), Exact Match) are unsuitable for this task because the generated claims may differ in wording or granularity (e.g., one sentence split into multiple atomic claims) while preserving the same semantic meaning.
+We defined two key metrics: (1) Semantic Recall: Measures whether each ground truth claim is semantically covered by the set of extracted claims.
+The LLM Judge compares the ground truth claim against all candidate claims extracted from the same source sentence and determines if the core information is present.
+(2) Classification F1: Measures whether the LLM correctly classified the claim type for the extracted claims.
+The implementation code and LLM prompts used for the evaluation are included in the supplementary material.
 Comparison with Sliding Window Baseline.
@@
 This result demonstrates that the LLM’s “Evidence Position” prediction provides a precise pointer to the supporting evidence, which is crucial for efficient verification.
-Table 16: Baseline comparison on correctly predicted B/C claims ($N=131$). The GPT-5-mini method achieves the best balance of precision and accuracy (Jaccard).
 Method Jaccard Precision Recall
@@
 Sliding ($k=15$) 0.5627 0.5627 0.9341
+Table 15: Baseline comparison on correctly predicted B/C claims ($N=131$). Backtracking achieves the best balance of precision, recall, and Jaccard.
 G.3 Claim Verification
@@
 Finally, the citation set used for verification becomes $R^{\prime}_{i}=R_{i}\cup f_{backtrack}(c_{i})$.
-Relevant Context Filtering (Retrieval-Augmented Verification)
+Context Retrieval
 Using the entire report or long retrieved documents as input in the verification step is costly and can induce hallucinations due to unnecessary information (Noise).
-To solve this, we apply Relevant Context Filtering.
+To solve this, we apply Context Retrieval.
 For the verification target claim $q$ and the retrieved document $D_{retrieved}$, the document is divided into chunks $K=\{k_{1},k_{2},\dots,k_{r}\}$ (Size $\approx$ 1000 tokens).
 Then, the relevance score $Sim(q,k_{j})$ between each chunk and the claim is calculated, and only the top $N$ (Top-K) chunks are selected and used as input for the verification model.
-In this study, the BM25 robertson2009probabilistic algorithm was used as the default Selector: The selected chunk set $K_{selected}=\{k\in K\mid rank(Sim(q,k))\leq N\}$ is combined while maintaining the original document order to form the final context $C_{final}$.
+In this study, the BM25 Robertson et al.
+(2009) and OpenAI’s text-embedding-3-large OpenAI was used as the embedding model: The selected chunk set $K_{selected}=\{k\in K\mid rank(Sim(q,k))\leq N\}$ is combined while maintaining the original document order to form the final context $C_{final}$.
 $C_{final}=\text{Concatenate}(\text{SortByPosition}(K_{selected}))$
@@
 The LLM infers whether the claim’s core content and numerical information match the source to make a final judgment.
+Augmented Dataset and Robustness Evaluation
+While the human-annotated dataset provides high-quality ground truth, it exhibits significant class imbalance, with 82 “Supported” claims and only 5 “Not Supported” claims among 100 examples.
+This skew limits the ability to effectively evaluate the model’s capacity to discern unsupported claims and mitigate hallucinations.
+To address this, we constructed an adversarial augmented dataset.
+This dataset was generated by systematically perturbing initially supported claims—specifically by negating semantic meanings or altering numerical values—to create plausible but factually incorrect statements (i.e., “Not Supported”).
+All augmented examples were rigorously reviewed by human evaluators to ensure they are strictly false or unsupported by the source text.
+Ablation Study on Retrieval Parameters
+To identify the optimal configuration for the cost-efficient gpt-5-mini model, we conducted an ablation study varying Batch Size (10, 20), Reasoning Effort (Low, Medium, High), Retrieval Method (BM25 Robertson et al.
+(2009), OpenAI’s text-embedding-3-large OpenAI ), and Context Size (Top-K=2, 4).
+Table 16 summarizes the results on both the Original and Adversarial (Augmented) datasets.
+We observed that increasing the retrieved context size from Top-K=2 to 4 improved accuracy on the Original dataset (e.g., 77.0% $\rightarrow$ 79.3% for OpenAI, Low Effort), but increased the cost by approximately 35% (\$0.95 $\rightarrow$ \$1.28).
+The Low Reasoning Effort setting proved to be highly cost-effective, achieving comparable or superior performance to Medium/High effort while costing significantly less.
+Notably, the model demonstrated high robustness on the Adversarial dataset, maintaining high accuracy (>>88%) across most configurations, suggesting that it effectively distinguishes unsupported claims even under perturbations.
+Consequently, prioritizing feasibility of large-scale verification (cost/throughput) while retaining strong robustness, we selected the configuration Batch 20, Low Effort, Top-K=2, and OpenAI Embedding.
+This setup offers a minimal cost of \$0.95 per 1k claims while maintaining a strong accuracy of 77.0% (Original) and 88.5% (Adversarial), making it the most balanced choice for our resource-constrained high-volume verification pipeline.
+Original Adversarial
+Batch Effort Embed. Top-K Cost/1k (\$) Acc (%) F1 Acc (%) F1
+10 high BM25 4 3.61 77.01 87.25 87.36 85.16
+10 high OpenAI 4 3.67 78.16 88.00 88.46 86.62
+10 low BM25 4 0.91 79.31 88.00 87.36 85.16
+10 low OpenAI 4 1.03 78.16 88.00 89.01 87.34
+10 medium BM25 4 1.75 73.56 84.93 87.36 85.16
+10 medium OpenAI 4 1.62 73.56 84.93 87.91 85.90
+20 high BM25 2 3.08 79.31 88.16 88.46 86.79
+20 high BM25 4 3.38 79.31 88.74 89.56 88.05
+20 high OpenAI 2 3.34 77.01 87.25 85.71 83.33
+20 high OpenAI 4 3.31 75.86 86.49 88.46 86.79
+20 low BM25 2 1.02 73.56 84.93 90.11 88.61
+20 low BM25 4 1.30 81.61 90.20 89.01 87.34
+20 low OpenAI 2 0.95 77.01 87.25 88.46 86.79
+20 low OpenAI 4 1.28 79.31 88.74 89.56 87.90
+20 medium BM25 2 1.57 74.71 85.71 86.81 84.42
+20 medium BM25 4 1.79 79.31 88.74 88.46 86.62
+20 medium OpenAI 2 1.43 73.56 84.93 87.91 86.08
+20 medium OpenAI 4 1.71 74.71 85.71 88.46 86.79
+Table 16: Ablation study of GPT-5-mini on Context Retrieval setting.
+Showing the impact of Batch Size, Reasoning Effort, Embedding Method, and Top-K chunks on performance.
+Original refers to the standard dataset, and Adversarial refers to the augmented dataset.
 Example.
@@
 $\Rightarrow$ Final Result: Supported
-Fair Use and Direct Quote Detection
-To ensure compliance with copyright and research ethics, the appropriateness of direct quotes is evaluated.
-First, the similarity between sentences in the report and the original text (Reference) is calculated using RapidFuzz and SequenceMatcher.
-If the similarity ($\tau$) is 0.9 or higher, it is considered a direct quote.
-Sentences detected as direct quotes are primarily checked for quotation marks (" ") or block quote formatting using regular expressions.
-If the format is insufficient or additional judgment is needed, the LLM is presented with the sentence and the original text to comprehensively evaluate compliance with "Fair Use Guidelines" (citation indication, minimal quotation, etc.).
-The final score is calculated by deducting points from a maximum of 10 based on the non-compliant ratio among all sentences.
 G.4 Evaluation Metrics
@@
 Integrity Metrics
-  • External Claim Accuracy: The proportion of claims verified as factual among claims requiring external evidence (Type A, B, C).
+  • Claim Factuality: The proportion of claims verified as factual among claims requiring external evidence (Type A, B, C).
 […]
-  • Citation Accuracy: The proportion of citations that correctly support the corresponding claim among all citations.
-    $\text{Cit. Acc}=\frac{|\text{Supported Citations}|}{|\text{Total Citations}|}$
-  • Reference Accuracy (Supported per Shown): The proportion of references that actually contributed to content verification (Supported) among unique references shown in the report.
-    $\text{Ref. Acc}=\frac{|\text{Supported Unique References}|}{|\text{Total Unique References Shown}|}$
+  • Citation Support: The proportion of citations that correctly support the corresponding claim among all citations.
+    $\text{Score}=\frac{|\text{Supported Citations}|}{|\text{Total Citations}|}$
+  • Reference Support: The proportion of references that actually contributed to content verification (Supported) among unique references shown in the report.
+    $\text{Score}=\frac{|\text{Supported Unique References}|}{|\text{Total Unique References Shown}|}$
   • Reference Reproducibility: The proportion of references that were accessible during the actual verification process and for which the webpage in markdown format could be successfully retrieved using the Jina API (not Error).
-    $\text{Reproducibility}=1-\frac{|\text{Error References}|}{|\text{Used References}|}$
-  • Information Reliability: The proportion of references that are both reliable sources and support the content among the used references.
-    $\text{Reliability}=\frac{|\text{Reliable \& Supported References}|}{|\text{Used References}|}$
-  • Reference Diversity (Citations CV): Uses the Coefficient of Variation of citation counts to measure dependency on specific references. (Lower is better; inverted for scoring)
-    $\text{CV}=\frac{\sigma(\text{Citation Counts})}{\mu(\text{Citation Counts})}$
+    $\text{Score}=1-\frac{|\text{Error References}|}{|\text{Used References}|}$
+  • Reference Reliability: The proportion of references that are both reliable sources and support the content among the used references.
+    $\text{Score}=\frac{|\text{Reliable \& Supported References}|}{|\text{Used References}|}$
+  • Reference Diversity (Normalized HHI): Measures how evenly citations are distributed across used references using the Normalized Herfindahl-Hirschman Index Rhoades (1993).
+    $\text{Score}=10\times\left(1-\frac{\text{HHI}-1/N}{1-1/N}\right)$
 Sufficiency Metrics
-  • Externally Verifiable Claims Ratio: The proportion of claims verifiable with external evidence (Type A, B, C) among all claims.
-    $\text{Verifiable Ratio}=\frac{|\text{Claims (A, B, C)}|}{|\text{Total Claims}|}$
-  • Information Quantity: The total number of claims verified as factual (Supported). (Raw Count before normalization)
-    $\text{Info. Qty}=|\text{Accurate Verifiable Claims}|$
-  • Citation Quantity: The total number of valid citations (Supported Citation) supporting claims.
-    $\text{Cit. Qty}=|\text{Supported Citations}|$
-  • Reference Quantity: The total number of valid references (Supported Reference) supporting claims.
-    $\text{Ref. Qty}=|\text{Supported References}|$
+  • Evidence Coverage: The proportion of claims verifiable with external evidence (Type A, B, C) among all claims.
+    $\text{Score}=\frac{|\text{Claims (A, B, C)}|}{|\text{Total Claims}|}$
+  • Information Amount: The total number of claims verified as factual (Supported).
+    $\text{Score}=|\text{Accurate Verifiable Claims}|$
+  • Citation Amount: The total number of valid citations (Supported Citation) supporting claims.
+    $\text{Score}=|\text{Supported Citations}|$
+  • Reference Amount: The total number of valid references (Supported Reference) supporting claims.
+    $\text{Score}=|\text{Supported References}|$
+Final Score Calculation
+The final scores for Integrity and Sufficiency are computed by hierarchically aggregating the metrics (Metric $\rightarrow$ Criterion $\rightarrow$ Dimension), consistent with the score_avgs and criteria_avgs structure in the output.
+1. Normalization (Metric Level)
+Each raw metric value is first converted to a 0–10 scale:
+  • Ratio-based metrics (e.g., Factuality): Scaled linearly.
+    $\text{Score}=\min(\max(R,0),1)\times 10$
+  • Quantity-based metrics (e.g., Counts): Scored via step function with divisors $D$ (Info=15, Cit=10, Ref=4).
+    $\text{Score}=\min\left(\left\lfloor\frac{\max(N-1,0)}{D}\right\rfloor+1,10\right)$
+2. Aggregation
+  • Criterion Level: Average of normalized metric scores within each criterion.
+  • Dimension Level: Average of criterion scores within each dimension.
+Appendix H Baseline Model Details
+We use the following backbone model families in our experiments: Qwen3-235B, Gemini 2.5, Claude Opus 4.5, and GPT-5.
+For readability, we refer to the GPT family as GPT-5 in the paper, while the actual backbone used in our runs was GPT-5.2.
+Since Gemini 2.5 Pro includes reasoning by default, we use Gemini 2.5 Flash for the fast (non-reasoning) setting, and Gemini 2.5 Pro for the other settings.
+For think, we use each service’s default reasoning budget.
+For think+search, we do not build a custom retrieval pipeline; instead, we use the built-in web search system provided by each service.
+We exclude Gemini think+search because citation information is not provided in its outputs.
+For WebThinker, we use Qwen3-235B as an auxiliary model.
+For OpenAI Deep Research, we collect reports generated from the service environment as of August 2025.
+Appendix I Human-Correlation Experiment Setup
+We measure alignment between LLM judges and expert human judgments on 45 reports.
+We consider five domains and sample three tasks from each, yielding 15 tasks in total.
+For each task, we randomly sample three reports from those produced by five Deep Research systems—OpenAI Deep Research, Gemini 2.5 Pro Deep Research, Claude Opus 4.1 Deep Research, WebThinker, and Qwen3-235B Deep Research—resulting in $15\times 3=45$ reports overall.
+WebThinker Li et al.
+(2025b) uses Qwen3-235B as an auxiliary model.
+Each report was independently evaluated by two domain experts matched to the report’s topic, supporting both LLM–human correlation and human–human reliability analyses (90 ratings total; 45 reports × 2 experts).
+Experts had at least a master’s degree in a relevant field or comparable professional experience and, taking various factors into account, assigned an overall report quality score on a 1–5 scale (fractional values allowed).
+We used the mean of the two expert ratings as the report-level human score for LLM–human correlation, and we also reported agreement between the two experts’ ratings.
+Evaluating a report took 1.5 hours on average.
+Compensation followed the vendor’s standard payment framework, and we verified adherence to relevant procedures and policies.
+We measure correlation with human judgments for five evaluator models (GPT-5, GPT-5-mini, Claude Opus 4.1, Claude Sonnet 4.5, and Gemini 2.5 Pro).
+Each model evaluates the same set of reports under each evaluation setting (Vanilla, +Dimensions, +Granular Rubrics, +Expert Guidance).
+Under the +Dimensions setting, evaluation is performed at the level of five report-quality dimensions.
+Under +Granular Rubrics and +Expert Guidance, evaluation is performed down to the level of rubric items that further break down those dimensions.
+To measure alignment with expert human judgments, we report Pearson correlation ($r$), Spearman rank correlation ($\rho$), and pairwise agreement (PA).
+Each task corresponds to a single query and contains three reports.
+For each task, we compute Pearson $r$ and Spearman $\rho$ between human and model scores across the three reports, and report the mean over the 15 tasks.
+Following DeepBench Du et al.
+(2025), PA is the fraction of report pairs within a task for which the model’s relative preference matches the human relative preference; with three reports, each task has $\binom{3}{2}=3$ pairs.
+We compute PA per task and report the mean across tasks.
+To match the human 1–5 scale, we rescale model scores to 1–5 by dividing 1–10 scores by 2.
+For PA, a pair is counted as an agreement if the model and human judgments induce the same relation (>>, <<, or ==) for that pair.
+Tasks with undefined correlations (e.g., zero variance) are excluded from the corresponding averages.
+For each setting, we compute the three metrics ($r$, $\rho$, and PA) independently for each of the five evaluator models, and then report the average across the models.
+Appendix J Ethical Considerations and Potential Risks
+DEER is designed as a research evaluation benchmark rather than a deployment or decision-making system, and its scores should be interpreted accordingly.
+Because the framework partially relies on LLM-based judges, evaluation outcomes may reflect systematic biases or limitations of the underlying models; however, this risk is mitigated through expert-authored rubrics, task-specific expert guidance, and validation against human expert judgments.
+Unless otherwise specified, reported results correspond to single-run evaluations under fixed configurations, rather than aggregates over multiple random seeds.
+DEER evaluates the quality and evidence alignment of generated reports, not the real-world correctness, safety, or deployability of the underlying claims or recommendations, and thus should not be used as an endorsement for practical application in high-stakes domains.
+In addition, while the information verification module checks both explicit and implicit claims against external sources, its reliability is inherently bounded by the quality, availability, and temporal stability of retrieved web evidence.
+Overall, DEER is intended for comparative research analysis and diagnostic evaluation, and any use beyond this scope should be accompanied by appropriate human oversight and domain-specific safeguards.
+Appendix K Information About Use of AI Assistants
+AI assistants were used in a limited and supportive manner during the preparation of this work.
+Specifically, they were employed to assist with coding-related tasks (e.g., drafting boilerplate scripts, refactoring, and debugging) and with writing-related support, such as grammar checking, stylistic refinement, and clarity improvement.
+Importantly, all conceptual contributions, experimental design, analysis, interpretation of results, and final manuscript were carried out by the authors.

### v3

@@
 Abstract
-As large language models advance, deep research systems capable of generating expert-level reports through multi-step reasoning and evidence-based synthesis are emerging.
-However, evaluating such reports remains challenging.
-Existing benchmarks often lack systematic evaluation criteria, rely heavily on LLM-based judges that may miss issues requiring expert judgment, and verify only a limited subset of explicitly cited statements rather than report-wide factual reliability.
-To address these limitations, we introduce DEER, a benchmark for evaluating expert-level deep research reports.
-DEER comprises 50 report-writing tasks spanning 13 domains, along with an expert-grounded evaluation taxonomy with seven dimensions and 25 subdimensions, operationalized into 101 fine-grained rubric items.
-To improve evaluation consistency, DEER provides task-specific Expert Evaluation Guidance to support LLM-based judging.
-Complementing rubric-based assessment, we propose a document-level fact-checking architecture that verifies both cited and uncited claims and quantifies the quality and reliability of the supporting evidence.
-Experimental results show that DEER exhibits strong correlation with human expert judgments and yields interpretable diagnostics of system strengths and weaknesses.
-DEER: A Benchmark for Evaluating Deep Research Agents
-on Expert Report Generation
-Janghoon Han1, Heegyu Kim1, Changho Lee1, Dahm Lee1, Min Hyung Park1, Hosung Song1, Stanley Jungkyu Choi1, Moontae Lee1,2, Honglak Lee1,3 1LG AI Research, 2University of Illinois Chicago, 3University of Michigan, Ann Arbor Correspondence: hanjh9439@gmail.com
+Recent advances in large language models have enabled deep research systems that generate expert-level reports through multi-step reasoning and evidence-based synthesis.
+However, evaluating such reports remains challenging: report quality is multifaceted, making it difficult to determine what to assess and by what criteria; LLM-based judges may miss errors that require domain expertise to identify; and because deep research relies on retrieved evidence, report-wide claim verification is also necessary.
+To address these issues, we propose DEER, a benchmark for evaluating expert-level deep research reports.
+DEER systematizes evaluation criteria with an expert-developed taxonomy (7 dimensions, 25 subdimensions) operationalized as 101 fine-grained rubric items.
+We also provide task-specific Expert Evaluation Guidance to support LLM-based judging.
+Alongside rubric-based assessment, we propose a claim verification architecture that verifies both cited and uncited claims and quantifies evidence quality.
+Experiments show that while current deep research systems can produce structurally plausible reports that cite external evidence, there is room for improvement in fulfilling expert-level user requests and achieving logical completeness.
+Beyond simple performance comparisons, DEER makes system strengths and limitations interpretable and provides diagnostic signals for improvement.
+[^1] https://github.com/hanjanghoon/DEER.git
+Machine Learning, ICML
 1 Introduction
 Figure 1: Deep Research System Performance Comparison. Results for five systems on the proposed benchmark.
-Driven by rapid advances in large language models (LLMs), automated deep research systems are emerging as a core technology in both academia and industry OpenAI (2025); Google (2025); Anthropic (2025); Yang et al.
-(2025); Li et al.
-(2025b); Huang et al.
-(2025); Li et al.
-(2025c).
-Unlike traditional information retrieval systems, these systems address complex research queries by decomposing them into multiple steps and dynamically seeking additional information based on intermediate results.
-Through this process, they integrate information from diverse sources and synthesize multiple perspectives to produce reliable, evidence-based research reports Xu and Peng (2025); Zhang et al.
-(2025); Java et al.
-(2025).
-As a result, deep research systems demonstrate strong performance even on challenging benchmark tasks.
-Early evaluations of deep research systems relied primarily on complex reasoning benchmarks Rein et al.
-(2023); Phan et al.
-(2025), which indirectly assessed information-gathering, hypothesis testing, and multi-step reasoning through task performance.
-Subsequently, deep web search QA benchmarks were introduced to more directly measure systems’ web browsing and information retrieval abilities—core capabilities of deep research—by evaluating multi-step search, information integration, and answer derivation Wei et al.
-(2025a); Krishna et al.
-(2025); Mialon et al.
-(2023); Chen et al.
-(2025); Gou et al.
-(2025).
-More recently, deep research report benchmarks have emerged to evaluate the quality of generated reports from multiple perspectives, moving beyond simple short-answer-based evaluation Li et al.
-(2025b); Du et al.
-(2025); Consult (2025); Coelho et al.
-(2025).
+Driven by rapid advances in large language models (LLMs), automated deep research systems are emerging as a core technology in both academia and industry (OpenAI, 2025b; Google, 2025; Anthropic, 2025; Yang et al., 2025; Li et al., 2025b; Huang et al., 2025; Li et al., 2025c).
+Unlike conventional web search, these systems address complex research queries by decomposing them into multiple steps and dynamically seeking additional information based on intermediate results.
+Through this process, they integrate information from diverse sources and synthesize multiple perspectives to produce reliable, evidence-based research reports (Xu and Peng, 2025; Zhang et al., 2025; Java et al., 2025).
+As a result, deep research systems can achieve strong performance even on challenging benchmark tasks (Mialon et al., 2023; Phan et al., 2025).
+Early evaluations of deep research systems relied primarily on complex reasoning benchmarks (Rein et al., 2023; Mialon et al., 2023; Phan et al., 2025), which indirectly assessed information gathering, hypothesis testing, and multi-step reasoning through task performance.
+Subsequently, deep web search QA benchmarks were introduced to more directly measure systems’ web browsing and information retrieval abilities—core capabilities of deep research—by evaluating multi-step search, information integration, and answer derivation (Wei et al., 2025a; Krishna et al., 2025; Mialon et al., 2023; Chen et al., 2025; Gou et al., 2025).
+More recently, deep research report benchmarks have emerged to evaluate the quality of generated reports from multiple perspectives, moving beyond simple short-answer-based evaluation (Consult, 2025; Coelho et al., 2025; Du et al., 2025; Wan et al., 2025).
 Despite these advancements, existing methodologies for evaluating deep research systems continue to face important limitations when applied to expert-level reports.
-First, although prior work has proposed high-level dimensions for report evaluation, these frameworks do not comprehensively specify the diverse elements required for expert reporting.
-In particular, many approaches Du et al.
-(2025); Patel et al.
-(2025); Wang et al.
-(2025); Wan et al.
-(2025) rely on LLMs rather than human experts to define detailed subdimensions and concrete evaluation items, which can introduce inconsistencies and reduce evaluation reliability.
-Second, even when a fixed rubric is provided, evaluations that rely on LLM judges may fail to identify issues that require expert judgment.
-Third, current approaches to source verification are typically restricted to sections with explicit citation markers, leaving the factual reliability of claims across the full report insufficiently examined.
+First, evaluation criteria are often underspecified, leaving it unclear what aspects of report quality should be assessed.
+Prior benchmarks typically evaluate reports using coarse, high-level dimensions, which do not provide sufficiently fine-grained criteria for precise assessment of report quality (Consult, 2025; Coelho et al., 2025).
+Moreover, even when fine-grained evaluation items are introduced, they are often generated or structured by LLMs, which can undermine consistency and reliability (Du et al., 2025; Wan et al., 2025; Wang et al., 2025).
+Second, even with well-specified rubric items, evaluations that rely on LLM judges may fail to identify issues that require domain expertise.
+Third, current approaches to source verification are typically restricted to claims with explicit citation markers, leaving factual reliability across the full report insufficiently examined.
 Together, these limitations hinder comprehensive and reliable evaluation of deep research systems.
-To address these limitations, we propose DEER (the DEep research Expert Report benchmark), which provides comprehensive and reliable evaluation of deep research reports through 50 report-generation tasks spanning 13 domains.
-DEER analyzes a broad range of expert report assessment criteria and consolidates them via expert consensus into a Deep Research Report Evaluation Taxonomy comprising 7 dimensions and 25 subdimensions.
-Based on this taxonomy, DEER evaluates each report across two complementary components: (i) report-quality dimensions that require holistic, document-level judgment, and (ii) external-information dimensions that can be assessed at the claim level by fact verification.
-For report quality, we operationalize the taxonomy into a fixed set of 101 fine-grained rubric items and supplement these with report-specific Expert Evaluation Guidance authored by domain experts to improve the consistency and validity of LLM-based scoring.
-For external information, we introduce a fact-verification module that examines both cited and uncited claims across the report and produces quantitative measures of evidence quality.
-Finally, DEER integrates rubric-based LLM evaluations with the resulting metric-based signals to yield a unified, multi-dimensional assessment of each expert report.
-Figure 1 presents an overview of dimension-level performance for major deep research systems evaluated using DEER.
+To address these limitations, we propose DEER (the DEep research Expert Report benchmark), which evaluates deep research reports through 50 report-generation tasks spanning 13 domains.
+DEER surveys established reporting norms and evaluation criteria across domains and synthesizes them, through an expert consensus process, into a Deep Research Report Evaluation Taxonomy comprising 7 dimensions and 25 subdimensions.
+Based on this taxonomy, DEER evaluates each report across two complementary components: (i) report-quality assessment, which requires holistic, document-level judgment, and (ii) external-information verification, which can be assessed at the claim level by checking against external sources.
+For report quality, we operationalize the taxonomy into a fixed set of 101 fine-grained rubric items and supplement them with task-specific evaluation guidance authored by domain experts to improve the consistency and validity of LLM-based scoring.
+For external information, we introduce an information-verification module that examines both cited and uncited claims across the report and produces quantitative measures of evidence quality.
+DEER then integrates rubric-based scores with these metric-based signals to yield a unified, multidimensional assessment of each expert report.
+Figure 1 presents an overview of dimension-level performance across deep research systems evaluated using DEER.
 The main contributions of this work are as follows:
   • We present DEER, a systematic and interpretable benchmark for evaluating deep research reports, grounded in a hierarchical evaluation taxonomy.
-  • We operationalize this taxonomy into a fixed set of 101 fine-grained rubric items and provide report-specific Expert Evaluation Guidance, improving the consistency and reliability of LLM-based evaluation.
-  • We propose a report-level fact verification architecture that backtracks inter-claim dependencies to retrieve and verify relevant citations, enabling more complete evaluation of factual reliability across full reports.
+  • We translate the taxonomy into a standardized set of 101 fine-grained rubric items and provide task-specific Expert Evaluation Guidance to support consistent and reliable LLM-based report scoring.
+  • We propose a report-level information-verification architecture that backtracks inter-claim dependencies to retrieve citations and verify claims, enabling more complete evaluation of claim reliability across full reports.
 Figure 2: Overview of the DEER evaluation framework.
@@
 2 Related Works
-With the proliferation of high-performing LLMs, LLM-as-a-Judge approaches—using LLMs as evaluators—have been widely proposed and studied Zheng et al.
-(2023); Liu et al.
-(2023b); Chiang et al.
-(2024); Kim et al.
-(2024a, b).
-In this line of work, early evaluations for deep research primarily focused on how well models solve expert-level, high-difficulty questions Phan et al.
-(2025); Rein et al.
-(2023).
-Subsequently, to directly assess a core property of deep research—accessing and leveraging external information from the open web—benchmarks were proposed to measure models’ ability to browse/search the web, synthesize the required information, and construct answers grounded in that evidence Gou et al.
-(2025); Wei et al.
-(2025a); Huang et al.
-(2025).
-More recently, beyond simple short-answer evaluation, studies have sought to evaluate the ability to generate long-form reports that require expert-level analysis, reasoning, and interpretation by integrating information from multiple sources Consult (2025); Coelho et al.
-(2025); Du et al.
-(2025).
+With the proliferation of high-performing LLMs, LLM-as-a-Judge approaches—using LLMs as evaluators—have been widely proposed and studied (Zheng et al., 2023; Liu et al., 2023b; Chiang et al., 2024; Kim et al., 2024a, b).
+In this line of work, early evaluations for deep research primarily focused on how well models solve expert-level, high-difficulty questions (Rein et al., 2023; Phan et al., 2025).
+Subsequently, to directly assess a core property of deep research—accessing and leveraging external information from the open web—benchmarks were proposed to measure models’ ability to browse/search the web, synthesize the required information, and construct answers grounded in that evidence (Gou et al., 2025; Wei et al., 2025a; Huang et al., 2025).
+More recently, beyond simple short-answer evaluation, studies have sought to evaluate the ability to generate long-form reports that require expert-level analysis, reasoning, and interpretation by integrating information from multiple sources (Consult, 2025; Coelho et al., 2025; Du et al., 2025; Wan et al., 2025; Wang et al., 2025; Yao et al., 2025; Sharma et al., 2025).
 Despite this progress, existing report-evaluation methods are not based on expert-defined, fine-grained criteria.
-Some approaches rely on only a few coarse axes, leaving judgments largely to the evaluator LLM’s implicit standards Consult (2025); Coelho et al.
-(2025).
-Even when rubrics are subdivided, their specification and application can still depend substantially on the evaluator LLM Du et al.
-(2025), making it unclear whether the resulting scores correctly reflect report quality or align with expert assessment.
-In addition, source verification is often omitted Consult (2025); Coelho et al.
-(2025) or limited to a subset of cited claims Du et al.
-(2025), which may be insufficient to assess report-level factuality.
+Some approaches rely on only a few coarse axes, leaving judgments largely to the evaluator LLM’s implicit standards (Consult, 2025; Coelho et al., 2025).
+Even when rubrics are subdivided, their specification and application can still depend substantially on the evaluator LLM (Du et al., 2025; Wan et al., 2025), making it unclear whether the resulting scores correctly reflect report quality or align with expert assessment.
+In addition, source verification is often omitted (Consult, 2025; Coelho et al., 2025) or limited to a subset of cited claims (Du et al., 2025; Wan et al., 2025), which may be insufficient to assess report-level factuality.
 To address these limitations, we propose an evaluation framework that (i) scores long-form research reports along multiple dimensions using expert-systematized, fine-grained criteria, and (ii) performs systematic source verification for claims throughout the report.
@@
 For topic classification, we adopt the taxonomy of Wettig et al.
-(2025), consistent with prior work Du et al.
-(2025).
-We use Humanity’s Last Exam (HLE) Phan et al.
-(2025) as the source of seed questions because it provides expert-written, high-difficulty, multidisciplinary questions that align well with the expert-level topics addressed in deep research reports.
+(2025), consistent with prior work (Du et al., 2025).
+We use Humanity’s Last Exam (HLE) (Phan et al., 2025) as the source of seed questions because it provides expert-written, high-difficulty, multidisciplinary questions that align well with the expert-level topics addressed in deep research reports.
 To match our target topic distribution, we map our topics to the 13 HLE subject domains and sample 50 seed questions accordingly.
@@
 4.1 Overview
-To systematically evaluate deep research reports, we establish an evaluation framework grounded in two core aspects of deep research: external evidence acquisition and report-level synthesis Java et al.
-(2025); Xu and Peng (2025); Zhang et al.
-(2025).
+To systematically evaluate deep research reports, we establish an evaluation framework grounded in two core aspects of deep research: external evidence acquisition and report-level synthesis (Java et al., 2025; Xu and Peng, 2025; Zhang et al., 2025).
 Based on these aspects, we construct a Deep Research Report Evaluation Taxonomy comprising seven major dimensions (§ 4.2), grouped into five report-quality dimensions and two external-information dimensions.
@@
 First, they typically apply broad, high-level criteria, granting LLM judges substantial discretion over what aspects of a report to attend to.
-As a result, judges may (i) examine only parts of a report, (ii) consider only a subset of the many sub-aspects implied by each criterion, and (iii) focus on non-critical or superficial elements, leading to increased variance across judges and reduced evaluation consistency Li et al.
-(2025b); Consult (2025); Coelho et al.
-(2025).
+As a result, judges may (i) examine only parts of a report, (ii) consider only a subset of the many sub-aspects implied by each criterion, and (iii) focus on non-critical or superficial elements, leading to increased variance across judges and reduced evaluation consistency (Li et al., 2025b; Consult, 2025; Coelho et al., 2025).
 Second, even when the evaluation focus is clearly defined, assessing correctness and completeness often requires domain expertise that LLM judges may lack.
-Consequently, they may miss subtle but important issues, including non-obvious logical leaps, domain-specific misinterpretations, and fine-grained inaccuracies Du et al.
-(2025).
+Consequently, they may miss subtle but important issues, including non-obvious logical leaps, domain-specific misinterpretations, and fine-grained inaccuracies (Du et al., 2025).
 To address these limitations, this study (1) makes evaluation criteria explicit through a fixed set of fine-grained rubric items and (2) provides task-specific Expert Evaluation Guidance that includes domain-specific context and reference points, enabling LLM judges to identify hard-to-detect errors and omissions.
@@
 Broad evaluation criteria can lead to inconsistent LLM-based scoring.
-Recent work addresses this issue by decomposing high-level criteria into more granular rubrics Lee et al.
-(2025); Wei et al.
-(2025b); Ruan et al.
-(2025).
+Recent work addresses this issue by decomposing high-level criteria into more granular rubrics (Lee et al., 2025; Wei et al., 2025b; Ruan et al., 2025).
 For deep research report evaluation, Du et al.
 (2025) follows a related approach by using an LLM to generate task-specific evaluation criteria.
-While this strategy narrows the evaluation focus, it raises two important concerns: whether the generated criteria are sufficiently comprehensive and capture what matters most in expert reports, and whether dynamically varying criteria enable interpretable and comparable scores that support systematic diagnosis of system strengths and weaknesses across tasks.
+While this strategy narrows the evaluation focus, it raises two important concerns: whether the set of generated criteria is sufficiently comprehensive and captures what matters most in expert reports, and whether dynamically varying criteria enable interpretable and comparable scores that support systematic diagnosis of system strengths and weaknesses across tasks.
 This study uses a fixed, expert-designed rubric to ensure reliable and interpretable evaluation.
@@
 [^2] The 46 criteria were validated through independent review by 10 experts following the procedure in § 4.2.
-The rubric is applied identically to all deep research reports across 50 tasks, enabling diagnosis of the system’s overall strengths and weaknesses with richer, more interpretable signals (see Appendix E for the rubric structure and examples).
+The rubric is applied identically to all deep research reports across 50 tasks, enabling diagnosis of the system’s overall strengths and weaknesses with richer, more interpretable signals (see Appendix D for the rubric structure and examples).
 Expert Evaluation Guidance.
@@
 Through this process, Expert Evaluation Guidance complements the fixed rubric by anchoring evaluation in task-specific expert knowledge while maintaining consistency and interpretability across tasks.
-Appendix B provides detailed guidance construction procedures and illustrative examples.
-Model Req. Ful. Analyt. Sound. Struct. Coh. Format & Style Info. Int. Info. Suff. Ethics Mean
+Appendix B.4 provides detailed guidance construction procedures and illustrative examples.
+Model Request. FulFill. Analyt. Sound. Struct. Cohere. Format & Style Inform Int. Inform. Suff. Ethics Mean
 General LLMs
@@
 Deep Research
-WebThinker 4.11 4.64 5.51 7.35 6.21 6.40 7.13 5.91
+WebThinker (Li et al., 2025b) 4.11 4.64 5.51 7.35 6.21 6.40 7.13 5.91
 Qwen3-235B (deep) 4.13 4.69 4.85 7.06 6.55 7.90 7.43 6.09
@@
 OpenAI (deep) 4.67 5.29 6.28 7.66 7.14 6.89 7.48 6.49
-Table 1: Baseline models’ expert report evaluation results (updated). In each column, the highest score is marked in bold and the second highest score is underlined.
+Table 1: Evaluation results for expert reports generated by baseline methods. The best score in each column is shown in bold, and the second highest score is underlined.
 4.4 Information Verification
-To automate information verification in expert reports, we must overcome the limitations of existing systems that focus solely on explicit citations.
-In reality, many sentences implicitly inherit preceding evidence or construct new claims through contextual connections.
-Excluding these implicit claims from verification makes the reliability assessment of the entire document incomplete.
-Our Information Verification Module proposes an architecture that extends the verification scope to the entire document by tracking sentence-level logical flows.
-Implicit Claim & Back-Tracking
-This module analyzes contextual correlations to convert even uncited sentences into verifiable units.
-Specifically, when sentence $s_{i}$ implicitly cites the content of a previous sentence or paragraph, we use a Back-Tracking mechanism to trace the reference source.
-To this end, the LLM identifies the set of preceding sentences $R(s_{i})$ that $s_{i}$ depends on, and the valid citation set $\mathcal{V}(s_{i})$ used for actual verification is defined as follows:
-$\mathcal{V}(s_{i})=\mathcal{C}(s_{i})\cup\bigcup_{k\in R(s_{i})}\mathcal{C}(s_{k})$ (1)
-where $\mathcal{C}(s_{i})$ is the explicit citation of $s_{i}$.
-This structure allows verification by inheriting citations from referenced preceding sentences even if the sentence itself has no citation ($\mathcal{C}(s_{i})=\emptyset$).
-This is a key element that substantially expands the verification scope in policy and research reports containing many implicit claims.
-Batch Processing for Claim Extraction
-When processing long documents as a single input, LLMs suffer from the Lost-in-the-Middle Liu et al.
-(2023a) problem, which leads to the omission of important claims.
-To address this, we apply a Batch Extraction strategy that maintains the context of the entire document $D$ while parallel-processing only the target sentences divided into small batches $B_{j}$:
-$\mathcal{A}=\bigcup_{j=1}^{M}E(D,B_{j}),\quad D=\bigcup_{j=1}^{M}B_{j}$ (2)
-This increases the model’s focus to prevent missing important claims, and experimental results showed a significant improvement in extraction recall compared to the single processing ($E(D,D)$) method.
-Although the number of API calls increases, we confirmed that this is an essential strategy for securing high coverage in reports dealing with hundreds of claims.
-Scalable Verification Architecture
-The extracted claims are passed to the fact verification stage, where the biggest technical challenge is the trade-off between accuracy and cost.
-Verifying every claim individually yields high accuracy, but cost increases linearly with long documents.
-To address this, we introduce two scaling strategies.
-First, we apply Grouped Claim Verification, which bundles multiple claims citing the same source into a single input for verification.
-Experimental results showed that increasing Group Size from 1 to 20 raised F1 from 87.25 to 91.61 and reduced cost per 1K claims by 72%.
-This indicates that grouped input improves understanding the overall structure of the document and verifying claims, while reducing verification overhead.
-Second, greater API cost reduction is possible through the Context Retrieval strategy, which selectively provides only the context necessary for fact verification.
-This method achieved an API cost reduction of approximately 72%, but a 3.4% drop in F1 was observed, indicating a clear trade-off.
-This hierarchical pipeline (Batch Extraction $\rightarrow$ Grouped Verification $\rightarrow$ Context Retrieval) is designed to achieve both high-precision verification and cost efficiency even in long reports containing hundreds of claims.
+To evaluate the external-information dimensions (Information Integrity and Information Sufficiency) in a reproducible way, we verify claims against evidence across the entire report and summarize the results as quantitative metrics, rather than relying on free-form LLM scoring.
+Specifically, our module (i) extracts atomic claims and links each verifiable claim to the sources it should be checked against (including uncited claims via implicit-citation recovery), and (ii) verifies whether the linked evidence supports each claim and aggregates the outcomes into Integrity/Sufficiency metrics (Appendix F.4).
+Claim Types and Verification Scope.
+Existing citation checkers typically verify only explicitly cited sentences, restricting verification to a narrow subset of report content.
+Following the human protocol (Appendix E), we classify extracted claims into six types (A–F) and focus external evidence verification on verifiable claim types (A–C) (Appendix F).
+This typing makes the verification scope explicit: (A) claims with inline citations; (B/C) uncited claims whose support is available elsewhere in the report; and (D–F) claims that are structural, non-verifiable, or lack identifiable support.
+Implicit Claim Back-Tracking.
+A key challenge is that expert reports often contain implicit claims whose supporting citations appear in earlier sentences (or even earlier sections), so the claim itself has no explicit marker.
+To verify such claims, we introduce a semantic Back-Tracking mechanism that recovers the citation set needed for verification.
+For a sentence $s_{i}$, the LLM identifies the set of preceding sentences $R(s_{i})$ that $s_{i}$ semantically depends on, and defines the valid citation set used for verification as:
+$\mathcal{V}(s_{i})=\mathcal{C}(s_{i})\cup\bigcup_{k\in R(s_{i})}\mathcal{C}(s_{k}),$ (1)
+where $\mathcal{C}(s_{i})$ denotes the explicit citations of $s_{i}$.
+This enables verification by inheriting citations from previously referenced sentences, even when $\mathcal{C}(s_{i})=\emptyset$, substantially expanding verification coverage beyond explicitly cited sentences.
+Evidence-Grounded Verification and Metrics.
+For each verifiable claim (Types A–C), we retrieve the evidence documents (URLs) cited in $\mathcal{V}(s_{i})$ and assess whether they support the claim under a strict support criterion (Appendix F).
+We then summarize claim-level outcomes into fine-grained metrics aligned with the Integrity/Sufficiency subdimensions (e.g., Claim Factuality, Citation Support, Evidence Coverage, and Reference Reliability/Diversity) and aggregate them into the two dimension-level scores used in our final evaluation (Appendix F.4).
+Implementation details for long-document processing and efficiency-oriented engineering (e.g., batching and grouping) are provided in Appendix F.4.
 5 Experiment Setup
@@
 It assigns a 1–10 score and a brief rationale to each item and returns the results as a JSON object mapping each item to its score and rationale.
-In parallel, Information Integrity and Information Sufficiency are assessed by the Information Verification Module, which extracts claims from the report and verifies them against external evidence retrieved via web search, producing metrics such as Claim Factuality and Citation Support.
-Details are provided in Appendix G.4.
-Finally, we compute the overall quality score hierarchically by aggregating scores bottom-up from rubric items to higher-level categories.
-Details of the scoring procedure are provided in Appendix E.3.
-Figure 3: Heatmap visualizations of expert report evaluation results.
-(a) Criteria-wise scores across detailed evaluation categories.
-(b) Domain-wise scores averaged over two representative tasks randomly sampled from each domain.
+In parallel, Information Integrity and Information Sufficiency are assessed by the Information Verification Module, which extracts and type-classifies claims (Types A–F) and verifies the verifiable subset (Types A–C) against evidence documents retrieved from the report‘s cited sources (URLs), with semantic back-tracking to recover omitted citations for implicit claims.
+The module outputs quantitative metrics aligned with the Integrity/Sufficiency subdimensions (e.g., Claim Factuality, Citation Support, Evidence Coverage, Reference Reliability/Diversity), and hierarchically aggregates them into the two dimension-level scores (Appendix D.4).
+We report results using GPT-5.2 for Report Quality Assessment and GPT-5-mini for the Information Verification module, with an average evaluation cost of approximately \$0.5–\$1.0 per report.
+Figure 3: Heatmap visualizations of expert report evaluation results. (a) Criteria-wise scores across detailed evaluation categories. (b) Domain-wise scores averaged from each domain.
 5.2 Baseline Models
 To compare expert report generation performance, we consider four baseline families: General LLMs (fast), LLMs+Reasoning (think), LLMs+Reasoning+WebSearch (think+search), and Deep Research (deep).
-Each family is instantiated using multiple model families/backbones (Qwen, Gemini, Claude, and GPT), and they differ in whether they (i) use reasoning, (ii) use web search, and (iii) employ research-system orchestration.
-Detailed model configurations are provided in Appendix H.
+Each family is instantiated using multiple model families/backbones—Qwen, Gemini, Claude, and GPT (Yang et al., 2025; Google, 2025; Anthropic, 2025; OpenAI, 2025a), and they differ in whether they (i) use reasoning, (ii) use web search, and (iii) employ research-system orchestration.
+Detailed model configurations are provided in Appendix G.
 6 Experiments
@@
 Fulfillment, Analytical Soundness).
-This suggests that report-writing quality is largely mature, whereas expert-level intent alignment and evidence-based justification remain limited.
+This suggests that report-writing quality is largely mature, whereas expert-level intent alignment and reasoning completeness remain limited.
 Across baseline families, adding reasoning in think improves overall performance over fast.
@@
 Figure 3(b) shows domain-level performance of Deep Research agents.
-For each domain, we report results on two tasks randomly selected from that domain.
-Performance tends to drop uniformly across agents on relatively difficult tasks, and Physics consistently records the lowest scores compared to other domains.
+We observe that performance varies substantially across disciplines: agents generally achieve higher success in Philosophy, Psychology, and Engineering.
+Conversely, scores are notably lower in Computer Science, History, and Physics, reflecting the inherent difficulty of handling highly technical content, specific historical contexts, and complex scientific reasoning in these fields.
+Overall, these results demonstrate that our evaluation framework effectively captures fine-grained and domain-specific performance, enabling a granular diagnostic analysis that goes beyond simple aggregate scoring to identify the varying strengths and weaknesses of agents across diverse academic and technical fields.
 6.3 Correlation with Human Evaluation
@@
 We computed Pearson’s r, Spearman’s ρ, and LLM–human pairwise agreement for each of the five LLM-based evaluator models, and report the average of each metric across models.
-Additional details on the experimental setup and human evaluation protocol are provided in Appendix I.
+Additional details on the experimental setup and human evaluation protocol are provided in Appendix H.
 Table 2 shows how alignment with human judgments changes as evaluation components are added step by step.
@@
 Table 3: Inter-evaluator reliability across five LLM-based evaluation models. Krip. $\alpha$: Krippendorff’s alpha. Higher values indicate more consistent evaluations.
-To verify that the proposed evaluation method yields consistent results across different LLM evaluators, we measure inter-evaluator reliability Artstein (2017); Lee et al.
-(2025).
+To verify that the proposed evaluation method yields consistent results across different LLM evaluators, we measure inter-evaluator reliability (Artstein, 2017; Lee et al., 2025).
 Specifically, we compute Krippendorff’s $\alpha$ and the intraclass correlation coefficient (ICC) using the scores from the five LLM judges described in Section 6.3.As shown in Table 3, Vanilla achieves moderate inter-evaluator reliability, but reliability drops substantially under +Dimensions.
@@
 6.5 Information Verification Module Evaluation
-In this section, we evaluate the proposed Information Verification module by dividing it into two stages: claim extraction and claim verification.  [^2] Detailed content is covered in Appendix F.
+To evaluate the proposed Information Verification module, we present results on claim extraction and claim verification.
+These results quantify how our batch extraction and grouped verification design achieves strong performance while improving efficiency.
+Detailed results are provided in Appendix E and F.
 Claim Extraction Analysis
@@
 Furthermore, although our current benchmark focuses on text-based reports, this specialization enables a deeper, more rigorous analysis of information integrity and logical coherence, laying a solid foundation for future extensions to multimodal research tasks.
+Impact Statement
+This paper presents methods for improving the efficiency and reliability of machine learning model evaluation and training.
+The proposed approach may support more systematic model comparison and selection, which can influence downstream deployment decisions.
+As with any automated evaluation framework, inappropriate or uncritical use may lead to over-reliance on quantitative metrics without sufficient human judgment.
+We therefore emphasize that the proposed methods are intended to complement, not replace, human oversight in model development and evaluation.
 Appendix A Comparison with Deep Research Benchmarks
 In this section, we compare existing deep research benchmarks with our methodology.
-Table 6 extends and modifies the comparison table proposed in ResearchRubrics Sharma et al.
-(2025), and adds new evaluation axes to summarize how DEER differs from prior work.
+Table 6 extends and modifies the comparison table proposed in ResearchRubrics (Sharma et al., 2025), and adds new evaluation axes to summarize how DEER differs from prior work.
 The ResearchRubrics table uses five axes (whether rubrics are human-written, whether experts curate tasks, whether tasks are open-ended, whether non-technical domains are included, and whether an LLM-as-judge is used).
-On top of this, we add three axes that we consider essential from the perspective of deep research report evaluation: (1) Claim fact-check, (2) Interpretability, and (3) Systematic criteria.
-Systematic criteria indicate whether the evaluation criteria form a shared hierarchical ontology across tasks, grounded in external expert norms and guidelines.
+On top of this, we add two axes that are essential for deep research report evaluation: (1) Claim fact-check and (2) Interpretability.
+A.1 Benchmark Landscape
+Search and browsing agent benchmarks.
 The first block of the table consists of benchmarks for evaluating search and browsing agents.
-AcademicBrowse Zhou et al.
-(2025), BrowseComp Wei et al.
-(2025a), and Mind2Web2 Gou et al.
-(2025) evaluate, respectively, the ability to browse an academic corpus or the open web to produce short answers to complex queries, the ability to perform agentic search across diverse websites, and the consistency between generated answers and cited sources.
+AcademicBrowse (Zhou et al., 2025), BrowseComp (Wei et al., 2025a), and Mind2Web2 (Gou et al., 2025) evaluate, respectively, the ability to browse an academic corpus or the open web to produce short answers to complex queries, the ability to perform agentic search across diverse websites, and the consistency between generated answers and cited sources.
 These benchmarks focus on “how well the system finds information,” i.e., search and browsing strategies, and thus are one step removed from the deep research report quality evaluation that we study.
+Benchmarks for partial abilities or conceptualizations of deep research.
 The second block covers benchmarks that focus on specific component abilities or conceptualizations of deep research.
-ResearchBench Liu et al.
-(2025) and ResearcherBench Xu et al.
-(2025) evaluate, respectively, the ability to generate inspiration and hypotheses from papers, and to provide long-form advice on frontier AI research questions based on expert rubrics.
-Characterizing Deep Research Java et al.
-(2025) and LiveDRBench evaluate the recovery of correct claims in search tasks that require many information units and non-trivial reasoning.
-SPOT Son et al.
-(2025) measures how well a system can detect critical errors in published papers, and ExpertLongBench Ruan et al.
-(2025) evaluates the ability to generate long-form expert text without external search.
-These benchmarks finely probe partial abilities such as hypothesis generation, claim discovery, error detection, and expert writing, but do not directly evaluate the quality of full reports produced by web- or literature-based deep research agents.
+ExpertLongBench (Ruan et al., 2025) evaluates the ability to generate long-form expert text without external search.
+ResearchBench (Liu et al., 2025) evaluates the ability to extract inspirations from papers and generate research hypotheses.
+ResearcherBench (Xu et al., 2025) evaluates long-form responses to frontier AI research questions using a dual framework: expert-rubric insight quality and citation-based factuality (faithfulness/groundedness).
+ReportBench (Li et al., 2025a) and DeepScholar-Bench (Patel et al., 2025) assess academic survey/related-work reports, focusing primarily on literature selection and citation-grounded verifiability of report content.
+LiveDRBench (Java et al., 2025) evaluates the recovery of correct claims in search tasks that require many information units and non-trivial reasoning.
+SPOT (Son et al., 2025) measures how well a system can detect critical errors in published papers.
+These benchmarks finely evaluate partial abilities such as expert writing, hypothesis generation, claim discovery, and error detection, but it is difficult to view them as evaluating the overall quality of reports produced by web- or literature-based deep research agents.
+Benchmarks for deep research report quality.
 The third block targets benchmarks that evaluate the quality of deep research reports themselves.
-ReportBench Li et al.
-(2025a) and DeepScholar-Bench Patel et al.
-(2025) measure report quality in academic survey and related-work settings using citation coverage, overlap with gold surveys, and verifiability of citations.
-DeepResearchBench Du et al.
-(2025) and LiveResearchBench Wang et al.
-(2025) evaluate reports for multi-domain web-based deep research tasks using LLM-generated evaluation criteria (RACE), citation-based fact-checking (FACT), and dimensions such as coverage, depth, presentation, and citation accuracy.
-DeepResearchArena Wan et al.
-(2025) automatically constructs large numbers of research tasks from seminar transcripts, and measures factual alignment and overall quality via coverage/conflict/omission (KSR/KCR/KOR) against evidence keypoints extracted from cited URLs and LLM-generated task-specific rubrics (ACE).
-DeepResearchGym Coelho et al.
-(2025) provides an offline search environment over a public web corpus and an LLM-as-judge protocol, offering a sandbox for evaluating deep research agents’ reports with respect to alignment with information needs, search faithfulness, and overall quality.
-ResearchRubrics Sharma et al.
-(2025) provides 101 deep research prompts and over 2,500 human-written rubrics, and is designed so that an LLM-as-judge assigns per-item scores by following these detailed human criteria.
-Among these, we analyze DeepResearchBench, LiveResearchBench, and ResearchRubrics in more detail, as they are closest to DEER in terms of evaluation setting.
-DeepResearchBench and LiveResearchBench are most similar to our web-based deep research report evaluation setting, but both rely heavily on LLMs to define evaluation criteria and assign scores, and their fact-checking is applied only to a limited subset of claims with explicit citations.
-ResearchRubrics has the advantage that domain experts write 20–40 rubric items for each deep research prompt, enabling fine-grained capture of what constitutes a “good” answer, but it does not perform source-level claim fact-checking, and the rubric structure differs across prompts, making it challenging to interpret system behavior consistently along shared dimensions and rubrics.
-In contrast, DEER extracts, as exhaustively as possible, all explicit and implicit claims in a report that can be grounded in evidence documents and performs claim-level fact-checking over this set, and evaluates reports using a shared ontology of 7 high-level areas, 25 criteria, and 101 rubric items derived from expert report evaluation norms and guidelines, augmented with task-specific expert guidance.
-This allows DEER to avoid designing a new rubric from scratch for each task while still applying a dense set of rubric items to each report, and to interpret system performance along standard dimensions and rubric items across tasks.
+DeepResearchGym (Coelho et al., 2025) provides an offline web-corpus sandbox with an LLM-as-judge protocol.
+DeepResearchBench (Du et al., 2025) evaluates multi-domain web-based deep research reports using LLM-generated evaluation criteria (RACE), citation-based fact-checking (FACT), and dimensions including coverage, depth, presentation, and citation accuracy.
+Concurrent work.
+Concurrently with DEER, several benchmarks have further advanced report-level evaluation.
+DeepResearchArena (Wan et al., 2025) derives tasks from seminar transcripts and evaluates evidence–keypoint alignment (KSR/KCR/KOR) with task-specific checklists (ACE).
+RigorousBench (Yao et al., 2025) uses expert-curated queries and two-level human rubrics (GRR/QSR), and adds a trustworthiness signal via matching citations to curated trustworthy-source links (TSL).
+LiveResearchBench (Wang et al., 2025) evaluates web-based deep research reports in a live, multi-domain setting using LLM-judged criteria, e.g., presentation & organization, coverage & comprehensiveness, and citation accuracy.
+ResearchRubrics (Sharma et al., 2025) provides 2,500+ expert-written rubric items spanning axes such as explicit requirements, synthesis, and reference use, with mandatory vs.
+optional criteria per task.
+A.2 Key Differences from Prior Work
+Alignment with expert standards.
+A key concern in deep-research evaluation is whether the reported score truly reflects expert notions of report quality.
+In several benchmarks (Du et al., 2025; Wang et al., 2025; Wan et al., 2025), LLMs are used not only as judges but also to instantiate parts of the evaluation criteria (e.g., LLM-generated dimensions or task-specific checklists), which can leave ambiguity about how closely evaluation aligns with expert standards; moreover, even with well-defined criteria, LLM judges may apply them unreliably due to limited domain knowledge and weak evidence judgment.
+In contrast, DEER anchors evaluation in a shared rubric system grounded in established expert reporting norms and guidelines, and further provides task-specific expert guidance so that LLM-based scoring better aligns with expert judgment rather than ad hoc scoring heuristics.
+Claim-level fact checking.
+Some benchmarks (Du et al., 2025; Wang et al., 2025) perform citation-based verification only for cited claims, leaving uncited claims unchecked.
+Another approach (Wan et al., 2025) scores alignment to keypoints extracted from cited URLs, making verification conditional on what is cited and less suited to detecting missing evidence.
+In contrast, DEER extracts all claims from a report and, for each claim, (i) determines whether evidence is required, (ii) links not only explicitly cited sources but also recovers omitted citation links by tracing each claim back to earlier cited context in the report, and (iii) verifies whether the linked evidence supports the claim.
+For more detailed comparisons, see Table 7.
+Systematic interpretability.
+Beyond a single overall score, deep-research evaluation should support consistent diagnosis of failure modes across tasks.
+When benchmarks use task- or prompt-specific sub-criteria, fine-grained diagnostics are not standardized across tasks, so results tend to be interpretable mainly at coarse, high-level dimensions and are difficult to compare at a shared checklist level (Du et al., 2025; Wang et al., 2025; Wan et al., 2025; Sharma et al., 2025).
+DEER instead uses a hierarchical, shared rubric taxonomy grounded in established expert report-writing norms and guidelines, applying a fixed set of rubric sub-dimensions and items across tasks.
+As a result, DEER evaluates each report against a dense set of rubric items for more thorough, fine-grained assessment, and supports rubric-item-level diagnosis of system weaknesses.
 Benchmark Explicit Verif. Implicit Verif. Global Context Key Limitation vs. DEER
 LiveDRBench
-Java et al. (2025)
+(Java et al., 2025)
 Relies on matching against static Ground Truth claims; lacks dynamic verification of citations against web sources.
 DeepResearch Bench
-Du et al. (2025)
+(Du et al., 2025)
 Ignores uncited sentences; fails to detect hallucinations in transitional logic.
 DeepResearchGym
-Coelho et al. (2025)
+(Coelho et al., 2025)
 Limited to explicitly cited claims; only calculates citation precision/recall metrics.
 ResearcherBench
-Xu et al. (2025)
+(Xu et al., 2025)
 Treats uncited claims simply as “ungrounded” (empty URL) without attempting context-based verification.
 DeepResearch Arena
-Wan et al. (2025)
+(Wan et al., 2025)
 Evaluates Source $\to$ Report coverage (summarization) rather than Report $\to$ Source verification; misses uncited hallucinations.
 LiveResearchBench
-Wang et al. (2025)
+(Wang et al., 2025)
 Checklist- and judge-based evaluation depends on task-specific annotations and LLM judgment; implicit claims and cross-sentence dependencies are not systematically enumerated or verified at the claim level.
 SPOT
-Son et al. (2025)
+(Son et al., 2025)
 $\triangle$ Evaluates internal error detection against static ground truth; penalizes valid but unannotated error predictions (false positives).
 ReportBench
-Li et al. (2025a)
+(Li et al., 2025a)
 Verifies non-cited claims via external web search voting, ignoring internal document grounding.
 DeepScholar-Bench
-Patel et al. (2025)
+(Patel et al., 2025)
 Relies on physical distance (sliding window $w$); misses long-range semantic dependencies.
@@
 Based on this analysis, we finalized 11 topic domains: the 10 most frequent domains and an Others category that aggregates all remaining domains.
+We used only aggregated topic counts/statistics derived from in-house queries; no raw queries were released or included, and all queries complied with the organization’s privacy policy, including the removal of any personally identifiable information.
 B.2 HLE Subject Mapping
@@
 Therefore, using these queries directly as evaluation tasks would likely mismatch the high-difficulty, expert-level report-generation setting we target in both scope and format.
-Accordingly, we used Humanity’s Last Exam (HLE) Phan et al.
-(2025) as a source of expert-written, high-difficulty seed items that align with expert-level report generation.
+Accordingly, we used Humanity’s Last Exam (HLE) (Phan et al., 2025) as a source of expert-written, high-difficulty seed items that align with expert-level report generation.
 To preserve the 11-topic domain composition derived from actual deep research logs, we mapped each domain to one of HLE’s 13 subject domains and sampled HLE items from the corresponding subjects as domain-specific seeds.
@@
 Experts were individuals with a master’s degree in the relevant field or equivalent domain expertise.
-An example research task is provided in Box 5.
+An example research task is provided in Figure 5.
 Example Research Task (MDP / Value Iteration) Write a research report that explores the conditions under which value iteration in Markov Decision Processes converges geometrically to the optimal value function.
@@
 During cross-review, we checked and revised whether (i) the topic’s mandatory content elements were included at a sufficiently detailed level and written in an evaluable way, (ii) optional content or stylistic preferences were not included as mandatory evaluation elements, and (iii) requirements specified in the task prompt’s writing requirements (writing-direction instructions/additional requested constraints) were reflected without omission in the Guidance.
-An example Expert Evaluation Guidance is provided in Box 6.
+An example Expert Evaluation Guidance is provided in Figure 6.
 Example Expert Evaluation Guidance (MDP / Value Iteration) 1. Formal MDP and Value Iteration Setup
@@
 The following presents the mapping of which external standards were referenced for each evaluation dimension.
-Appendix D Foundations of Deep Research Report Evaluation Taxonomy
 Request Fulfillment
-Completeness aligns with guideline-level completeness checks that evaluate whether all required components of a report are present according to prescribed inclusion mandates across authoritative standards Bastian and Moher (2021); Boutron et al.
-(2010); International Organization for Standardization (2011); What Works Clearinghouse (2022); Percie du Sert et al.
-(2020); Global Reporting Initiative (2023); von Elm and others (2007).
-Scope reflects boundary-setting requirements that assess whether a report explicitly specifies its operative limits—such as eligibility conditions, assumptions, and constraints—within the structural fields defined in major evaluative frameworks Bastian and Moher (2021); National Institute of Standards and Technology (2023); American Historical Association (2024); American Economic Association (2024); Center for Open Science (2024); NASA (2016); U.S.
-Securities and Exchange Commission (2024); Chan and others (2013).
-Helpfulness corresponds to coherence and utility standards that evaluate whether the report’s core conclusions not only align logically with the evidence base but also deliver sufficient specificity and practical feasibility to comprehensively address the user’s inquiry Guyatt et al.
-(2013); American Psychological Association (2025); Chang et al.
-(2024); Institute of Education Sciences (2022); OECD (2020).
+Completeness aligns with guideline-level completeness checks that evaluate whether all required components of a report are present according to prescribed inclusion mandates across authoritative standards (Bastian and Moher, 2021; Boutron et al., 2010; International Organization for Standardization, 2011; What Works Clearinghouse, 2022; Percie du Sert et al., 2020; Global Reporting Initiative, 2023; von Elm and others, 2007).
+Scope reflects boundary-setting requirements that assess whether a report explicitly specifies its operative limits—such as eligibility conditions, assumptions, and constraints—within the structural fields defined in major evaluative frameworks (Bastian and Moher, 2021; National Institute of Standards and Technology, 2023; American Historical Association, 2024; American Economic Association, 2024; Center for Open Science, 2024; NASA, 2016; U.S.
+Securities and Exchange Commission, 2024; Chan and others, 2013).
+Helpfulness corresponds to coherence and utility standards that evaluate whether the report’s core conclusions not only align logically with the evidence base but also deliver sufficient specificity and practical feasibility to comprehensively address the user’s inquiry (Guyatt et al., 2013; American Psychological Association, 2025; Chang et al., 2024; Institute of Education Sciences, 2022; OECD, 2020).
 Analytical Soundness
-Quantification reflects evaluative requirements that check whether quantitative statements in a report correspond to verifiable computations, prespecified analytic procedures, and reproducible numerical evidence as established in major methodological frameworks Guyatt et al.
-(2013); Joint Committee for Guides in Metrology (2008); Association for Computing Machinery (2025); International Union of Pure and Applied Chemistry (2007); International Union of Pure and Applied Physics (2010); The Econometric Society (2024); American Educational Research Association et al.
-(2014); Mohr and others (2024); OECD (2011); International Accounting Standards Board (2024); International Organization for Standardization (2018); Intergovernmental Panel on Climate Change (2019).
-Reasoning aligns with reasoning-assessment criteria that evaluate whether inferential steps are explicitly grounded in evidence, free of unsupported leaps, and consistent with theoretical or mathematical rigor European Mathematical Society (2025); American Philosophical Association (2024); American Educational Research Association (2006); American Mathematical Society (2022); U.S.
-Congress (2024); Tong et al.
-(2007).
+Quantification reflects evaluative requirements that check whether quantitative statements in a report correspond to verifiable computations, prespecified analytic procedures, and reproducible numerical evidence as established in major methodological frameworks (Guyatt et al., 2013; Joint Committee for Guides in Metrology, 2008; Association for Computing Machinery, 2025; International Union of Pure and Applied Chemistry, 2007; International Union of Pure and Applied Physics, 2010; The Econometric Society, 2024; American Educational Research Association et al., 2014; Mohr and others, 2024; OECD, 2011; International Accounting Standards Board, 2024; International Organization for Standardization, 2018; Intergovernmental Panel on Climate Change, 2019).
+Reasoning aligns with reasoning-assessment criteria that evaluate whether inferential steps are explicitly grounded in evidence, free of unsupported leaps, and consistent with theoretical or mathematical rigor (European Mathematical Society, 2025; American Philosophical Association, 2024; American Educational Research Association, 2006; American Mathematical Society, 2022; U.S.
+Congress, 2024; Tong et al., 2007).
 Structural Coherence
-Introduction corresponds to structural-orientation requirements that assess whether a report’s opening section presents its scope and analytic pathway in accordance with prescribed organizational fields in established reporting and specification frameworks Bastian and Moher (2021); Boutron et al.
-(2010); IEEE Computer Society (2018); IFRS Foundation (2021); Gagnier and others (2013).
-Body aligns with structural-progression criteria that evaluate whether the main sections develop the promised analytical sequence without omission or drift relative to the ordered components defined in recognized guideline structures Bastian and Moher (2021); Boutron et al.
-(2010); IEEE Computer Society (2018); Eaton and others (2024); Gagnier and others (2013).
-Conclusion reflects closure-consistency requirements that examine whether final statements integrate preceding evidence without introducing unsupported expansions, consistent with the conclusion-governance rules embedded in major evaluative standards Bastian and Moher (2021); Boutron et al.
-(2010); Gagnier and others (2013).
-Section-Level corresponds to intra- and inter-section coherence checks that assess alignment, ordering, and non-duplication of content in accordance with structured specification and reporting templates in authoritative frameworks Bastian and Moher (2021); IEEE Computer Society (2018); IFRS Foundation (2021).
+Introduction corresponds to structural-orientation requirements that assess whether a report’s opening section presents its scope and analytic pathway in accordance with prescribed organizational fields in established reporting and specification frameworks (Bastian and Moher, 2021; Boutron et al., 2010; IEEE Computer Society, 2018; IFRS Foundation, 2021; Gagnier and others, 2013).
+Body aligns with structural-progression criteria that evaluate whether the main sections develop the promised analytical sequence without omission or drift relative to the ordered components defined in recognized guideline structures (Bastian and Moher, 2021; Boutron et al., 2010; IEEE Computer Society, 2018; Eaton and others, 2024; Gagnier and others, 2013).
+Conclusion reflects closure-consistency requirements that examine whether final statements integrate preceding evidence without introducing unsupported expansions, consistent with the conclusion-governance rules embedded in major evaluative standards (Bastian and Moher, 2021; Boutron et al., 2010; Gagnier and others, 2013).
+Section-Level corresponds to intra- and inter-section coherence checks that assess alignment, ordering, and non-duplication of content in accordance with structured specification and reporting templates in authoritative frameworks (Bastian and Moher, 2021; IEEE Computer Society, 2018; IFRS Foundation, 2021).
 Format & Style
-Report Format corresponds to format-governance requirements that evaluate whether a report adheres to prescribed structural conventions for scientific communication as codified in established editorial and reporting standards Bastian and Moher (2021); Boutron et al.
-(2010); International Committee of Medical Journal Editors (2025); Comrie, B.
+Report Format corresponds to format-governance requirements that evaluate whether a report adheres to prescribed structural conventions for scientific communication as codified in established editorial and reporting standards (Bastian and Moher, 2021; Boutron et al., 2010; International Committee of Medical Journal Editors, 2025; Comrie, B.
 and Haspelmath, M.
-and Bickel, B.
-(2015); Society for Industrial and Applied Mathematics (2024); International Phonetic Association (1999); American Political Science Association (2018); American Sociological Association (2022).
-Writing Quality aligns with language-precision criteria that assess clarity, specificity, and terminological consistency according to recognized guidelines for accurate and unbiased scholarly expression Boutron et al.
-(2010); International Committee of Medical Journal Editors (2025); University of Chicago Press (2024); Garner (2019, 2013).
-Paragraph Quality reflects cohesion-assessment rules that examine whether individual paragraphs follow a coherent internal logic—anchoring topic statements, supporting evidence, and transitional structure—in line with authoritative reporting and specification frameworks Bastian and Moher (2021); IEEE Computer Society (2018); Harvard Law Review Association (2020); American Sociological Association (2022).
-Readability corresponds to accessibility-oriented requirements that evaluate whether the narrative facilitates comprehension through appropriate signaling, explanatory devices, and presentation practices as outlined in major editorial and reporting guidelines Bastian and Moher (2021); International Committee of Medical Journal Editors (2025); Garner (2013); Global Reporting Initiative (2023).
+and Bickel, B., 2015; Society for Industrial and Applied Mathematics, 2024; International Phonetic Association, 1999; American Political Science Association, 2018; American Sociological Association, 2022).
+Writing Quality aligns with language-precision criteria that assess clarity, specificity, and terminological consistency according to recognized guidelines for accurate and unbiased scholarly expression (Boutron et al., 2010; International Committee of Medical Journal Editors, 2025; University of Chicago Press, 2024; Garner, 2019, 2013).
+Paragraph Quality reflects cohesion-assessment rules that examine whether individual paragraphs follow a coherent internal logic—anchoring topic statements, supporting evidence, and transitional structure—in line with authoritative reporting and specification frameworks (Bastian and Moher, 2021; IEEE Computer Society, 2018; Harvard Law Review Association, 2020; American Sociological Association, 2022).
+Readability corresponds to accessibility-oriented requirements that evaluate whether the narrative facilitates comprehension through appropriate signaling, explanatory devices, and presentation practices as outlined in major editorial and reporting guidelines (Bastian and Moher, 2021; International Committee of Medical Journal Editors, 2025; Garner, 2013; Global Reporting Initiative, 2023).
 Ethics & Compliance
-Sensitive Handling corresponds to ethical-screening requirements that evaluate whether sensitive domains are addressed with neutrality, respect, and responsible language in accordance with established editorial and reporting ethics standards International Committee of Medical Journal Editors (2025); Bastian and Moher (2021); Oral History Association (2024); Organization of American Historians (2018); British Psychological Society (2021); National Commission for the Protection of Human Subjects of Biomedical and Behavioral Research (1979); American Political Science Association (2012a).
-Safety & Impact aligns with harm-assessment provisions that examine whether potential adverse effects, misuse risks, or disproportionate impacts are identified and mitigated under recognized guidelines for responsible scientific communication International Committee of Medical Journal Editors (2025); Guyatt et al.
-(2013); National Research Council (2011); World Medical Association (2013); Ecological Society of America (2021).
-Perspective Balance reflects fairness-evaluation criteria that assess whether multiple viewpoints and counterpositions are represented without undue bias following principles of impartiality articulated in authoritative publication and reporting frameworks International Committee of Medical Journal Editors (2025); Bastian and Moher (2021); British Philosophical Association (2024); Australasian Association of Philosophy (2023); CFA Institute (2024); American Bar Association (2023); American Association for Public Opinion Research (2021).
+Sensitive Handling corresponds to ethical-screening requirements that evaluate whether sensitive domains are addressed with neutrality, respect, and responsible language in accordance with established editorial and reporting ethics standards (International Committee of Medical Journal Editors, 2025; Bastian and Moher, 2021; Oral History Association, 2024; Organization of American Historians, 2018; British Psychological Society, 2021; National Commission for the Protection of Human Subjects of Biomedical and Behavioral Research, 1979; American Political Science Association, 2012a).
+Safety & Impact aligns with harm-assessment provisions that examine whether potential adverse effects, misuse risks, or disproportionate impacts are identified and mitigated under recognized guidelines for responsible scientific communication (International Committee of Medical Journal Editors, 2025; Guyatt et al., 2013; National Research Council, 2011; World Medical Association, 2013; Ecological Society of America, 2021).
+Perspective Balance reflects fairness-evaluation criteria that assess whether multiple viewpoints and counterpositions are represented without undue bias following principles of impartiality articulated in authoritative publication and reporting frameworks (International Committee of Medical Journal Editors, 2025; Bastian and Moher, 2021; British Philosophical Association, 2024; Australasian Association of Philosophy, 2023; CFA Institute, 2024; American Bar Association, 2023; American Association for Public Opinion Research, 2021).
 Information Sufficiency
-Evidence Coverage aligns with evidence-provision requirements that evaluate whether claims are accompanied by verifiable supporting sources in accordance with established evidentiary and reporting guidelines Bastian and Moher (2021); Data Citation Synthesis Group (2014); Chang et al.
-(2024); International Society for Stem Cell Research (2025); CFA Institute (2020).
-Claim Amount corresponds to content-adequacy criteria that assess whether a report supplies all necessary factual and contextual material needed to justify its reasoning under recognized completeness and transparency standards Bastian and Moher (2021); International Organization for Standardization (2011); EQUATOR Network (2025); Linguistic Society of America (2024); The Journal of Organic Chemistry (2025); U.S.
-Geological Survey (2024); Global Reporting Initiative (2023).
-Citation Amount align with source-attribution rules that examine whether in-text citations are provided at appropriate argumentative locations following authoritative norms for evidentiary traceability International Committee of Medical Journal Editors (2025); Wilkinson et al.
-(2016); IEEE (2025); American Chemical Society (2021); Harvard Law Review Association (2020).
-Reference Amount are evaluated to determine whether the number of actually used references is appropriate and whether they collectively form a well-documented, traceable, and accessible evidence base Stanford HAI (2025); Data Citation Synthesis Group (2014); American Political Science Association (2012b).
+Evidence Coverage aligns with evidence-provision requirements that evaluate whether claims are accompanied by verifiable supporting sources in accordance with established evidentiary and reporting guidelines (Bastian and Moher, 2021; Data Citation Synthesis Group, 2014; Chang et al., 2024; International Society for Stem Cell Research, 2025; CFA Institute, 2020).
+Claim Amount corresponds to content-adequacy criteria that assess whether a report supplies all necessary factual and contextual material needed to justify its reasoning under recognized completeness and transparency standards (Bastian and Moher, 2021; International Organization for Standardization, 2011; EQUATOR Network, 2025; Linguistic Society of America, 2024; The Journal of Organic Chemistry, 2025; U.S.
+Geological Survey, 2024; Global Reporting Initiative, 2023).
+Citation Amount align with source-attribution rules that examine whether in-text citations are provided at appropriate argumentative locations following authoritative norms for evidentiary traceability (International Committee of Medical Journal Editors, 2025; Wilkinson et al., 2016; IEEE, 2025; American Chemical Society, 2021; Harvard Law Review Association, 2020).
+Reference Amount are evaluated to determine whether the number of actually used references is appropriate and whether they collectively form a well-documented, traceable, and accessible evidence base (Stanford HAI, 2025; Data Citation Synthesis Group, 2014; American Political Science Association, 2012b).
 Information Integrity
-Claim Factuality reflects evidence-verification provisions that assess whether factual assertions in a report correspond to verifiable sources and documented evidence traces within established evaluative frameworks Wilkinson et al.
-(2016); American Physical Society (2023); National Institute of Standards and Technology (2023); Navas and others (2024); National Institute of Standards and Technology (2024); U.S.
-Securities and Exchange Commission (2024).
-Citation Support corresponds to source-justification checks that evaluate whether cited references substantively support the claims they accompany according to recognized standards for evidential accountability Wilkinson et al.
-(2016); American Sociological Association (2018); American Historical Association (2024); Berez-Kroeker and others (2018).
-Reference Support aligns with source-quality and provenance requirements that examine whether referenced materials reliably and transparently support the arguments for which they are cited, in line with authoritative data-governance and reporting criteria Wilkinson et al.
-(2016); Chang et al.
-(2024); Data Citation Synthesis Group (2014); American Psychological Association (2025); International Accounting Standards Board (2024).
-Reference Quality evaluates whether the set of sources consists of reproducible, trustworthy, and methodologically sound information sources, ensuring that the report’s evidentiary foundation is grounded in reliable materials American Chemical Society (2021); IEEE (2025); Stanford Encyclopedia of Philosophy (2025); World Medical Association (2013); National Commission for the Protection of Human Subjects of Biomedical and Behavioral Research (1979).
-Reference Diversity evaluates whether the evidence base avoids excessive concentration on a small number of sources and instead maintains a balanced level of source diversity, reducing the risk of narrow, biased, or selectively framed arguments EQUATOR Network (2025); Linguistic Society of America (2024); American Association for Public Opinion Research (2021).
-To establish a robust and universally applicable Deep Research Report Evaluation Taxonomy, this framework comprehensively integrates authoritative standards and guidelines from a wide spectrum of disciplines, ranging from quantitative sciences and engineering to the humanities, social sciences, and professional fields such as law and finance (see Table 10).
+Claim Factuality reflects evidence-verification provisions that assess whether factual assertions in a report correspond to verifiable sources and documented evidence traces within established evaluative frameworks (Wilkinson et al., 2016; American Physical Society, 2023; National Institute of Standards and Technology, 2023; Navas and others, 2024; National Institute of Standards and Technology, 2024; U.S.
+Securities and Exchange Commission, 2024).
+Citation Support corresponds to source-justification checks that evaluate whether cited references substantively support the claims they accompany according to recognized standards for evidential accountability (Wilkinson et al., 2016; American Sociological Association, 2018; American Historical Association, 2024; Berez-Kroeker and others, 2018).
+Reference Support aligns with source-quality and provenance requirements that examine whether referenced materials reliably and transparently support the arguments for which they are cited, in line with authoritative data-governance and reporting criteria (Wilkinson et al., 2016; Chang et al., 2024; Data Citation Synthesis Group, 2014; American Psychological Association, 2025; International Accounting Standards Board, 2024).
+Reference Quality evaluates whether the set of sources consists of reproducible, trustworthy, and methodologically sound information sources, ensuring that the report’s evidentiary foundation is grounded in reliable materials (American Chemical Society, 2021; IEEE, 2025; Stanford Encyclopedia of Philosophy, 2025; World Medical Association, 2013; National Commission for the Protection of Human Subjects of Biomedical and Behavioral Research, 1979).
+Reference Diversity evaluates whether the evidence base avoids excessive concentration on a small number of sources and instead maintains a balanced level of source diversity, reducing the risk of narrow, biased, or selectively framed arguments (EQUATOR Network, 2025; Linguistic Society of America, 2024; American Association for Public Opinion Research, 2021).
+Overall, this framework comprehensively integrates authoritative standards and guidelines from a wide spectrum of disciplines, ranging from quantitative sciences and engineering to the humanities, social sciences, and professional fields such as law and finance (see Table 10).
 By synthesizing the core principles shared across these diverse domains, this taxonomy prioritizes criteria that are universally recognized as essential for high-quality intellectual work.
@@
 Table 10: References by Domain used in Deep Research Report Evaluation Taxonomy. (Total 80 listings from 78 unique sources; 2 interdisciplinary standards are cross-listed).
-Appendix E Rubric Structure
-In this appendix, we summarize the hierarchical structure and scoring method of the rubric used to evaluate the Deep Research report.
+Appendix D Rubric Structure
+In this appendix, we summarize the rubric taxonomy and scoring procedure used to evaluate Deep Research reports.
 Dim / Sub-dim Criterion / Rubric items Description
+1.
+Request Fulfillment └ 1.1 Completeness 1.1.1 Criterion Does the report include all required elements without omission and present each clearly?
+[1.1.1.1 (Coverage)] The report must include all elements required by the User Query and the Expert Evaluation Guidance (EG), and each element must be presented with clear and understandable explanations.
+Completeness is judged against the EG; if the explanation for any required element falls short of the EG standard, that element is considered omitted.
+[1.1.1.2 (Coverage)] Each major requirement from the User Query must be developed in at least one sufficiently substantial paragraph.
+This length must consist of explanations directly relevant to the User Query and the EG context; if the content is only filler without substantive relation, the requirement is considered unfulfilled.
+[1.1.1.3 (Quality)] For each required element, the report must be supported by appropriate evidence, reasoning, and validation, guided by the EG.
+[1.1.1.4 (Quality)] Each required element must be supported with sufficient evidence and depth.
 2.
 Analytical Soundness └ 2.2 Reasoning 2.2.5 Criterion All major claims follow logically from the previously presented facts, data, interpretations, and assumptions, without skipped steps or unsupported leaps.
-[2.2.5.1 Coverage] All claims that require logical support are explicitly linked to the relevant facts, data, and interpretations (including the key evidence specified in the Expert Evaluation Guidance, EG), and no central claim is left without the necessary supporting evidence.
-[2.2.5.2 Quality] Each major claim is explained through clear, well-structured reasoning that shows how the underlying facts, data, and interpretations lead to that conclusion, making the logical connection explicit and easy for an expert reader to follow.
-Table 11: Example Coverage/Quality rubric factors for a reasoning-focused element under 2. Analytical Soundness / 2.2 Reasoning.
-E.1 Hierarchical Levels
-We have the following 4-level hierarchical structure to systematically evaluate report quality in multiple dimensions.
+[2.2.5.1 (Coverage)] All claims that require logical support are explicitly linked to the relevant facts, data, and interpretations (including the key evidence specified in the EG), and no central claim is left without the necessary supporting evidence.
+[2.2.5.2 (Quality)] Each major claim is explained through clear, well-structured reasoning that shows how the underlying facts, data, and interpretations lead to that conclusion, making the logical connection explicit and easy for an expert reader to follow.
+3.
+Structural Coherence └ 3.1 Introduction 3.1.1 Criterion Does the introduction clearly present the report’s topic, problem, and significance, avoiding excessive generalization or irrelevant topic development?
+Does it also provide sufficient context and motivation for the reader?
+[3.1.1.1 (Coverage)] The introduction must include the report’s topic, problem, and significance, and provide sufficient background and motivation for the reader to understand the report’s context and rationale.
+[3.1.1.2 (Quality)] The introduction must be sufficiently developed for a professional report, and each component specified in 3.1.1.1 must be treated with adequate depth.
+[3.1.1.3 (Quality)] Each component must be described clearly and specifically, without excessive generalization or ambiguity.
+[3.1.1.4 (Quality)] The introduction must present its components in a logical, coherent flow so the reader can easily grasp the report’s overall direction.
+4.
+Format & Style └ 4.2 Writing Quality 4.2.3 Criterion Are technical terms defined when they first appear and used consistently thereafter?
+[4.2.3.1 (Coverage)] Technical terms and field-specific concepts must be defined when they are central to the argument, potentially ambiguous, or not guaranteed to be known by the intended audience.
+Well-established terms that are standard in the field do not require formal definitions if their meaning is clear from context.
+[4.2.3.2 (Coverage)] After being defined, technical terms must be used consistently with that definition throughout the document, including abbreviations and symbols.
+5.
+Ethics & Compliance └ 5.2 Safety & Impact 5.2.1 Criterion Are the potential impacts of proposed policies, technologies, strategies, or research outcomes sufficiently considered, including key implications, possible side-effects, and interpretations from multiple perspectives (when essential)?
+[5.2.1.1 (Coverage)] Potential side-effects or limitations are discussed.
+[5.2.1.2 (Coverage)] Multiple perspectives and relevant contextual considerations are included.
+[5.2.1.3 (Quality)] Key implications are presented in a balanced way, and relevant contexts are sufficiently considered.
+[5.2.1.4 (Quality)] Each identified impact is analyzed with adequate detail, supported by data, evidence, or clear reasoning.
+Table 11: Example criteria and rubric items from a four-level taxonomy (dimension, sub-dimension, criterion, and item: Coverage or Quality), with one example shown for each report-quality dimension.
+D.1 Hierarchical Levels
+We describe the rubric in two parts: (i) a shared taxonomy of dimensions and sub-dimensions, and (ii) the scoring instantiation for report-quality assessment, which specifies criteria and rubric items.
 Level 1: Evaluation Dimensions (7)
@@
 Level 2: Sub-dimensions (25)
-Detailed criteria that decompose each dimension into finer evaluation units; 2–5 sub-dimensions are defined under each dimension.
-This forms an intermediate-level unit that specifies which aspects to examine for each dimension.
-Level 3: Evaluation Elements (46)
-Evaluation elements that concretize each detailed criterion into units actually observable in the report.
-One or more elements are mapped under one detailed criterion, and are defined as evaluation units directly corresponding to report content, such as "Inclusion of requested items," "Specification of scope/limitations," and "Specificity of final message."
-Level 4: Rubric Factors (101)
-Rubric factors corresponding to the fourth level of the rubric are defined to evaluate each evaluation element (Level 3) from two perspectives.
-One element is subdivided into multiple factors from the perspectives of Coverage (C) and Quality (Q), and these factors serve as the minimum unit for actual scoring.
-Thus, with 7 dimensions, 25 sub-dimensions, and 46 elements as the framework, each element is divided into Coverage/Quality factors to constitute a total of 101 specific check items.
-Table 11 shows an example of one element and its corresponding Coverage/Quality factors in the 2.
-Analytical Soundness / 2.2 Reasoning dimension.
-E.2 Coverage and Quality Factors
-The rubric factors corresponding to Level 4 of the rubric evaluate each evaluation element (Level 3) from two perspectives: Coverage (C) and Quality (Q).
-The Coverage factor examines whether the element has been covered without omission as required throughout the document, i.e., how faithfully it includes content scattered in multiple locations or composed of multiple detailed requirements.
-On the other hand, the Quality factor evaluates whether the description possesses sufficient depth, logic, and rigor based on professional report standards, assuming the parts already written, i.e., “how well it is written.” By separating Coverage and Quality in this way, we can independently measure (1) what is missing and (2) what the level of the included content is in a long report.
-In our rubric, Level 4 consists of 101 factor-level items in total, comprising 66 Coverage items and 35 Quality items.
-Both factors use a 1–10 point scale, and sub-dimensions for scoring are presented in Table 12.
+Each dimension is decomposed into finer sub-dimensions (2–5 per dimension), specifying which aspects to examine within that dimension.
+D.2 Criteria and Rubric Items
+For the report-quality dimensions, we further specify criteria and rubric items.
+For the information-verification dimensions, we instead use metric-based measurements at the sub-dimension level; see Appendix F.4.
+Level 3: Criteria (46)
+Criteria operationalize each sub-dimension into evaluation requirements that are directly assessable from the report.
+One or more criteria are defined under each sub-dimension, and each criterion corresponds to a concrete aspect of report content, such as "Inclusion of requested items" and "Specification of scope/limitations."
+Level 4: Rubric Items (101)
+Rubric items at the fourth level decompose each criterion (Level 3) into atomic scoring units under two aspects, Coverage (C) and Quality (Q).
+Each rubric item is labeled as either (Coverage) or (Quality), and rubric items serve as the minimum unit for scoring.
+Table 11 shows example criteria and their associated rubric items.
+D.3 Coverage and Quality
+At Level 4, rubric items evaluate each criterion (Level 3) from two perspectives: Coverage (C) and Quality (Q).
+Coverage assesses whether the criterion is covered without omission as required throughout the document, including requirements distributed across multiple locations or composed of multiple detailed components.
+In contrast, Quality assesses whether the written content exhibits sufficient depth, logic, and rigor under professional report standards, conditioning on what is actually written (i.e., “how well it is written”).
+By separating Coverage and Quality, we can independently measure (1) what is missing and (2) the quality of what is included in a long report.
+In our rubric, Level 4 consists of 101 rubric items in total, comprising 66 Coverage items and 35 Quality items.
+Both use a 1–10 point scale, and the interpretation of score bands is summarized in Table 12.
 Score Range Coverage Quality
@@
 Table 12: Interpretation of 1–10 score ranges for Coverage (C) and Quality (Q) factors.
-E.3 Scoring and Aggregation
-The score aggregation flow in this evaluation framework consists of factor $\rightarrow$ element $\rightarrow$ criterion $\rightarrow$ dimension.
-First, rubric factors (the lowest items) are evaluated with a 1–10 score (or N/A).
-Coverage (C) and Quality (Q) are calculated separately up to the factor/element level, then integrated at the element level, and the resulting element score is raised to the criterion and dimension levels.
+D.4 Scoring and Aggregation
+The score aggregation flow in this evaluation framework consists of rubric item $\rightarrow$ criterion $\rightarrow$ sub-dimension $\rightarrow$ dimension.
+First, rubric items (the lowest-level units) are evaluated with a 1–10 score (or N/A).
+Coverage (C) and Quality (Q) are computed separately at the rubric-item level, then integrated at the criterion level, and the resulting criterion scores are aggregated to the sub-dimension and dimension levels.
 N/A items are excluded from all average calculations.
 Let $r$ be the report, and type $T\in\{C,Q\}$ represent Coverage and Quality, respectively.
-Let the score of each rubric factor $f$ be $s_{r,f}\in\{1,\dots,10\}$ (or N/A), and the set of non-N/A factors of type $T$ belonging to element $e$ be $F_{e,r}^{T}$.
-Let $E_{c,r}$ be the set of elements belonging to criterion $c$, and $\mathcal{C}_{d,r}$ be the set of criteria belonging to dimension $d$.
-Then, the factor $\rightarrow$ element $\rightarrow$ criterion $\rightarrow$ dimension aggregation is defined as follows:
-The Coverage/Quality score (element level) of report $r$ for element $e$ is
-$S_{r}^{T}(e)=\frac{1}{\lvert F_{e,r}^{T}\rvert}\sum_{f\in F_{e,r}^{T}}s_{r,f},\quad T\in\{C,Q\},$ (3)
-and if $F_{e,r}^{T}=\varnothing$, $S_{r}^{T}(e)$ is set to N/A.
-The integrated element score $S_{r}(e)$ of element $e$ is set as the average of defined values among C/Q.
-Let the set of types defined in element $e$ be $\mathcal{T}_{e}=\{T\in\{C,Q\}\mid S_{r}^{T}(e)\neq\text{N/A}\}$.
-$S_{r}(e)=\frac{1}{\lvert\mathcal{T}_{e}\rvert}\sum_{T\in\mathcal{T}_{e}}S_{r}^{T}(e),$ (4)
-If $\mathcal{T}_{e}=\varnothing$, $S_{r}(e)$ is set to N/A.
-The score of criterion $c$ is defined as the average of element scores belonging to that criterion.
-$S_{r}(c)=\frac{1}{\lvert E_{c,r}\rvert}\sum_{e\in E_{c,r}}S_{r}(e).$ (5)
-If $E_{c,r}=\varnothing$, $S_{r}(c)$ is set to N/A.
-The score of dimension $d$ is defined as the average of criterion scores belonging to that dimension.
-$S_{r}(d)=\frac{1}{\lvert\mathcal{C}_{d,r}\rvert}\sum_{c\in\mathcal{C}_{d,r}}S_{r}(c).$ (6)
-If $\mathcal{C}_{d,r}=\varnothing$, $S_{r}(d)$ is set to N/A.
-In summary, we calculate $S_{r}^{C}(e),S_{r}^{Q}(e)$ by averaging C/Q factors in each element, integrate them to create the element score $S_{r}(e)$, then average the element scores by criterion to calculate $S_{r}(c)$, and finally average the criterion scores to calculate the dimension score $S_{r}(d)$.
-Appendix F Human-based Information Verification Protocol
-F.1 Overview
+Let the score of each rubric item $i$ be $s_{r,i}\in\{1,\dots,10\}$ (or N/A), and let the set of non-N/A rubric items of type $T$ belonging to criterion $c$ be $I_{c,r}^{T}$.
+Let $\mathcal{C}_{s,r}$ be the set of criteria belonging to sub-dimension $s$, and let $\mathcal{S}_{d,r}$ be the set of sub-dimensions belonging to dimension $d$.
+Then, the rubric item $\rightarrow$ criterion $\rightarrow$ sub-dimension $\rightarrow$ dimension aggregation is defined as follows:
+The Coverage/Quality score (criterion level) of report $r$ for criterion $c$ is
+$S_{r}^{T}(c)=\frac{1}{\lvert I_{c,r}^{T}\rvert}\sum_{i\in I_{c,r}^{T}}s_{r,i},\quad T\in\{C,Q\},$ (2)
+and if $I_{c,r}^{T}=\varnothing$, $S_{r}^{T}(c)$ is set to N/A.
+The integrated criterion score $S_{r}(c)$ is set as the average of defined values among C/Q.
+Let the set of defined types for criterion $c$ be $\mathcal{T}_{c}=\{T\in\{C,Q\}\mid S_{r}^{T}(c)\neq\text{N/A}\}$.
+$S_{r}(c)=\frac{1}{\lvert\mathcal{T}_{c}\rvert}\sum_{T\in\mathcal{T}_{c}}S_{r}^{T}(c).$ (3)
+If $\mathcal{T}_{c}=\varnothing$, $S_{r}(c)$ is set to N/A.
+The score of sub-dimension $s$ is defined as the average of criterion scores belonging to that sub-dimension.
+$S_{r}(s)=\frac{1}{\lvert\mathcal{C}_{s,r}\rvert}\sum_{c\in\mathcal{C}_{s,r}}S_{r}(c).$ (4)
+If $\mathcal{C}_{s,r}=\varnothing$, $S_{r}(s)$ is set to N/A.
+The score of dimension $d$ is defined as the average of sub-dimension scores belonging to that dimension.
+$S_{r}(d)=\frac{1}{\lvert\mathcal{S}_{d,r}\rvert}\sum_{s\in\mathcal{S}_{d,r}}S_{r}(s).$ (5)
+If $\mathcal{S}_{d,r}=\varnothing$, $S_{r}(d)$ is set to N/A.
+In summary, we compute $S_{r}^{C}(c),S_{r}^{Q}(c)$ by averaging rubric items within each criterion, integrate them to obtain the criterion score $S_{r}(c)$, then average criterion scores to obtain the sub-dimension score $S_{r}(s)$, and finally average sub-dimension scores to obtain the dimension score $S_{r}(d)$.
+Appendix E Human-based Information Verification Protocol
+E.1 Overview
 DEER’s Information Verification Protocol is based on a stepwise information verification procedure performed by human evaluators.
 This section details the actual two-step procedure (Claim Extraction, Factual Accuracy Evaluation) performed by human evaluators.
-F.2 Step 1: Claim Extraction and Classification
+E.2 Step 1: Claim Extraction and Classification
 Human procedure.
-Evaluators segmented the report into sentences (in ‘Lx.Sy‘ format), reviewed each sentence individually to identify sentences containing claims, and extracted only the core claims.
+Evaluators segmented the report into paragraphs and sentences (in ‘Lx.Sy’ format, where L denotes the paragraph index and S denotes the sentence index), reviewed each sentence individually to identify sentences containing claims, and extracted only the core claims.
 Pronouns were replaced with explicit references according to the context, and if a single sentence contained multiple claims, they were separated.
@@
 Table 14: Example of human-annotated claims and corresponding LLM classification results.
-F.3 Step 2: Claim Verification
+E.3 Step 2: Claim Verification
 Human procedure.
@@
 This suggests that the proposed verification criteria are objective and reproducible.
-Appendix G LLM-based Information Verification Implementation
-G.1 Overview
+Appendix F LLM-based Information Verification Implementation
+F.1 Overview
 The Information Verification Module is designed to automate the human verification protocol described above.
@@
 In this process, algorithms such as Batch Extraction, Back-tracking, and Relevant Context Filtering were applied to achieve both cost efficiency and accuracy.
-G.2 Claim Extraction and Classification
+F.2 Claim Extraction and Classification
 LLM adaptation.
-The Information Verification Module automates the procedure of human evaluators.
-After tokenizing the report into sentences, the LLM evaluates the "claim likelihood" of each sentence and predicts A–F types based on the 6-Type definitions defined in Table 13.
-For Type B and C claims, evidence sentences are subsequently searched using Back-tracking.
+The Information Verification Module is designed to automate the human verification protocol described above.
+The LLM analyzes the report sentence by sentence to extract and classify claims, and, if necessary, retrieves external documents to verify their factual accuracy.
+In this process, algorithms such as Batch Extraction, Back-tracking, and Relevant Context Filtering were applied to achieve both cost efficiency and accuracy.
 Batch Extraction Strategy
-To mitigate the “Lost-in-the-Middle” phenomenon that occurs when processing long contexts and to increase cost efficiency, this study introduced a Batch Extraction strategy.
+To mitigate the “Lost-in-the-Middle (Liu et al., 2023a)” phenomenon that occurs when processing long contexts and to increase cost efficiency, this study introduced a Batch Extraction strategy.
 After dividing the entire report $D$ into sentence units $S=\{s_{1},s_{2},\dots,s_{m}\}$, they are processed in batches of a fixed size $B$ (in this study, $B=20$).
-Each batch processing provides the full report context ($D$) at the beginning of the prompt, but instructs the model to extract claims only for the sentences belonging to the current batch ($s_{i},\dots,s_{i+B-1}$).
-Batch Extraction Prompt Structure System: You are an expert fact-checker and claim extractor.
-User: # Full Report Context {Full_Report_Text} # Target Sentences to Extract Claims From L10.S1: … L10.S20: … Extract claims only from the target sentences above.
-Use the full report context for coreference resolution.
+Each batch of processing provides the full report context ($D$) at the beginning of the prompt, but instructs the model to extract claims only for the sentences in the current batch ($s_{i},\dots,s_{i+B-1}$).
+Figure 7 shows the simplified prompt structure used for batch extraction, along with an example JSON output.
+This method induces the model to focus on local sentences while remaining aware of the entire context, achieving human-level claim-extraction performance.
+Simplified Batch Extraction Prompt Structure System: You are an expert fact-checker and claim extractor.
+User:
+# Full Report Context
+{Full_Report_Text}
+# Target Sentences to Extract Claims From
+L10.S1: …
+L10.S20: …
+Extract claims only from the target sentences above. Use the full report context for coreference resolution.
 Example JSON Output
@@
   "evidence_position": null
-This method induces the model to focus on local sentences while remaining aware of the entire context, achieving human-level claim-extraction performance.
-Claim Extraction Evaluation Setup
-To rigorously evaluate the claim extraction performance, we employed an LLM-based Judge (GPT-5).
-Standard lexical metrics (e.g., ROUGE Lin (2004), Exact Match) are unsuitable for this task because the generated claims may differ in wording or granularity (e.g., one sentence split into multiple atomic claims) while preserving the same semantic meaning.
-We defined two key metrics: (1) Semantic Recall: Measures whether each ground truth claim is semantically covered by the set of extracted claims.
+Figure 7: Simplified batch extraction prompt structure and an example JSON output.
+Claim Extraction Evaluation Setup.
+To rigorously evaluate the claim extraction performance in Table 4, we employed an LLM-based Judge (GPT-5).
+Standard lexical metrics (e.g., ROUGE (Lin, 2004), Exact Match) are unsuitable for this task because the generated claims may differ in wording or granularity (e.g., one sentence split into multiple atomic claims) while preserving the same semantic meaning.
+We defined two key metrics: (1) Paragraph-level Semantic Recall: Measures whether each ground-truth claim is semantically covered by the extracted claims within the same source paragraph.
 The LLM Judge compares the ground truth claim against all candidate claims extracted from the same source sentence and determines if the core information is present.
@@
 The implementation code and LLM prompts used for the evaluation are included in the supplementary material.
-Comparison with Sliding Window Baseline.
-To validate the effectiveness of the LLM’s targeted evidence prediction, we compared it with a “Sliding Window” baseline on the subset of correctly classified B/C claims ($N=131$).
+Semantic Back-tracking for Citation Recovery
+The Information Verification Module uses a Backtracking algorithm to find evidence for claims without explicit citations (Types B and C).
+Type B (Same Section) and Type C (Previous Section) claims often share citations from previous sentences in context.
+The algorithm traces the ‘evidence_position‘ (location ID of the reference target sentence, e.g., L1.S3) recorded in the claim’s metadata and adds the citations held by that target sentence to the citations of the current claim.
+This restores omitted citation relationships and allows the corresponding source to be reviewed together in the subsequent verification step.
+Formally, for a set of claims $\mathcal{C}=\{c_{1},c_{2},\dots,c_{n}\}$, each claim $c_{i}$ has a position $p_{i}$, type $t_{i}$, explicit citations $R_{i}$, and a reference position $ref_{i}$ (if $t_{i}\in\{B,C\}$).
+The Back-tracking function $f_{backtrack}(c_{i})$ is defined as follows:
+$f_{backtrack}(c_{i})=\begin{cases}R_{j}&\text{if }t_{i}\in\{B,C\}\text{ and }\exists c_{j}\text{ s.t. }p_{j}=ref_{i}\\ \emptyset&\text{otherwise}\end{cases}$
+Finally, the citation set used for verification becomes $R^{\prime}_{i}=R_{i}\cup f_{backtrack}(c_{i})$.
+Semantic Back-tracking Evaluaton.
+To validate the effectiveness of the LLM’s targeted evidence prediction, we compared it with a “Sliding Window (Patel et al., 2025)” baseline on the subset of correctly classified B/C claims ($N=131$).
 The sliding window method collects all citations within a window of size $k$ centered on the claim.
@@
 Table 15: Baseline comparison on correctly predicted B/C claims ($N=131$). Backtracking achieves the best balance of precision, recall, and Jaccard.
-G.3 Claim Verification
-Semantic Back-tracking for Citation Recovery
-The Information Verification Module uses a Backtracking algorithm to find evidence for claims without explicit citations (Types B and C).
-Type B (Same Section) and Type C (Previous Section) claims often share citations from previous sentences in context.
-The algorithm traces the ‘evidence_position‘ (location ID of the reference target sentence, e.g., L1.S3) recorded in the claim’s metadata and adds the citations held by that target sentence to the citations of the current claim.
-This restores omitted citation relationships and allows the corresponding source to be reviewed together in the subsequent verification step.
-Formally, for a set of claims $\mathcal{C}=\{c_{1},c_{2},\dots,c_{n}\}$, each claim $c_{i}$ has a position $p_{i}$, type $t_{i}$, explicit citations $R_{i}$, and a reference position $ref_{i}$ (if $t_{i}\in\{B,C\}$).
-The Back-tracking function $f_{backtrack}(c_{i})$ is defined as follows:
-$f_{backtrack}(c_{i})=\begin{cases}R_{j}&\text{if }t_{i}\in\{B,C\}\text{ and }\exists c_{j}\text{ s.t. }p_{j}=ref_{i}\\ \emptyset&\text{otherwise}\end{cases}$
-Finally, the citation set used for verification becomes $R^{\prime}_{i}=R_{i}\cup f_{backtrack}(c_{i})$.
+F.3 Claim Verification
 Context Retrieval
@@
 Then, the relevance score $Sim(q,k_{j})$ between each chunk and the claim is calculated, and only the top $N$ (Top-K) chunks are selected and used as input for the verification model.
-In this study, the BM25 Robertson et al.
-(2009) and OpenAI’s text-embedding-3-large OpenAI was used as the embedding model: The selected chunk set $K_{selected}=\{k\in K\mid rank(Sim(q,k))\leq N\}$ is combined while maintaining the original document order to form the final context $C_{final}$.
+In this study, the BM25 (Robertson et al., 2009) and OpenAI’s text-embedding-3-large (OpenAI, ) was used as the embedding model: The selected chunk set $K_{selected}=\{k\in K\mid rank(Sim(q,k))\leq N\}$ is combined while maintaining the original document order to form the final context $C_{final}$.
 $C_{final}=\text{Concatenate}(\text{SortByPosition}(K_{selected}))$
@@
 The Information Verification Module’s automatic verifier is prompt-engineered to determine whether the given context supports the claim, identical to the human protocol above.
-The LLM infers whether the claim’s core content and numerical information match the source to make a final judgment.
+The LLM infers whether the claim’s core content and numerical information match the source, then makes a final judgment.
 Augmented Dataset and Robustness Evaluation
@@
 Ablation Study on Retrieval Parameters
-To identify the optimal configuration for the cost-efficient gpt-5-mini model, we conducted an ablation study varying Batch Size (10, 20), Reasoning Effort (Low, Medium, High), Retrieval Method (BM25 Robertson et al.
-(2009), OpenAI’s text-embedding-3-large OpenAI ), and Context Size (Top-K=2, 4).
+To identify the optimal configuration for the cost-efficient gpt-5-mini model, we conducted an ablation study varying Batch Size (10, 20), Reasoning Effort (Low, Medium, High), Retrieval Method (BM25 (Robertson et al., 2009), OpenAI’s text-embedding-3-large (OpenAI, )), and Context Size (Top-K=2, 4).
 Table 16 summarizes the results on both the Original and Adversarial (Augmented) datasets.
@@
 Evaluation: The reference explicitly states 46.2% efficiency, supporting the claim of "above 45%".
-$\Rightarrow$ Final Result: Supported
-G.4 Evaluation Metrics
-The Information Verification Module calculates the following detailed metrics to evaluate Integrity and Sufficiency.
+$\Rightarrow$
+Final Result: Supported
+F.4 Evaluation Metrics
+The evaluation metrics are designed to assess the Integrity and Sufficiency subdimensions by decomposing evidence use into complementary, claim-level signals.
+Rather than relying on a single aggregate score, we measure multiple failure modes of information use—factual incorrectness, unsupported attribution, unreliable or inaccessible sources, and insufficient evidence coverage—so that different weaknesses in evidence grounding can be diagnosed explicitly.
 Integrity Metrics
@@
     $\text{Score}=\frac{|\text{Reliable \& Supported References}|}{|\text{Used References}|}$
-  • Reference Diversity (Normalized HHI): Measures how evenly citations are distributed across used references using the Normalized Herfindahl-Hirschman Index Rhoades (1993).
+  • Reference Diversity (Normalized HHI): Measures how evenly citations are distributed across used references using the Normalized Herfindahl-Hirschman Index (Rhoades, 1993).
+We define $s_{i}$ as the share of citations for reference $i$ among total citations ($s_{i}=c_{i}/\sum c)$, and HHI as follows:
+    $\text{HHI}=\sum_{i=1}^{N}s_{i}^{2}$
+    Based on this, the Normalized HHI score (0–10) is calculated as:
     $\text{Score}=10\times\left(1-\frac{\text{HHI}-1/N}{1-1/N}\right)$
@@
   • Dimension Level: Average of criterion scores within each dimension.
-Appendix H Baseline Model Details
+Appendix G Baseline Model Details
 We use the following backbone model families in our experiments: Qwen3-235B, Gemini 2.5, Claude Opus 4.5, and GPT-5.
@@
 We exclude Gemini think+search because citation information is not provided in its outputs.
-For WebThinker, we use Qwen3-235B as an auxiliary model.
+For WebThinker (Li et al., 2025b), we use Qwen3-235B as an auxiliary model.
 For OpenAI Deep Research, we collect reports generated from the service environment as of August 2025.
-Appendix I Human-Correlation Experiment Setup
+Appendix H Human-Correlation Experiment Setup
 We measure alignment between LLM judges and expert human judgments on 45 reports.
@@
 For each task, we randomly sample three reports from those produced by five Deep Research systems—OpenAI Deep Research, Gemini 2.5 Pro Deep Research, Claude Opus 4.1 Deep Research, WebThinker, and Qwen3-235B Deep Research—resulting in $15\times 3=45$ reports overall.
-WebThinker Li et al.
-(2025b) uses Qwen3-235B as an auxiliary model.
+WebThinker (Li et al., 2025b) uses Qwen3-235B as an auxiliary model.
 Each report was independently evaluated by two domain experts matched to the report’s topic, supporting both LLM–human correlation and human–human reliability analyses (90 ratings total; 45 reports × 2 experts).
@@
 For each task, we compute Pearson $r$ and Spearman $\rho$ between human and model scores across the three reports, and report the mean over the 15 tasks.
-Following DeepBench Du et al.
-(2025), PA is the fraction of report pairs within a task for which the model’s relative preference matches the human relative preference; with three reports, each task has $\binom{3}{2}=3$ pairs.
+Following DeepBench (Du et al., 2025), PA is the fraction of report pairs within a task for which the model’s relative preference matches the human relative preference; with three reports, each task has $\binom{3}{2}=3$ pairs.
 We compute PA per task and report the mean across tasks.
@@
 For each setting, we compute the three metrics ($r$, $\rho$, and PA) independently for each of the five evaluator models, and then report the average across the models.
-Appendix J Ethical Considerations and Potential Risks
-DEER is designed as a research evaluation benchmark rather than a deployment or decision-making system, and its scores should be interpreted accordingly.
-Because the framework partially relies on LLM-based judges, evaluation outcomes may reflect systematic biases or limitations of the underlying models; however, this risk is mitigated through expert-authored rubrics, task-specific expert guidance, and validation against human expert judgments.
-Unless otherwise specified, reported results correspond to single-run evaluations under fixed configurations, rather than aggregates over multiple random seeds.
-DEER evaluates the quality and evidence alignment of generated reports, not the real-world correctness, safety, or deployability of the underlying claims or recommendations, and thus should not be used as an endorsement for practical application in high-stakes domains.
-In addition, while the information verification module checks both explicit and implicit claims against external sources, its reliability is inherently bounded by the quality, availability, and temporal stability of retrieved web evidence.
-Overall, DEER is intended for comparative research analysis and diagnostic evaluation, and any use beyond this scope should be accompanied by appropriate human oversight and domain-specific safeguards.
-Appendix K Information About Use of AI Assistants
-AI assistants were used in a limited and supportive manner during the preparation of this work.
-Specifically, they were employed to assist with coding-related tasks (e.g., drafting boilerplate scripts, refactoring, and debugging) and with writing-related support, such as grammar checking, stylistic refinement, and clarity improvement.
-Importantly, all conceptual contributions, experimental design, analysis, interpretation of results, and final manuscript were carried out by the authors.
+Expert contributors participated as paid contractors via a professional vendor; no sensitive personal data were collected, and participation followed the vendor’s consent and compensation policies.
+Appendix I Prompts
+This appendix provides the prompt templates used in our evaluation pipeline.
+Figure 8 shows an example evaluator prompt for the Request Fulfillment dimension, one of the report-quality dimensions.
+Figure 9 shows the full prompt for claim extraction and classification.
+Figure 9 shows the full prompt used for the claim extraction and classification task.
+Figure 10 shows the complete prompt for the claim verification and source reliability task.
+Evaluation Prompt Template for Request Fulfillment
+# 1. Overview
+Evaluate the report using the provided rubric, which operationalizes the requirements of a professional, expert-level long-form
+report.
+For each rubric item (e.g., C1-1, C1-2, Q1-1, Q1-2, …), provide systematic evaluation reasoning (including relevant evidence
+from the report) and assign an item-level score.
+Scores must be integers from 1 to 10. If the report contains no assessable material for a given item, enter “N/A” instead of a numeric score.
+# 2. Evaluation Method
+Evaluate each rubric item strictly using the provided rubric and the Expert Evaluation Guidance (EG), and do not make arbitrary
+judgments outside the rubric and EG.
+# 3. Expert Evaluation Guideline (EG)
+Expert Evaluation Guidance (EG) provides task-specific expert criteria: it enumerates required content elements and expert
+expectations as concrete, verifiable statements that can be checked directly against the report.
+The EG has absolute priority in every evaluation: each rubric item must be evaluated on the basis of the EG with all applicable EG
+requirements applied in full, and if EG requirements are not met, no high score (Perfect or Excellent) may be awarded regardless of
+supplementary strengths; supplementary strengths may only be considered once full compliance with the EG has been confirmed.
+## Expert Evaluation Guideline
+{expert guideline}
+# 4. Rubric Items
+[Omitted for brevity.]
+# 5. How to Score: Coverage (C) vs. Quality (Q)
+Each rubric item has either **C (Coverage) or Q (Quality)** attribute, and each is evaluated independently.
+## 5.1 Coverage (C) Evaluation
+This item is evaluated based on whether every required component is present and fully addressed wherever relevant in the report.
+**Evaluation Method (Coverage/C):**
+1. Identify all required elements for this rubric item.
+2. For each required element, evaluate the relevant parts of the report as Pass/Fail (met/not met).
+3. Classify Fails as core gaps vs minor omissions.
+4. Assign a 1–10 score based on the number and type of Fails.
+**Scoring Guidelines:**
+[Omitted for brevity; see Table 12.]
+**Core Principles:**
+* Even one core gap makes Excellent (7-8) impossible
+* Multiple core gaps make Good (5-6) impossible
+## 5.2 Quality (Q) Evaluation
+This item evaluates how well the report executes the relevant written content for the rubric item.
+**Evaluation Method:**
+1. Evaluate only the written parts relevant to this rubric item, and do not evaluate omissions.
+2. Make an overall-level judgment (academic/professional level) based on the written content, and assign a provisional 1–10 score.
+3. If a core element falls short of the provisional score level, adjust the final score downward to match the level of that core
+element.
+**Scoring Guidelines:**
+[Omitted for brevity; see Table 12.]
+**Core Principles:**
+* If an EG core element falls short, the overall Q score should be lowered accordingly; weaknesses in non-core elements have limited
+impact on the overall score.
+* Review the quality aspects specified by the rubric item (e.g., depth, logic, rigor, precision, clarity, accuracy, balance,
+methodological soundness).
+# 6. Output Format
+For each rubric item, write the "description" as score justification grounded in the scoring guidelines, not as a general pros/cons
+summary.
+(1) Band alignment: The description must match the score-band definition (Coverage or Quality) for the score you assigned.
+(2) Deficiency-focused: Describe what is inadequate or insufficient with concrete specifics, grounded in the report.
+(3) Consistency enforcement: Deficiency severity must be consistent with the assigned band; otherwise revise the score.
+[JSON schema omitted for brevity]
+[User Query]
+{query}
+[Expert Report]
+{doc}
+Figure 8: Abbreviated evaluator prompt template for Request Fulfillment. We omit the full rubric item list, detailed scoring bands, and the JSON output schema for space.
+Full Prompt for Claim Extraction and Classification
+You are an expert fact-checker. Your task is to extract distinct claims from the provided Target Sentences
+and classify each one into one of the following categories.
+Class Definitions:
+| Class | Definition |
+| A: Cited Claim | Claims that include explicit citations (e.g., [1] or (Kim, 2024)) directly within the sentence. |
+| B: Uncited - Same Section/Paragraph | Claims whose supporting citations appear within the same section or paragraph.
+Crucially, the Evidence Position MUST be in the same section or paragraph. If the evidence is in a previous section, it is Class C. |
+| C: Uncited – Previous Section/Paragraph | Claims whose supporting citations appear in previous sections or paragraphs. T1
+he Evidence Position MUST be in a previous section. |
+| D: Uncited – Recap/Structural | Simple mentions of introduction, conclusion, abstract, or structural summaries. |
+| E: Uncited – Citation not Required | Widely accepted facts, common knowledge, author’s own findings, subjective claims,
+experimental results, methodological contributions, etc. WARNING: Conclusions, results, or findings of the paper are NOT Class E.
+They are specific claims that require evidence (Class A/B/C) or are uncited (Class F). |
+| F: No-Citation – Unknown Source | Specific factual assertions that require verification but lack any identifiable
+supporting source in the text. If a claim is specific (e.g., "X has a radius of Y", "X is planar") and has no citation attached to it
+or the sentence immediately preceding/following it that covers this fact, it is Class F.
+Do not assume a nearby citation covers it unless it clearly does. |
+Instructions:
+1. Read the Report Context to understand the global context.
+2. Process the "Target Sentences":
+   - Break down the text into atomic claims. A single sentence may contain multiple claims
+   (e.g., "X is Y, and Z requires W" -> Claim 1: "X is Y", Claim 2: "Z requires W").
+   - Extract ALL statements, including facts, opinions, structural descriptions, and summaries.
+3. For each extracted claim, analyze its relationship with the context and citations:
+   - Step 1: Specificity Check. Contains numbers, chemical properties, specific results? -> Likely A, B, C, or F.
+   - Step 2: Citation Check.
+     - Citation in same sentence? -> Class A.
+     - Citation in same paragraph? -> Class B.
+     - Citation in previous section? -> Class C.
+   - Step 3: Uncited Check.
+     - Specific fact but no citation? -> Class F.
+     - General knowledge/methodology/author opinion? -> Class E.
+     - Structural summary? -> Class D.
+4. Determine Evidence Position:
+   - For Class B or C, identify the exact sentence index (e.g., "L1.S1") that contains the citation supporting this claim.
+   - Must contain an explicit citation.
+5. Output Format:
+   - Return a JSON object with a list of claims.
+   - Each claim must include: ‘position‘ (line/sent index from input), ‘claim‘ (text), ‘claim_type‘ (A-F), ‘rationale‘,
+   ‘numeric‘ (bool), ‘citations‘ (list of strings), ‘implicit_citations‘ (list), ‘cross_references‘ (list).
+Input Format:
+# Report Excerpt
+# Target Sentences
+L1.S1: ...
+L1.S2: ...
+Extraction and Classification Logic:
+- If a sentence is "Most perovskites are unstable[1], but our new material is stable.", extract TWO claims:
+  1. "Most perovskites are unstable." (Class A, citations=[’1’])
+  2. "Our new material is stable." (Class E - Author’s finding, or Class F if it lacks proof provided elsewhere)
+Examples:
+*Example Input:*
+L1.S1: Several studies[1] have shown that urban green spaces can reduce ambient air temperatures by up to 2 °C. This is crucial.
+*Example Output (Conceptual):*
+1. Claim: "Several studies have shown that urban green spaces can reduce ambient air temperatures by up to 2 °C."
+   - Class: A
+   - Citations: ["1"]
+   - Position: "L1.S1"
+2. Claim: "This is crucial."
+   - Class: B (supported by L1.S1)
+   - Evidence Position: "L1.S1"
+   - Position: "L1.S1"
+Figure 9: Full prompt for Claim Extraction and Classification.
+Full Prompt for Claim Verification
+You are an expert fact-checker. Verify the following claims against the provided context.
+Be extremely strict. High precision is required.
+For each claim, determine if it is supported by the context.
+Result should be based on semantic meaning, not just keyword matching. If the text implies the claim is true,
+mark it as supported.
+Verification Result (‘result‘):
+- supported: The text explicitly states or clearly implies the claim is true.
+- conflict: The text explicitly contradicts the claim.
+- not_supported: The text does not contain enough information.
+- error: Access/processing error (4xx/5xx, captcha, paywall, etc.)
+[Error Cases — Details]
+- Set result to "error" in the following cases:
+  - HTTP 4xx/5xx errors
+  - Access restrictions such as Captcha, Paywall, Authentication Required
+  - Other processing error messages
+- Briefly describe the specific error cause in explanation (e.g., "404 Not Found", "paywall restriction").
+[Document Reliability — Top-Level Fields]
+- reliable (true/false):
+  - Evaluate whether the entire URL is from a trustworthy source
+  - Example 1: Official statistics, academic journals, authoritative institutions -> true
+  - Example 2: Personal blogs, social media posts, etc. -> false
+- reliable_explanation (string or null):
+  - Rationale for reliability judgment (optional)
+# Examples:
+Claim: "GA+ prevents degradation."
+Context: "Guanidinium (GA+) ... helps passivate defects, stabilizing the structure against degradation."
+Result: supported
+# url: {url}
+# claims:
+{claims}
+# Context:
+{context}
+Figure 10: Full prompt for Claim Verification.